### v1

# Completeness and additive property for submeasures

Abstract.
Given an extended real-valued submeasure $\nu$ defined on a field of subsets $\Sigma$ of a given set, we provide necessary and sufficient conditions for which the pseudometric $d_{\nu}$ defined by $d_{\nu}(A,B):=\min\{1,\nu(A\bigtriangleup B)\}$ for all $A,B\in\Sigma$ is complete. As an application, we show that if $\varphi:\mathcal{P}(\omega)\to[0,\infty]$ is a lower semicontinuous submeasure and $\nu(A):=\lim_{n}\varphi(A\setminus\{0,1,\ldots,n-1\})$ for all $A\subseteq\omega$, then $d_{\nu}$ is complete. This includes the case of all weighted upper densities, fixing a gap in a proof by Just and Krawczyk in [Trans. Amer. Math. Soc. 285 (1984), 803–816]. In contrast, we prove that if $\nu$ is the upper Banach density (or an upper density greater than or equal to the latter) then $d_{\nu}$ is not complete.
Key words and phrases:
Completeness; submeasures; additive property; lower semicontinuous submeasures; upper Banach density.
2010 Mathematics Subject Classification:
Primary: 11B05, 54E50. Secondary: 28A12, 54A20.
1. Introduction
Let $\Sigma\subseteq\mathcal{P}(X)$ be a field of subsets of a nonempty set $X$, that is, a nonempty family of subsets of $X$ that is stable under finite unions and complements. Denote also by $\omega$ the set of nonnegative integers and by $\overline{\mathbb{R}}:=[-\infty,\infty]$ the two-point compactification of the reals, endowed with its natural total order. A map $\nu:\Sigma\to\overline{\mathbb{R}}$ is said to be a capacity if it is monotone (that is, $\nu(A)\leq\nu(B)$ for all $A,B\in\Sigma$ with $A\subseteq B$) and satisfies $\nu(\emptyset)=0$. If, in addition, the capacity $\nu$ is also subadditive (that is, $\nu(A\cup B)\leq\nu(A)+\nu(B)$ for all $A,B\in\Sigma$), then it is said to be a submeasure. Lastly, a capacity (or a submeasure) $\nu$ is called bounded if $\nu(X)<\infty$ and normalized if $\nu(X)=1$. It is clear that submeasures are necessarily nonnegative. If there is no risk of confusion, we will write also $\nu:\Sigma\to S$, for some $S\subseteq\overline{\mathbb{R}}$. Given a submeasure $\nu:\Sigma\to\overline{\mathbb{R}}$, it is easy to check that the map $d_{\nu}:\Sigma^{2}\to\mathbb{R}$ defined by
$\forall A,B\in\Sigma,\quad d_{\nu}(A,B):=\min\{1,\nu(A\bigtriangleup B)\}.$
is a pseudometric on the field $\Sigma$. The aim of this work is to provide necessary and sufficient conditions on $\nu$ such that the pseudometric space $(\Sigma,d_{\nu})$ is complete, hence to show completeness or non-completeness for large classes of submeasures.
Examples of normalized submeasures include the nonnegative finitely additive probability measures $\mu:\Sigma\to\mathbb{R}$ (for instance, the characteristic functions of free ultrafilters on $\omega$ and, of course, all the $\sigma$-additive probability measures), the upper asymptotic density $\mathsf{d}^{\star}:\mathcal{P}(\omega)\to\mathbb{R}$ defined by
$\forall A\subseteq\omega,\quad\mathsf{d}^{\star}(A):=\limsup_{n\to\infty}\frac{|A\cap n|}{n}$ (1)
(where, as usual, we identify each $n\in\omega$ with the set $\{0,1,\ldots,n-1\}$), and all upper densities on $\omega$ considered in [8, 15, 25], e.g. the upper analytic, upper Pólya, and upper Banach densities. In different contexts, we find also exhaustive submeasures [20, 33], nonpathological and lower semicontinuous submeasures [10, 31], and duals of exact games [28, 29].
1.1. Literature results: additive setting
Throughout, we reserve the symbol $\mu$ for finitely additive maps. The following result characterizes the completeness of the pseudometric space $(\Sigma,d_{\mu})$ in the case where the domain $\Sigma$ is a $\sigma$-field.
Theorem 1.1.
Let $\Sigma\subseteq\mathcal{P}(X)$ be a $\sigma$-field of subsets on a set $X$, and let $\mu:\Sigma\to\mathbb{R}$ be a nonnegative, finitely additive, bounded map. The following are equivalent:
  1. (i)
    The pseudometric space $(\Sigma,d_{\mu})$ is complete;
  2. (ii)
    For every increasing sequence $(A_{n}:n\in\omega)$ in $\Sigma$, there exists $A\in\Sigma$ such that $\mu(A_{n}\setminus A)=0$ for all $n\in\omega$ and $\mu(A)=\lim_{n}\mu(A_{n})$.
Proof.
See [18, Theorem 1] or [17, Lemma 2.2, Theorem 2.4, and Remark 2.5]. ∎
Additional equivalent conditions can be found in [18, Theorem 1], including completeness of the function space $L_{1}(\mu)$ or $L_{p}(\mu)$ for some $p\in[0,\infty)$ (see also [3, Theorem 3.4]). We refer the reader to the article [3] of Basile and Bhaskara Rao for a systematic study of completeness of $L_{p}(\mu)$ spaces and the related spaces $(\Sigma,d_{\mu})$. However, we will not consider these characterisations further here.
The property in item (ii) above is usually called AP0 and provides a weak form of continuity from below for $\mu$, cf. e.g. [5]. The next result shows that, even in the case $\Sigma=\mathcal{P}(\omega)$, one cannot replace the condition $\mu(A_{n}\setminus A)=0$ for all $n\in\omega$ in the latter property with $|A_{n}\setminus A|<\infty$ for all $n\in\omega$. To this aim, recall that the asymptotic density $\mathsf{d}(A)$ of a set $A\subseteq\omega$ is defined as in (1) by replacing $\limsup_{n\to\infty}$ with $\lim_{n\to\infty}$, provided that latter limit exists (cf. Section 2.2).
Theorem 1.2.
There exists an additive extension $\mu:\mathcal{P}(\omega)\to[0,1]$ of the asymptotic density $\mathsf{d}$ with the following properties:
  1. (i)
    The pseudometric space $(\mathcal{P}(\omega),d_{\mu})$ is complete;
  2. (ii)
    There exists an increasing sequence $(A_{n}:n\in\omega)$ in $\mathcal{P}(\omega)$ such that, if $A\subseteq\omega$ satisfies $|A_{n}\setminus A|<\infty$ for all $n\in\omega$, then $\mu(A)\neq\lim_{n}\mu(A_{n})$.
Proof.
It follows by [5, Corollary 4 or Theorem 5(b)] and Theorem 1.1. ∎
In contrast, the same authors also proved the following:
Theorem 1.3.
There exists an additive extension $\mu:\mathcal{P}(\omega)\to[0,1]$ of the asymptotic density $\mathsf{d}$ such that the pseudometric space $(\mathcal{P}(\omega),d_{\mu})$ is not complete.
Proof.
It follows by [5, Theorem 6] and the argument following its proof, together with Theorem 1.1. ∎
In both results above from [5], the authors consider special additive extensions of the asymptotic density $\mathsf{d}$ which are also called density measures, namely, maps $\mu^{\mathscr{F}}:\mathcal{P}(\omega)\to[0,1]$ defined by
$\forall A\subseteq\omega,\quad\mu^{\mathscr{F}}(A):=\mathscr{F}\text{-}\lim_{n\to\infty}\frac{|A\cap n|}{n},$ (2)
where $\mathscr{F}$ is a free ultrafilter on $\omega$. Kunisada proved in [22, Theorem 5.1] a necessary and sufficient condition for a density measure $\mu^{\mathscr{F}}$ to have AP0.
1.2. Literature results: non-additive setting
So far, only some rather special cases have been proved in the literature, all in the case $\Sigma=\mathcal{P}(\omega)$.
The first example is due to Solecki in his seminal work [31]: a map $\varphi:\mathcal{P}(\omega)\to[0,\infty]$ is said to be a lower semicontinuous submeasure (in short, lscsm) if it is a submeasure such that $\varphi(F)<\infty$ for all finite $F\subseteq\omega$ and
$\forall A\subseteq\omega,\quad\varphi(A)=\sup\{\varphi(A\cap n):n\in\omega\}.$
Notice that the above property is precisely the lower semicontinuity of the submeasure $\varphi$, regarding its domain $\mathcal{P}(\omega)$ as the Cantor space $2^{\omega}$, that is, if $A_{n}\to A$ then $\liminf_{n}\varphi(A_{n})\geq\varphi(A)$. Examples of lscsms include $\varphi(A)=|A|$ or $\varphi(A)=\sum_{n\in A}1/(n+1)$ or $\varphi(A)=\sup_{n\geq 1}|A\cap n|/n$, cf. also [10, Chapter 1].
Theorem 1.4.
Let $\varphi:\mathcal{P}(\omega)\to[0,\infty]$ be a lscsm. Then the pseudometric space $(\mathcal{P}(\omega),d_{\varphi})$ is complete.
Proof.
See the proof of implication (iii) $\implies$ (i) in [31, Theorem 3.1]. ∎
Of course, the upper asymptotic density $\mathsf{d}^{\star}$ defined in (1) is not a lscsm (since $\mathsf{d}^{\star}(A)=0$ for every finite $A\subseteq\omega$). However, it was recently proved that the analogous result holds for $\mathsf{d}^{\star}$:
Theorem 1.5.
The pseudometric space $(\mathcal{P}(\omega),d_{\mathsf{d}^{\star}})$ is complete.
Proof.
See [21, Theorem 1.1]. ∎
Just and Krawczyk claimed in [19, Lemma 3.1] that the analogue of Theorem 1.5 holds for all weighted upper densities, that is, maps $\nu_{f}:\mathcal{P}(\omega)\to\mathbb{R}$ defined by $\nu_{f}(A):=\limsup_{n}\sum_{i\in A\cap n}f(i)/\sum_{i\in n}f(i)$, where $f:\omega\to\mathbb{R}$ is an Erdős–Ulam function, that is, a nonnegative function such that $\lim_{n}\sum_{i\in n}f(i)=\infty$ and $\lim_{n}f(n)/\sum_{i\in n}f(i)=0$. It seems to us that their proof contains a gap that does not appear easily fixable. [^1] More precisely, the last centered upper bound of the distance $\rho_{h}(a_{n}/I_{h},a/I_{h})$ in the proof of [19, Lemma 3.1] is given by $2^{-n}+\sum_{k=1}^{\infty}2^{-(n+k)}$, which goes to 00 as $n\to\infty$. However, the latter should be replaced by $2^{-n}+k2^{-n}$, which does not necessarily go to 00 as $n\to\infty$ because in the construction there is no explicit dependence between $n$ and $k$. A further generalization was claimed by Farah in [10, Lemma 1.3.3(c)]. However, as confirmed by Farah in personal communications regarding the latter result, his argument also contains a gap. We show in Theorem 2.4 below, with a different proof, that their claims were, in fact, correct.
2. Main results
Our first main result characterizes the completeness of the pseudometric spaces $(\Sigma,d_{\nu})$, hence providing the analogue of Theorem 1.1 for submeasures.
Theorem 2.1.
Let $\Sigma\subseteq\mathcal{P}(X)$ be a field of subsets on a set $X$, and let $\nu:\Sigma\to\overline{\mathbb{R}}$ be a submeasure. Then the following are equivalent:
  1. (i)
    The pseudometric space $(\Sigma,d_{\nu})$ is complete;
  2. (ii)
    For every increasing sequence $(A_{n}:n\in\omega)$ in $\Sigma$ such that $\sum_{n}\nu(A_{n+1}\setminus A_{n})<\infty$, there exists $A\in\Sigma$ such that $\nu(A_{n}\setminus A)=0$ for all $n\in\omega$ and $\lim_{n}\nu(A\setminus A_{n})=0$.
Condition (ii) above is the analogue of AP0 for submeasures. Note that, even if $\nu(F)=0$ for all finite $F\in\Sigma$, the condition “$\nu(A_{n}\setminus A)=0$ for all $n\in\omega$”  cannot be replaced with “$|A_{n}\setminus A|<\infty$ for all $n\in\omega$,” (see Remark 3.1 below).
As a first immediate application, we show that Gangopadhyay’s characterization (namely, Theorem 1.1) holds, more generally, for fields of sets.
Corollary 2.2.
Let $\Sigma\subseteq\mathcal{P}(X)$ be a field of subsets on a set $X$, and let $\mu:\Sigma\to\mathbb{R}$ be a nonnegative, finitely additive, bounded map. The following are equivalent:
  1. (i)
    The pseudometric space $(\Sigma,d_{\mu})$ is complete;
  2. (ii)
    For every increasing sequence $(A_{n}:n\in\omega)$ in $\Sigma$, there exists $A\in\Sigma$ such that $\mu(A_{n}\setminus A)=0$ for all $n\in\omega$ and $\mu(A)=\lim_{n}\mu(A_{n})$.
Quite surprisingly, it seems that the above characterization cannot be found explicitly in the literature. As remarked by K. P. S. Bhaskara Rao in personal communications, [4, Theorem 3.3] is a relevant result in the finitely additive setting and may provide an alternative route to a direct proof of Corollary 2.2.
As another immediate application, we recover [21, Proposition 3.4]:
Corollary 2.3.
Let $\Sigma\subseteq\mathcal{P}(X)$ be a $\sigma$-field of subsets on a set $X$, and let $\nu:X\to\overline{\mathbb{R}}$ be a submeasure which is $\sigma$-subadditive. Then $(\Sigma,d_{\nu})$ is complete.
In Section 2.1, we provide a large class of submeasures $\nu:\mathcal{P}(\omega)\to[0,\infty]$ such that $(\mathcal{P}(\omega),d_{\nu})$ is complete (see Theorem 2.4), which will give a positive proof to the claims in [19, Lemma 3.1] and [10, Lemma 1.3.3(c)]. In Section 2.2, we provide another large class of submeasures $\nu:\mathcal{P}(\omega)\to[0,\infty]$ such that $(\mathcal{P}(\omega),d_{\nu})$ is not complete (see Theorem 2.5). Remarkably, this class includes the upper Banach density $\mathsf{bd}^{\star}$ (3) and the upper Buck density $\mathfrak{b}^{\star}$ (4) defined below. In both cases, some preliminaries will be necessary. Proofs of our results follow in Section 3.
2.1. Positive results: ideals and lscsms
Let $\mathcal{I}$ be an ideal on the nonnegative integers $\omega$, that is, a family of subsets of $\omega$ which is closed under subsets and finite unions. Unless otherwise stated, it is also assumed that $\mathcal{I}$ contains the family $\mathrm{Fin}$ of finite subsets of $\omega$, and that $\omega\notin\mathcal{I}$. Regarding ideals as subsets of the Cantor space $2^{\omega}$, we can speak about their topological complexity: for instance, $\mathcal{I}$ can be an analytic, Borel, or a $F_{\sigma}$-subset of $\mathcal{P}(\omega)$. In addition, we say that $\mathcal{I}$ is a $P$-ideal if it is $\sigma$-directed modulo finite sets, that is, for every sequence $(A_{n})$ of sets in $\mathcal{I}$ there exists $A\in\mathcal{I}$ such that $A_{n}\setminus A$ is finite for all $n\in\omega$. For instance, $\mathrm{Fin}$ is a $F_{\sigma}$ $P$-ideal, and the family of asymptotic density zero sets
$\mathcal{Z}:=\left\{A\subseteq\omega:\mathsf{d}^{\star}(A)=0\right\}$
is an analytic $P$-ideal which is not $F_{\sigma}$. Also, maximal ideals (i.e., the complements of free ultrafilters on $\omega$) do not have the Baire property, hence they are not analytic. We refer to [10] for an excellent textbook on the theory of ideals.
Given a lscsm $\varphi:\mathcal{P}(\omega)\to[0,\infty]$, define the family
$\mathrm{Exh}(\varphi):=\{S\subseteq\omega:\|S\|_{\varphi}=0\},\,\,\,\,\text{ where }\,\,\|S\|_{\varphi}:=\lim_{n\to\infty}\varphi(S\setminus n).$
Informally, $\|S\|_{\varphi}$ stands for the $\varphi$-mass at infinity of the set $S$. A classical result of Solecki [31, Theorem 3.1] states that an ideal $\mathcal{I}$ on $\omega$ is an analytic $P$-ideal if and only if there exists a lscsm $\varphi$ such that
$\mathcal{I}=\mathrm{Exh}(\varphi)\quad\text{ and }\quad\varphi(\omega)<\infty.$
We remark that the family of analytic $P$-ideals is large and includes, among others, all Erdős–Ulam ideals introduced by Just and Krawczyk in [19], ideals generated by nonnegative regular matrices [14, 16], the Fubini products ×$\emptyset\times\mathrm{Fin}$, which can be defined as $\{A\subseteq\omega:\forall n\in\omega,A\cap I_{n}\in\mathrm{Fin}\},$ where $(I_{n})$ is a given partition of $\omega$ into infinite sets, certain ideals used by Louveau and Velic̆ković [27], and, more generally, density-like ideals and generalized density ideals [7, 23]. Additional pathological examples can be found in [32]. It has been suggested in [6, 7] that the theory of analytic $P$-ideals may have some relevant yet unexploited potential for the study of the geometry of Banach spaces.
It is easy to check that, for each finite lscsm $\varphi$, the map $\|\cdot\|_{\varphi}$ is monotone, subadditive, and invariant under finite modifications, cf. [10, Lemma 1.3.3(a)-(b)]. In addition, the lscsm $\varphi$ in such a representation is not necessarily unique, cf. Example 3.3 below for the case of the ideal $\mathcal{Z}$. On a similar note, we will show in Proposition 3.5 that, if $\varphi_{1},\varphi_{2}$ are two lscsms with $\mathrm{Exh}(\varphi_{1})=\mathrm{Exh}(\varphi_{2})$, then the pseudometrics $d_{\|\cdot\|_{\varphi_{1}}}$ and $d_{\|\cdot\|_{\varphi_{2}}}$ are topologically equivalent, though not necessarily metrically equivalent (see Proposition 3.4).
With the above context, our main positive result follows:
Theorem 2.4.
Let $\varphi:\mathcal{P}(\omega)\to[0,\infty]$ be a lscsm. Then $(\mathcal{P}(\omega),d_{\|\cdot\|_{\varphi}})$ is complete.
Since $\mathcal{Z}$ is an analytic $P$-ideal, Theorem 2.4 provides a generalization of Theorem 1.5, cf. also Example 3.3 below. In addition, a special case of Theorem 2.4 follows by a result of Solecki in [31, Theorem 3.4], where he shows that the topology induced by $d_{\|\cdot\|_{\varphi}}$ is discrete if and only if $\mathrm{Exh}(\varphi)$ is a $F_{\sigma}$-ideal.
As remarked by Farah in personal communications, an instance of Theorem 2.4 can also be obtained as follows: suppose that $\varphi$ is a lscsm such that $\mathrm{Exh}(\varphi)$ is a generalized density ideal. Then the Boolean algebra $\mathcal{P}(\omega)/\mathrm{Exh}(\varphi)$ endowed with the metric induced by $d_{\|\cdot\|_{\varphi}}$ is countably saturated, see [13, Proposition 2.8] and cf. also [11, Theorem 16.5.1] and the argument preceeding [12, Corollary 6.5]; thus, completeness is a consequence of the latter property. However, it is an open question whether the analogous argument holds for all analytic $P$-ideals; see the last paragraph of [13, Section 2].
The density measures $\mu^{\mathscr{F}}$ in (2) used by Blass et al. [5] do not fall into the realm of Theorem 2.4: in fact, for each free ultrafilter $\mathscr{F}$ on $\omega$, the family $\{A\subseteq\omega:\mu^{\mathscr{F}}(A)=0\}$ is an ideal which is not analytic by the proof of [2, Theorem 2.10], hence it cannot coincide with $\mathrm{Exh}(\varphi)$ for any lscsm $\varphi$.
2.2. Negative results: upper densities
Following [25, Definition 1], we say that a map $\mu^{\star}:\mathcal{P}(\omega)\to\mathbb{R}$ is an upper density on $\omega$ if it satisfies the following properties:
  1. (f1)
    $\mu^{\star}(\omega)=1$;
  2. (f2)
    $\mu^{\star}(A)\leq\mu^{\star}(B)$ for all $A\subseteq B\subseteq\omega$;
  3. (f3)
    $\mu^{\star}(A\cup B)\leq\mu^{\star}(A)+\mu^{\star}(B)$ for all $A,B\subseteq\omega$;
  4. (f4)
    $\mu^{\star}(k\cdot A)=\mu^{\star}(A)/k$ for all $A\subseteq\omega$ and all nonzero $k\in\omega$;
  5. (f5)
    $\mu^{\star}(A+h)=\mu^{\star}(A)$ for all $A\subseteq\omega$ and $h\in\omega$.
It is readily seen that, if $\mu^{\star}$ is an upper density, then $0\leq\mu^{\star}(A)\leq 1$ for all $A\subseteq\omega$, and $\mu^{\star}(F)=0$ for every finite $F\subseteq\omega$, see [25, Proposition 2 and Proposition 6]; in particular, each $\mu^{\star}$ is a normalized submeasure.
Examples of upper densities include the upper asymptotic density $\mathsf{d}^{\star}$ defined in (1), the upper $\alpha$-densities with $\alpha\geq-1$ [25, Example 4], the upper Banach density $\mathsf{bd}^{\star}$ defined by
$\forall A\subseteq\omega,\quad\mathsf{bd}^{\star}(A):=\lim_{n\to\infty}\,\max_{k\in\omega}\,\frac{|A\cap[k,k+n)|}{n},$ (3)
the upper analytic density [25, Example 6], the upper Polya density [25, Example 8], and the upper Buck density $\mathfrak{b}^{\star}$ defined by
$\forall A\subseteq\omega,\quad\mathfrak{b}^{\star}(A):=\inf_{X\in\mathscr{A},A\subseteq X}\,\mathsf{d}^{\star}(A),$ (4)
where $\mathscr{A}$ stands for the family of finite unions of infinite arithmetic progressions $k\cdot\omega+h$ (with $k,h\in\omega$ and $k\neq 0$). With these premises, recall that
$\mathsf{d}^{\star}(A)\leq\mathsf{bd}^{\star}(A)\quad\text{ and }\quad\mu^{\star}(A)\leq\mathfrak{b}^{\star}(A)$ (5)
for all $A\subseteq\omega$ and all upper densities $\mu^{\star}$ on $\omega$, see [25, Theorem 3].
For any upper density $\mu^{\star}$ on $\omega$, we denote by $\mu_{\star}$ its lower dual, that is, the map defined by $\mu_{\star}(A):=1-\mu^{\star}(\omega\setminus A)$ for all $A\subseteq\omega$. We denote by $\mu$ the density induced by the pair $(\mu^{\star},\mu_{\star})$, that is, the restriction of $\mu^{\star}$ to the family $\mathrm{dom}(\mu):=\{A\subseteq\omega:\mu^{\star}(A)=\mu_{\star}(A)\}.$ For instance, if $\mu^{\star}$ is the upper asymptotic density $\mathsf{d}^{\star}$, then $\mathsf{d}_{\star}$ is the lower asymptotic density, $\mathsf{d}$ is the asymptotic density, and $\mathrm{dom}(\mathsf{d})$ is the family of sets $A\subseteq\omega$ which admits asymptotic density (that is, $\mathsf{d}^{\star}(A)=\mathsf{d}_{\star}(A)$).
Our main negative result follows:
Theorem 2.5.
Let $\mu^{\star}$ be an upper density on $\omega$ such that $\mathsf{bd}^{\star}(A)\leq\mu^{\star}(A)$ for all $A\subseteq\omega$. Then $(\mathcal{P}(\omega),d_{\mu^{\star}})$ is not complete.
Thanks to (5), the above result applies to the upper Banach density $\mathsf{bd}^{\star}$ and the upper Buck density $\mathfrak{b}^{\star}$. In addition, thanks to [25, Proposition 10], it applies also to the upper density $(\alpha(\textsf{bd}^{\star})^{q}+(1-\alpha)(\mathfrak{b}^{\star})^{q})^{1/q}$ for every $q\in[1,\infty)$ and every $\alpha\in[0,1]$ (more generally, the set of upper densities satisfying the hypothesis of Theorem 2.5 is countably $q$-convex for every $q\in[1,\infty))$.
3. Proofs of the main results in Section 2
Let us start with the proof of our main characterization.
Proof of Theorem 2.1.
(i) $\implies$ (ii). Let $S$ be the Stone space of the Boolean algebra $\Sigma$, that is, the set of ultrafilters on $\Sigma$, equipped with its usual topology. Recall that $S$ is compact and totally disconnected. Denote by $\mathcal{C}$ the field of all clopen (that is, simultaneously open and closed) subsets of $S$. By Stone’s representation theorem, there exists a Boolean isomorphism $\phi:\Sigma\to\mathcal{C};$ see e.g. [30, Chapter 1, Section 8] for a textbook exposition on Boolean algebras.
For each $B\subseteq S$ let $\Delta(B)$ be the family of all sequences $(A_{n}:n\in\omega)$ with values in $\Sigma$ such that $B\subseteq\bigcup_{n}\phi(A_{n})$ (of course, each $\Delta(B)$ is nonempty since $\phi(X)=S$). Thus, define the map $\hat{\nu}:\mathcal{P}(S)\to\overline{\mathbb{R}}$ by
$\forall B\subseteq S,\quad\hat{\nu}(B):=\inf\left\{\sum_{n\in\omega}\nu(A_{n}):(A_{n})\in\Delta(B)\right\},$
and observe that $\hat{\nu}$ is $\sigma$-subadditive by the very same argument used in the first part of the proof of [9, Lemma III.5.5].
First, we claim that
$\nu=\hat{\nu}\circ\phi.$ (6)
For, pick $A\in\Sigma$. Since $(A,\emptyset,\emptyset,\ldots)\in\Delta(\phi(A))$, we have $\hat{\nu}(\phi(A))\leq\nu(A)$. Conversely, suppose $(A_{n}:n\in\omega)\in\Delta(\phi(A))$. Since $\{\phi(A_{n}):n\in\omega\}$ is an open cover of the compact set $\phi(A)$, there exists $F\in\mathrm{Fin}$ such that $\phi(A)\subseteq\bigcup_{n\in F}\phi(A_{n})$. It follows that $A\subseteq\bigcup_{n\in F}A_{n}$, hence by monotonicity and subadditivity we get $\nu(A)\leq\sum_{n\in F}\nu(A_{n})\leq\sum_{n\in\omega}\nu(A_{n}).$ Therefore $\nu(A)\leq\hat{\nu}(\phi(A))$.
Now, let $(A_{n}:n\in\omega)$ be an increasing sequence with values in $\Sigma$ such that $\sum_{n}\nu(A_{n+1}\setminus A_{n})<\infty$. Define $B_{n}:=\phi(A_{n})$ for each $n\in\omega$, and $B:=\bigcup_{n}B_{n}$. It follows by the $\sigma$-subadditivity of $\hat{\nu}$ and identity (6) that
$\forall n\in\omega,\quad\hat{\nu}(B\setminus B_{n})\leq\sum_{k\geq n}\hat{\nu}(B_{k+1}\setminus B_{k})=\sum_{k\geq n}\nu(A_{k+1}\setminus A_{k}).$
Hence $\lim_{n}\hat{\nu}(B\setminus B_{n})=0$. Since $B_{n}\subseteq B$ for all $n\in\omega$, then $(B_{n})$ is $d_{\hat{\nu}}$-convergent to $B$. In particular, $(B_{n})$ is a $d_{\hat{\nu}}$-Cauchy sequence. Thanks to (6), the sequence $(A_{n})$ is $d_{\nu}$-Cauchy, and therefore converges to some $A\in\Sigma$. But then $(B_{n})$ converges also to $\phi(A)$, so that $\hat{\nu}(\phi(A)\bigtriangleup B)=0$. Using again (6), we get that
$\forall n\in\omega,\quad\nu(A_{n}\setminus A)=\hat{\nu}(B_{n}\setminus\phi(A))\leq\hat{\nu}(B\setminus\phi(A))=0.$
On a similar note, and using that $\hat{\nu}$ is a submeasure, we obtain
$\forall n\in\omega,\quad\nu(A\setminus A_{n})=\hat{\nu}(\phi(A)\setminus B_{n})\leq\hat{\nu}(\phi(A)\setminus B)+\hat{\nu}(B\setminus B_{n}),$
which implies that $\lim_{n}\nu(A\setminus A_{n})=0$.
(ii) $\implies$ (i). Let $(A_{n}:n\in\omega)$ be a Cauchy sequence in the pseudometric space $(\Sigma,d_{\nu})$. Without loss of generality, we can assume that $\nu(A_{i}\triangle A_{j})<2^{-i}$ for all $i,j\in\omega$ with $i\leq j$, since there is a subsequence with this property and a Cauchy sequence converges to the same limits as any subsequence. For all $i,j\in\omega$ with $i\leq j$, define the $\Sigma$-measurable set
$B_{i,j}:=\bigcap_{k=i}^{j}A_{k}$
and note that
$\nu(A_{i}\triangle B_{i,j})\leq\sum_{k=i}^{j-1}\nu(A_{k}\setminus A_{k+1})<2^{1-i}.$ (7)
In addition, for each $i\in\omega$, $(B_{i,j}^{c}:j\geq i)$ is an increasing sequence with
$\sum_{j\geq i}\nu(B_{i,j+1}^{c}\setminus B_{i,j}^{c})\leq\sum_{j\geq i}\nu(A_{j}\setminus A_{j+1})<2^{1-i}.$
Hence, by assumption, for each $i\in\omega$ there exists $B_{i}\in\Sigma$ such that $\nu(B_{i}\setminus B_{i,j})=\nu(B_{i,j}^{c}\setminus B_{i}^{c})=0$ for all $j\geq i$, and $\lim_{j}\nu(B_{i,j}\setminus B_{i})=\lim_{j}\nu(B_{i}^{c}\setminus B_{i,j}^{c})=0$. Thus
$\lim_{j\to\infty}\nu(B_{i,j}\bigtriangleup B_{i})=0.$ (8)
At this point, for all $i\in\omega$, define the $\Sigma$-measurable set
$C_{i}:=\bigcup_{k=0}^{i}B_{k}.$
Observe that, for all integers $i,k\in\omega$ with $k\geq i$, we have
$\begin{split}\nu(B_{i}\triangle C_{i})&\leq\sum_{j=0}^{i-1}\nu(B_{j}\setminus B_{j+1})\\ &\leq\sum_{j=0}^{i-1}(\nu(B_{j}\setminus B_{j,k})+\nu(B_{j,k}\setminus B_{j+1,k})+\nu(B_{j+1,k}\setminus B_{j+1}))\\ &=\sum_{j=0}^{i-1}\nu(B_{j+1,k}\setminus B_{j+1}).\end{split}$
By the hypothesis as applied to each $B_{j}$ and letting $k\to\infty$, this implies that
$\forall i\in\omega,\quad\nu(B_{i}\bigtriangleup C_{i})=0.$ (9)
Moreover, $(C_{i}:i\in\omega)$ is an increasing sequence and, for all $i,j\in\omega$ with $i\leq j$,
$C_{i+1}\setminus C_{i}$$\subseteq$ $B_{i+1}\setminus B_{i}$
$\subseteq$ $(B_{i+1}\setminus B_{i+1,j})\cup(B_{i+1,j}\setminus B_{i,j})\cup(B_{i,j}\setminus B_{i})$
$\subseteq$ $(B_{i+1}\setminus B_{i+1,j})\cup(A_{i+1}\setminus A_{i})\cup(B_{i,j}\setminus B_{i}).$
Since $\nu$ is a submeasure then
$\nu(C_{i+1}\setminus C_{i})\leq\nu(B_{i+1}\setminus B_{i+1,j})+\nu(A_{i+1}\setminus A_{i})+\nu(B_{i,j}\setminus B_{i})<2^{-i}$
for sufficiently large $j\in\omega$ and thus $\sum_{i}\nu(C_{i+1}\setminus C_{i})<\infty.$ By the standing assumption, there exists $A\in\Sigma$ such that $\nu(C_{i}\setminus A)=0$ for all $i\in\omega$, and $\lim_{i}\nu(A\setminus C_{i})=0$. In particular,
$\lim_{i\to\infty}\nu(C_{i}\bigtriangleup A)=0.$ (10)
To conclude, we claim that $(A_{n})$ is $d_{\nu}$-convergent to $A$. In fact, observe that for all $i,j\in\omega$ with $i\leq j$, we have
$A_{i}\triangle A\subseteq(A_{i}\triangle B_{i,j})\cup(B_{i,j}\triangle B_{i})\cup(B_{i}\triangle C_{i})\cup(C_{i}\triangle A).$
Since $\nu$ is a submeasure, it follows by inequalities (7), (8), and (9) that
$\forall i\in\omega,\quad\nu(A_{i}\triangle A)<2^{1-i}+\nu(C_{i}\triangle A).$
Finally, it follows by (10) that $\lim_{i}\nu(A_{i}\triangle A)=0$. Hence $(\Sigma,d_{\nu})$ is complete. ∎
We continue with the proofs of its applications.
Proof of Corollary 2.2.
Pick an increasing sequence $(A_{n})$ in $\Sigma$. Observe that $\sum_{n}\mu(A_{n+1}\setminus A_{n})<\infty$: otherwise there would exist $n_{0}\in\omega$ such that $\mu(X)<\sum_{n\leq n_{0}}\mu(A_{n+1}\setminus A_{n})=\mu(A_{n_{0}+1}\setminus A_{0})\leq\mu(A_{n_{0}+1})$, which is impossible. Suppose that there exists $A\in\Sigma$ such that $\mu(A_{n}\setminus A)=0$ for all $n\in\omega$ and $\lim_{n}\mu(A\setminus A_{n})=0$. Since $|\mu(A)-\mu(A_{n})|\leq\mu(A_{n}\setminus A)+\mu(A\setminus A_{n})$ for all $n\in\omega$, then $\lim_{n}\mu(A_{n})=\mu(A)$. Conversely, suppose that there exists $A\in\Sigma$ such that $\mu(A_{n}\setminus A)=0$ for all $n\in\omega$ and $\lim_{n}\mu(A_{n})=\mu(A)$. Since $\mu(A\setminus A_{n})=\mu(A)-\mu(A_{n})+\mu(A_{n}\setminus A)$, then $\lim_{n}\mu(A\setminus A_{n})=0$. This proves that item (ii) of Corollary 2.2 is a rewriting of item (ii) of Theorem 2.1 for nonnegative finitely additive bounded maps. The conclusion follows by Theorem 2.1. ∎
Proof of Corollary 2.3.
Pick an increasing sequence $(A_{n})$ in $\Sigma$ such that $\sum_{n}\nu(A_{n+1}\setminus A_{n})<\infty$ and define $A:=\bigcup_{k}A_{k}\in\Sigma$. Then $A_{n}\setminus A=\emptyset$ for all $n\in\omega$. In addition, since $A\setminus A_{n}=\bigcup_{k\geq n}(A_{k+1}\setminus A_{k})$ for all $n\in\omega$, we obtain
$\limsup_{n\to\infty}\nu(A\setminus A_{n})\leq\limsup_{n\to\infty}\sum_{j\geq n}\nu(A_{j+1}\setminus A_{j})=0.$
The conclusion follows by Theorem 2.1. ∎
Remark 3.1.
Item (ii) in Theorem 2.1 cannot be replaced by:
  1. (ii)
    For every increasing sequence $(A_{n}:n\in\omega)$ in $\Sigma$ such that $\sum_{n}\nu(A_{n+1}\setminus A_{n})<\infty$, there exists $A\in\Sigma$ such that $A_{n}\setminus A\in\mathrm{Fin}$ for all $n\in\omega$ and $\lim_{n}\nu(A\setminus A_{n})=0$.
In fact, choose $\Sigma=\mathcal{P}(\omega)$ and let $\mu$ be as in the statement of Theorem 1.2. Then $(\mathcal{P}(\omega),d_{\mu})$ is complete, i.e., item (i) in Theorem 2.1 holds. On the other hand, there exists an increasing sequence $(A_{n}:n\in\omega)$ in $\mathcal{P}(\omega)$ such that, if $A\subseteq\omega$ satisfies $A_{n}\setminus A\in\mathrm{Fin}$ for all $n\in\omega$, then $\mu(A)\neq\lim_{n}\mu(A_{n})$. Since $\mu$ is invariant modulo finite sets then there exists $\kappa>0$ such that $\mu(A)\geq\mu(A_{n})+\kappa$ for all $n\in\omega$. Since $\mu(A\setminus A_{n})=\mu(A)-\mu(A_{n})+\mu(A_{n}\setminus A)$ for all $n\in\omega$, we obtain $\liminf_{n}\mu(A\setminus A_{n})\geq\kappa$. In addition, $\sum_{n}\mu(A_{n+1}\setminus A_{n})<\infty$, by the same argument as in the proof of Corollary 2.2. Therefore item (ii) above fails.
Remark 3.2.
In the same vein as Remark 3.1, there exists a finitely additive map $\mu:\mathcal{P}(\omega)\to[0,1]$ satisfying property AP0 and such that its zero set :=$\mathcal{I}:=\{A\subseteq\omega:\mu(A)=0\}$ is not a $P$-ideal.
In fact, let $\mu$ be an additive extension of the asymptotic density $\mathsf{d}$ satisfying the claim of Theorem 1.2. Then by item (i) the induced pseudometric $d_{\mu}$ is complete, hence AP0 holds by Corollary 2.2. In addition, by item (ii), it is possible to pick an increasing sequence $(A_{n}:n\in\omega)$ in $\mathcal{P}(\omega)$ such that, if $B\subseteq\omega$ satisfies $A_{n}\setminus B\in\mathrm{Fin}$ for all $n\in\omega$, then $\mu(B)\neq\lim_{n}\mu(A_{n})$. Thanks to AP0, there exists $A\subseteq\omega$ such that $C_{n}:=A_{n}\setminus A\in\mathcal{I}$ for all $n\in\omega$ and $\mu(A)=\lim_{n}\mu(A_{n})$. Suppose for the sake of contradiction that $\mathcal{I}$ is a $P$-ideal. Then there exists $C\in\mathcal{I}$ such that $C_{n}\setminus C\in\mathrm{Fin}$ for all $n\in\omega$. Lastly, define $B:=A\cup C$. Then $A_{n}\setminus B=C_{n}\setminus C\in\mathrm{Fin}$ for all $n\in\omega$, and $\mu(B)=\mu(A)=\lim_{n}\mu(A_{n})$. This provides the claimed contradiction.
It is natural to investigate whether the topological and uniform properties of the pseudometric $d_{\|\cdot\|_{\varphi}}$ induced by an lscsm $\varphi$ can be deduced from some property of the exhaustive ideal $\mathrm{Exh}(\varphi)$. For example, below we identify two lscsms with the same exhaustive ideal that also induce metrically equivalent pseudometrics.
Example 3.3.
Define $I_{n}:=\omega\cap[2^{n},2^{n+1})$ for all $n\in\omega$, and let $\varphi:\mathcal{P}(\omega)\to[0,\infty]$ and $\psi:\mathcal{P}(\omega)\to[0,\infty]$ be the lscsms given by
$\varphi(A):=\sup_{n\geq 1}\,\frac{|A\cap n|}{n}\quad\text{ and }\quad\psi(A):=\sup_{n\in\omega}\,\frac{|A\cap I_{n}|}{|I_{n}|}$
for all $A\subseteq\omega$. It is clear that $\|A\|_{\varphi}=\mathsf{d}^{\star}(A)$ and $\|A\|_{\psi}=\limsup_{n}\,|A\cap I_{n}|/|I_{n}|$ for each $A\subseteq\omega$. Note
$\mathcal{Z}=\mathrm{Exh}(\varphi)=\mathrm{Exh}(\psi)$
(see [1, Lemma 3.1] or [24, Theorem 2]; cf. also the proof of [10, Theorem 1.13.3(a)], which shows that $\mathcal{Z}$ is a density-ideal).
In addition, $d_{\|\cdot\|_{\psi}}$ and $d_{\mathsf{d}^{\star}}$ are metrically equivalent or, more precisely,
$\forall A\subseteq\omega,\quad\frac{\|A\|_{\psi}}{2}\leq\mathsf{d}^{\star}(A)\leq 16\|A\|_{\psi}.$
[…]
However, it is not always the case that lscsms with the same exhaustive ideal induce metrically equivalent pseudometrics.
Proposition 3.4.
There exists a lscsm $\varphi:\mathcal{P}(\omega)\to[0,\infty]$ such that:
  1. (i)
    $\mathrm{Exh}(\varphi)=\mathcal{Z}$;
  2. (ii)
    $\|\cdot\|_{\varphi}$ is an upper density;
  3. (iii)
    $d_{\|\cdot\|_{\varphi}}$ is not metrically equivalent to $d_{\mathsf{d}^{\star}}$.
Proof.
For each real $\alpha>-1$, let $\varphi_{\alpha}:\mathcal{P}(\omega)\to[0,\infty]$ be the lscsm defined by
$\forall A\subseteq\omega,\quad\varphi_{\alpha}(A):=\sup_{n\geq 1}\frac{\sum_{i\in A\cap[1,n]}i^{\alpha}}{\sum_{i=1}^{n}i^{\alpha}}.$
Then $\varphi_{0}$ coincides with the lscsm $\varphi$ used in Example 3.3 and $\|\cdot\|_{\varphi_{\alpha}}$ is the classical upper $\alpha$-density, cf. [25, Example 4]. In addition, it follows by [26, Theorem 4.1 and Theorem 4.3] that $\mathcal{Z}=\mathrm{Exh}({\varphi_{\alpha}})$ for each $\alpha>-1$. Thus, define the map $\varphi_{\infty}:\mathcal{P}(\omega)\to[0,\infty]$ by
$\forall A\subseteq\omega,\quad\varphi_{\infty}(A):=\sum_{\alpha\in\omega}\frac{\varphi_{2^{\alpha}}(A)}{2^{\alpha}}.$
Then $\varphi_{\infty}$ is a lscsm that satisfies our claim.
It is immediate to see that $\varphi_{\infty}$ is monotone, subadditive, and $\varphi_{\infty}(\omega)=1$. Moreover, for each $\varepsilon>0$ and $\alpha>-1$ there exists $n_{\alpha,\varepsilon}\in\omega$ such that $\varphi_{2^{\alpha}}(A)\leq\varphi_{2^{\alpha}}(A\cap n_{\alpha,\varepsilon})+\varepsilon$ for $n\geq n_{\alpha,\varepsilon}$ (since $\varphi_{2^{\alpha}}$ is an lscsm). Now, fix $\varepsilon>0$, pick $\alpha_{0}\in\omega$ such that $32^{-\alpha_{0}}\leq\varepsilon/$ 3, and define $n_{\varepsilon}:=\max\{n_{\alpha,\varepsilon/3}:\alpha\in\alpha_{0}+1\}$. It follows that
$\begin{split}\varphi_{\infty}(A)&\leq\sum_{\alpha=0}^{\alpha_{0}}\frac{\varphi_{2^{\alpha}}(A)}{2^{\alpha}}+\sum_{\alpha>\alpha_{0}}\frac{1}{2^{\alpha}}=\sum_{\alpha=0}^{\alpha_{0}}\frac{\varphi_{2^{\alpha}}(A)}{2^{\alpha}}+\frac{1}{2^{\alpha_{0}}}\\ &\leq\sum_{\alpha=0}^{\alpha_{0}}\frac{\varphi_{2^{\alpha}}(A\cap n_{\alpha,\varepsilon/3})+\varepsilon/3}{2^{\alpha}}+\frac{\varepsilon}{3}\\ &\leq\sum_{\alpha=0}^{\alpha_{0}}\frac{\varphi_{2^{\alpha}}(A\cap n_{\varepsilon})}{2^{\alpha}}+\frac{\varepsilon}{3}\left(1+\sum_{\alpha=0}^{\alpha_{0}}\frac{1}{2^{\alpha}}\right)\leq\varphi_{\infty}(A\cap n_{\varepsilon})+\varepsilon.\end{split}$
This proves that $\varphi_{\infty}$ is a (bounded) lscsm.
By the Dominated Convergence theorem, we have that
$\forall A\subseteq\omega,\quad\|A\|_{\varphi_{\infty}}=\lim_{n\to\infty}\sum_{\alpha\in\omega}\frac{\varphi_{2^{\alpha}}(A\setminus n)}{2^{\alpha}}=\sum_{\alpha\in\omega}\frac{\|A\|_{\varphi_{2^{\alpha}}}}{2^{\alpha}}.$ (11)
Since each $\|\cdot\|_{\varphi_{\alpha}}$ is an upper density and the set of upper densities is countably convex by [25, Proposition 10], then $\|\cdot\|_{\varphi_{\infty}}$ is an upper density as well, i.e., item (ii) holds. Moreover, by the above observations, identity (11) implies that $\mathcal{Z}=\mathrm{Exh}(\varphi_{\infty})$, i.e., item (i) holds.
Lastly, we claim that there is no constant $C>0$ such that $\|A\|_{\varphi_{\infty}}\leq C\|A\|_{\psi}$ for all $A\subseteq\omega$, where $\psi$ is the lscsm used in Example 3.3. For, define $A_{n}:=\omega\cap\bigcup_{i\in\omega}[2^{i},2^{i}(1+2^{-n}))$ for each $n\in\omega$, and observe that $\|A_{n}\|_{\psi}=2^{-n}$. Then, for each $n\in\omega$ and each $\alpha>-1$, we obtain
$\begin{split}\|A_{n}\|_{\varphi_{\alpha}}&=\limsup_{k\to\infty}\frac{\sum_{i\in A_{n}\cap[1,2^{k}(1+2^{-n})]}i^{\alpha}}{\sum_{i=1}^{2^{k}(1+2^{-n})}i^{\alpha}}\\ &\geq\limsup_{k\to\infty}\frac{\sum_{i=2^{k}}^{2^{k}(1+2^{-n})}i^{\alpha}}{\sum_{i=1}^{2^{k}(1+2^{-n})}i^{\alpha}}=1-(1+2^{-n})^{-{\alpha}-1}.\end{split}$
Putting it all together, it follows that
$\forall\alpha>-1,\quad\liminf_{n\to\infty}\frac{\|A_{n}\|_{\varphi_{\alpha}}}{\|A_{n}\|_{\psi}}\geq\lim_{n\to\infty}\frac{1-(1+2^{-n})^{-\alpha-1}}{2^{-n}}=\alpha+1.$ (12)
Now, fix $C\in\omega$. Thanks to (12), there exists a sufficiently large $n_{C}\in\omega$ such that
$\forall\alpha\in 2C,\quad\frac{\|A_{n_{C}}\|_{\varphi_{2^{\alpha}}}}{\|A_{n_{C}}\|_{\psi}}\geq\frac{2^{\alpha}+1}{2}.$
Since $\frac{2^{\alpha}+1}{2}>2^{\alpha-1}$, it follows by (11) that
$\|A_{n_{C}}\|_{\varphi_{\infty}}\geq\sum_{\alpha\in 2C}\frac{\|A_{n_{C}}\|_{\varphi_{2^{\alpha}}}}{2^{\alpha}}>\|A_{n_{C}}\|_{\psi}\sum_{\alpha\in 2C}\frac{2^{\alpha-1}}{2^{\alpha}}=C\|A_{n_{C}}\|_{\psi}.$
Therefore $d_{\|\cdot\|_{\varphi_{\infty}}}$ and $d_{\|\cdot\|_{\psi}}$ are not metrically equivalent. Since $d_{\|\cdot\|_{\psi}}$ and $d_{\mathsf{d}^{\star}}$ are metrically equivalent by Example 3.3, we conclude that item (iii) holds. ∎
It is true that lscsms with the same exhaustive ideal induce pseudometrics that are topologically equivalent.
Proposition 3.5.
Let $\varphi_{1},\varphi_{2}$ be two lscsms such that $\mathrm{Exh}(\varphi_{1})=\mathrm{Exh}(\varphi_{2})$. Then the pseudometrics $d_{\|\cdot\|_{\varphi_{1}}}$ and $d_{\|\cdot\|_{\varphi_{2}}}$ are topologically equivalent.
Proof.
It is easy to check that $\mathrm{Exh}(\varphi_{1})=\mathrm{Exh}(\varphi_{2})$ is equivalent to the fact that, for every disjoint sequence $(F_{n})$ of finite sets of $\omega$, we have $\lim_{n}\varphi_{1}(F_{n})=0$ if and only if $\lim_{n}\varphi_{2}(F_{n})=0$, cf. [7, Remark 2.3]. Now, pick a sequence of sets $(A_{n})$ and $A\subseteq\omega$. Since $\varphi_{1}$ and $\varphi_{2}$ are two lscsms, there exists a partition $(P_{n}:n\in\omega)$ of $\omega$ into nonempty finite sets such that
$|\varphi_{i}((A_{n}\bigtriangleup A)\cap P_{n})-\|A_{n}\bigtriangleup A\|_{\varphi_{i}}|\leq 2^{-n}$
for each $i\in\{1,2\}$ and for all sufficiently large $n$. It follows that $\lim_{n}\|A\bigtriangleup A_{n}\|_{\varphi_{1}}=0$ if and only if $\lim_{n}\|A\bigtriangleup A_{n}\|_{\varphi_{2}}=0$, concluding the proof. ∎
Next we prove Theorem 2.4 using Theorem 2.1.
Proof of Theorem 2.4.
Let $(A_{n}:n\in\omega)$ be an increasing sequence of sets such that $\sum_{n}\|A_{n+1}\setminus A_{n}\|_{\varphi}<\infty$. Thanks to Theorem 2.1, it will be enough to show that there exists $A\subseteq\omega$ such that $A_{n}\setminus A$ is finite for all $n\in\omega$ and $\lim_{n}\|A\setminus A_{n}\|_{\varphi}=0$ (note that this is stronger than item (ii) in Theorem 2.1).
To this aim, define $B_{j}:=A_{j+1}\setminus A_{j}$ for all $j\in\omega$. Let $(n_{j}:j\in\omega)$ be a strictly increasing sequence in $\omega$ such that
$\forall j\in\omega,\quad\varphi(B_{j}\setminus n_{j})\leq 2\|B_{j}\|_{\varphi}.$ (13)
We claim that $A:=A_{0}\cup\bigcup_{j\in\omega}(B_{j}\setminus n_{j})$ satisfies our claim. For, note that $A_{k}\setminus A\subseteq n_{k}$ for all $k\in\omega$. To complete the proof, we need to show that $\lim_{n}\|A\setminus A_{n}\|_{\varphi}=0$. To this aim, fix $\varepsilon>0$ and pick $k_{0}\in\omega$ such that $\sum_{j\geq k_{0}}\|B_{j}\|_{\varphi}<\varepsilon/4$. Pick an integer $k\geq k_{0}$. Taking into account that $\|\cdot\|_{\varphi}$ is invariant under finite modifications and that the elements of $(B_{j})$ are pairwise disjoint, it follows that
$\begin{split}\|A\bigtriangleup A_{k}\|_{\varphi}\leq\left\|\,n_{k}\cup\bigcup_{j\geq k}((B_{j}\setminus n_{j})\setminus A_{k})\right\|_{\varphi}=\left\|\,\bigcup_{j\geq k}(B_{j}\setminus n_{j})\right\|_{\varphi}.\end{split}$ (14)
Since $\varphi$ is a lscsm, there exists an integer $m_{k}>n_{k+1}$ such that
$\left\|\,\bigcup_{j\geq k}(B_{j}\setminus n_{j})\right\|_{\varphi}\leq 2\varphi\left(m_{k}\cap\bigcup_{j\geq k}(B_{j}\setminus n_{j})\right).$ (15)
Fix an integer $i_{k}\in\omega$ such that $n_{i_{k}}>m_{k}$. Since $\min(B_{j}\setminus n_{j})\geq n_{j}>m_{k}$ for all $j\geq i_{k}$, it follows by (13), (14), (15), and the above observations that
$\begin{split}\|A\bigtriangleup A_{k}\|_{\varphi}&\leq 2\varphi\left(\bigcup_{k\leq j<i_{k}}(B_{j}\setminus n_{j})\right)\\ &\leq 2\sum_{k\leq j<i_{k}}\varphi(B_{j}\setminus n_{j})\leq 2\sum_{j\geq k}\varphi(B_{j}\setminus n_{j})\\ &\leq 4\sum_{j\geq k}\|B_{j}\|_{\varphi}\leq 4\sum_{j\geq k_{0}}\|B_{j}\|_{\varphi}<\varepsilon.\end{split}$
Therefore $\lim_{n}\|A\bigtriangleup A_{n}\|_{\varphi}=0$. ∎
Remark 3.6.
Solecki proved also in [31, Theorem 3.1] that, if $\varphi:\mathcal{P}(\omega)\to[0,\infty]$ is a lscsm, then $\mathrm{Exh}(\varphi)$ is $d_{\varphi}$-closed and separable. Since $d_{\|\cdot\|_{\varphi}}(A,B)\leq d_{\varphi}(A,B)$ for all $A,B\subseteq\omega$, then $\mathrm{Exh}(\varphi)$ is $d_{\|\cdot\|_{\varphi}}$-closed and separable as well. Alternatively, this can be seen directly: since $\|\cdot\|_{\varphi}$ is $d_{\|\cdot\|_{\varphi}}$-continuous, then the $\|\cdot\|_{\varphi}$-counterimage of $\{0\}$, that is, $\mathrm{Exh}(\varphi)$, is closed (in this case, the separability is obvious).
We conclude with the proof of our main negative result.
Proof of Theorem 2.5.
Let $\mu^{\star}$ be an upper density on $\omega$ such that $\mathsf{bd}^{\star}(A)\leq\mu^{\star}(A)$ for all $A\subseteq\omega$. Thanks to Theorem 2.1, it will be enough to show that for each $\kappa\in(0,1)$ there exists an increasing sequence $(A_{n}:n\in\omega)$ of subsets of $\omega$ such that $\sum_{n}\mu^{\star}(A_{n+1}\setminus A_{n})<\infty$ and, if a set $A\subseteq\omega$ satisfies $\mu^{\star}(A_{n}\setminus A)=0$ for all $n\in\omega$, then $\mu^{\star}(A\setminus A_{n})\geq\kappa$ for all $n\in\omega$.
To this aim, fix $\kappa\in(0,1)$ and a positive integer $N>e^{2}$ such that
$\frac{\log^{2}(N)}{N}<\frac{1-\kappa}{2}.$ (16)
Define the positive integer $C:=\lceil 1/\kappa\rceil$ and let $(a_{n}:n\in\omega)$ be a strictly increasing sequence of positive integers with the property that
$\sum_{j\in\omega}\frac{j}{a_{j}!}<\frac{1-\kappa}{2}\quad\text{ and }\quad a_{n}!>\max\left\{\frac{n}{\kappa},\,Ce^{2(n+1)},\,(N+1)^{2}\right\}$ (17)
for all $n\in\omega$. Thus, define recursively three sequences of sets of nonnegative integers $(A_{n}:n\in\omega)$, $(B_{n}:n\in\omega)$ and $(H_{n}:n\in\omega)$ as follows:
  1. (i)
    Set $H_{0}:=B_{0}:=\emptyset$.
  2. (ii)
    Set $H_{1}:=\{0\}$, $A_{0}:=B_{1}:=a_{1}!\cdot(\omega\setminus\{0\})$.
    Observe that $A_{0}=a_{2}!\cdot\omega+\{a_{1}!j:j\in\{1,2,\ldots,a_{2}!/a_{1}!\}\}$ and $\mathsf{d}(A_{0})=\mathsf{d}(B_{1})=1/a_{1}!$; in addition, we have $|A_{0}\cap n|/n<1/a_{1}!$ for all $n\geq 1$, hence choosing $n=a_{2}!$ we get
    $|a_{2}!\setminus A_{0}|>a_{2}!(1-1/a_{1}!)\geq a_{2}>2.$
  3. (iii)
    Given a positive integer $n\geq 1$, suppose that all sets $(A_{i}:i\in n)$, $(B_{i}:i\in n+1)$, and $(H_{i}:i\in n+1)$ have been defined satisfying the following properties:
    1. (a)
      all sets $B_{i}$ are pairwise disjoint;
    2. (b)
      $H_{i}\subseteq a_{i}!$ and $|H_{i}|=i$ for all $i\in n+1$;
    3. (c)
      $i\geq\kappa\ell_{i}$ for all nonzero $i\in n+1$, where
      $\ell_{i}:=1+\max H_{i}-\min H_{i};$
    4. (d)
      $B_{i}=a_{i}!\cdot(\omega\setminus\{0\})+H_{i}$ for all $i\in n+1$;
    5. (e)
      $A_{i}=B_{0}\cup\cdots\cup B_{i+1}$ for all $i\in n$;
    6. (f)
      $|A_{i}\cap m|/m<\mathsf{d}(A_{i})=\sum_{j\in i+2}j/a_{j}!$ for all $m\geq 1$ and $i\in n$;
    (In particular, such conditions hold for $n=1$, thanks to item (ii).)
  4. (iv)
    It follows by item (iii) that there exists a unique $K_{n+1}\subseteq a_{n+1}!$ such that
    $A_{n-1}\setminus a_{n+1}!=\bigcup_{i\in n+1}B_{i}\setminus a_{n+1}!=a_{n+1}!\cdot(\omega\setminus\{0\})+K_{n+1}$
    and, by hypothesis (17), also
    $a_{n+1}!-|K_{n+1}|=a_{n+1}!\left(1-\sum_{i\in n+1}\frac{i}{a_{i}!}\right)\geq\kappa a_{n+1}!>n+1.$
    Hence it is possible to pick a set $H_{n+1}\subseteq a_{n+1}!\setminus K_{n+1}$ such that $|H_{n+1}|=n+1$ and $\max H_{n+1}$ is minimized.
  5. (v)
    Lastly, define
    $B_{n+1}:=a_{n+1}!\cdot(\omega\setminus\{0\})+H_{n+1}\,\,\text{ and }\,\,A_{n}:=\bigcup_{i\in n+2}B_{i}.$ (18)
    To complete the induction, we have to show that the conditions in item (iii) hold for $n+1$. For, observe by item (iv) that $B_{n+1}\cap A_{n-1}=\emptyset$, hence $B_{n+1}\cap B_{i}=\emptyset$ for all $i\in n+1$. Together with the inductive step, this proves condition (a). Condition (b) follows by the construction in item (iv), while conditions (d) and (e) follow by the above definitions of $B_{n+1}$ and $A_{n}$ in (18). Since $B_{n+1}$ is a finite union of infinite arithmetic progressions (i.e., $B_{n+1}\in\mathscr{A}$) and it is disjoint from $A_{n-1}$, we obtain
    $\mathsf{d}(A_{n})=\mathsf{d}(A_{n-1})+\mathsf{d}(B_{n+1})=\sum_{j\in n+1}\frac{j}{a_{j}!}+\frac{|H_{n+1}|}{a_{n+1}!}=\sum_{j\in n+2}\frac{j}{a_{j}!},$
    cf. also [25, Proposition 7 and Proposition 8]. In addition, by the definitions in (18), it is not difficult to observe (since $B_{i}\cap a_{i}!=\emptyset$ for all $i$) that $|A_{n}\cap m|/m<\mathsf{d}(A_{n})$ for all $m\geq 1$, proving also condition (f). Now, note that for all nonzero $i\in n+2$ we have by (17) that
    $\begin{split}\ell_{i}&\leq\max H_{i}\leq\sum_{j=0}^{i}j\left\lceil\frac{\ell_{i}}{a_{j}!}\right\rceil\\ &\leq\sum_{j=0}^{i}j+\ell_{i}\sum_{j\in\omega}\frac{j}{a_{j}!}\leq i^{2}+\ell_{i}(1-\kappa),\end{split}$
    which implies that
    $\ell_{i}\leq\frac{i^{2}}{\kappa}\quad\text{ and }\quad\max H_{i}\leq i^{2}+\ell_{i}(1-\kappa)\leq\frac{i^{2}}{\kappa}.$
    Recalling that $C=\lceil 1/\kappa\rceil$, it follows that
    $\forall i\in n+2,\quad B_{i}\subseteq a_{i}!\cdot\omega+\{0,1,\ldots,Ci^{2}\}.$ (19)
    In addition, since $a_{0}!>(N+1)^{2}$ by (17), it follows by the construction of the sets $H_{i}$ that
    $\forall i\in N+2,\quad B_{i}=a_{i}!\cdot(\omega\setminus 0)+\left\{\binom{i}{2}+j:j\in i\right\}.$
    The latter identity implies that, if $n\in\{1,\ldots,N\}$, then $A_{n-1}\subseteq a_{0}!\cdot\omega+[0,1,\ldots,\max H_{n}]$ and $(a_{n+1}!+[\min H_{n+1},\max H_{n+1}])\subseteq a_{0}!\cdot\omega+[\min H_{n+1},\max H_{n+1}]$. Considering also that $\max H_{n}<\min H_{n+1}$ by construction, we get
    $\forall n\in\{1,\ldots,N\},\quad A_{n-1}\cap(a_{n+1}!+[\min H_{n+1},\max H_{n+1}])=\emptyset.$
    In particular, if $n\in\{1,\ldots,N\}$, then $\ell_{n+1}=n+1$. Suppose now that $n\geq N+1$ and note by (16) that $\log^{2}(n)/n\leq\log^{2}(N)/N<(1-\kappa)/2$. Observe that, if $i>\log(n)$ then $i>\log((n+1)/e)$, which is equivalent to $n+1<e^{i+1}$. Hence it follows by (17) and (19) that, if $i>\log(n)$, then
    $\max H_{n+1}\leq C(n+1)^{2}<Ce^{2(i+1)}\leq a_{i}!.$
    This implies that
    $\begin{split}|A_{n-1}\cap(a_{n+1}!&+[\min H_{n+1},\max H_{n+1}])|\\ &\leq\sum_{i=0}^{n}|B_{i}\cap(a_{n+1}!+[\min H_{n+1},\max H_{n+1}])|\\ &\leq\ell_{n+1}\left(\sum_{i=0}^{n}\frac{i}{a_{i}!}\right)+\sum_{i=0}^{\lfloor\log(n)\rfloor}i\\ &\leq\ell_{n+1}\left(\sum_{i\in\omega}\frac{i}{a_{i}!}+\frac{\log^{2}(n)}{\ell_{n+1}}\right)\\ &\leq\ell_{n+1}\left(\frac{1-\kappa}{2}+\frac{\log^{2}(n)}{n}\right)\\ &\leq\ell_{n+1}(1-\kappa).\end{split}$
    Hence, in both cases (i.e., if $n\leq N$ or if $n\geq N+1$), we get that $|A_{n-1}\cap(a_{n+1}!+[\min H_{n+1},\max H_{n+1}])|\leq\ell_{n+1}(1-\kappa)$. Therefore
    $\begin{split}n+1&=|H_{n+1}|=|B_{n+1}\cap(a_{n+1}!+[\min H_{n+1},\max H_{n+1}])|\\ &=\ell_{n+1}-|A_{n-1}\cap(a_{n+1}!+[\min H_{n+1},\max H_{n+1}])|\geq\kappa\ell_{n+1}.\end{split}$
    This proves condition (c) and completes the induction.
To complete the proof, we claim that the sequence of sets $(A_{n}:n\in\omega)$ in the above construction satisfies our claim. For, we have by construction that $A_{n+1}\setminus A_{n}=B_{n+2}\in\mathscr{A}$ for all $n\in\omega$. Since $\mu^{\star}(B_{n})=|H_{n}|/a_{n}!$ by [25, Proposition 7], it follows by (17) and the conditions in item (v) that
$\sum_{n\in\omega}\mu^{\star}(A_{n+1}\setminus A_{n})\leq\sum_{n\in\omega}\mu^{\star}(B_{n})=\sum_{n\in\omega}\frac{n}{a_{n}!}<\infty.$
Now, pick a set $A\subseteq\omega$ such that $\mu^{\star}(A_{n}\setminus A)=0$ for all $n\in\omega$. Define also
$\forall n\in\omega,\quad C_{n}:=B_{n}\setminus A\quad\text{ and }\quad D_{n}:=B_{n}\setminus C_{n},$
so that $\{C_{n},D_{n}\}$ is a partition of $B_{n}$, and $D_{n}\subseteq A$ for all $n\in\omega$. In addition, we get by monotonicity of $\mu^{\star}$ that $\mu^{\star}(C_{n})\leq\mu^{\star}(A_{n}\setminus A)=0$ for all $n\in\omega$, hence $\mu^{\star}(C_{n})=0$. Thanks to (5), we get $\mathsf{d}^{\star}(C_{n})=0$, i.e.,
$\forall n\in\omega,\forall\varepsilon>0,\exists N_{n,\varepsilon}\geq 1,\forall m\geq N_{i,\varepsilon},\quad\frac{|C_{n}\cap m|}{m}<\varepsilon.$ (20)
It follows that for any positive integer $n$ and sufficiently large $m$ we have $|C_{n}\cap m|/m<1/a_{n}!$. This implies there exists $j_{n}\in\omega$ such that
$C_{n}\cap(a_{n}!j_{n}+[\min H_{n},\max H_{n}])=\emptyset.$
Hence, by the condition (c) in item (iii), we get
$\frac{|D_{n}\cap(a_{n}!j_{n}+[\min H_{n},\max H_{n}])|}{\ell_{n}}=\frac{|H_{n}|}{\ell_{n}}\geq\kappa.$
It follows by the definition of the upper Banach density that
$\forall n\geq 1,\quad\mathrm{bd}^{\star}\left(\bigcup_{i\geq n}D_{i}\right)\geq\kappa.$ (21)
Therefore, for each $n\in\omega$, we get
$\mu^{\star}(A\setminus A_{n})\geq\mu^{\star}\left(\bigcup_{i\geq n+2}D_{i}\right)\geq\mathrm{bd}^{\star}\left(\bigcup_{i\geq n+2}D_{i}\right)\geq\kappa,$
which completes the proof. ∎
Remark 3.7.
The argument provided in the proof of Theorem 2.5 shows that, if $\mathscr{A}_{\infty}$ stands for the family of countably infinite unions of infinite arithmetic progressions $k\cdot\omega+h$, then
$\mathscr{A}_{\infty}\setminus\mathsf{dom}(\mathsf{bd})\neq\emptyset.$
In fact, suppose that $\kappa=1/2$, so that $\sum_{n}n/a_{n}!\leq 1/4$ by (17). Then the set $B:=\bigcup_{n}B_{n}\in\mathscr{A}_{\infty}$ contains $\bigcup_{i\geq 1}D_{n}$, hence $\mathsf{bd}^{\star}(B)\geq 1/2$ by (21). On the other hand, for each $m\geq 1$, there exists $k_{m}\geq 1$ for which $B\cap m=\bigcup_{i\in k_{m}}B_{i}\cap m$, so that by the definition of the sets $B_{i}$ we get
$\frac{|B\cap m|}{m}\leq\sum_{i\in k_{m}}\frac{|B_{i}\cap m|}{m}\leq\sum_{i\in k_{m}}\frac{i}{a_{i}!}\leq\frac{1}{4}.$
This implies that $\mathsf{bd}_{\star}(B)\leq\mathsf{d}_{\star}(B)\leq 1/4$, cf. (5). Therefore $B\notin\mathrm{dom}(\mathsf{bd})$.
This example also shows that the restriction of the upper Banach density $\mathsf{bd}^{\star}$ to $\mathcal{P}(\omega)\setminus\{A\subseteq\omega:\mathsf{bd}^{\star}(A)=0\}$ is not $\sigma$-subadditive. This comment applies to all upper densities $\mu^{\star}$ for which $\mu^{\star}(A)\geq\mathsf{bd}^{\star}(A)$ for all $A\subseteq\omega$.
3.1. Acknowledgments
The authors are grateful to K. P. S. Bhaskara Rao (Indiana University Northwest, US) and Ilijas Farah (York University, CA) for several helpful communications during the preparation of this manuscript.