### v2

# Completeness and additive property for submeasures

Abstract.
Given an extended real-valued submeasure $\nu$ defined on a field of subsets $\Sigma$ of a given set, we provide necessary and sufficient conditions for which the pseudometric $d_{\nu}$ defined by $d_{\nu}(A,B):=\min\{1,\nu(A\bigtriangleup B)\}$ for all $A,B\in\Sigma$ is complete. As an application, we show that if $\varphi:\mathcal{P}(\omega)\to[0,\infty]$ is a lower semicontinuous submeasure and $\nu(A):=\lim_{n}\varphi(A\setminus\{0,1,\ldots,n-1\})$ for all $A\subseteq\omega$, then $d_{\nu}$ is complete. This includes the case of all weighted upper densities, fixing a gap in a proof by Just and Krawczyk in [Trans. Amer. Math. Soc. 285 (1984), 803–816]. In contrast, we prove that if $\nu$ is the upper Banach density (or an upper density greater than or equal to the latter) then $d_{\nu}$ is not complete. We conclude with several characterizations of completeness in terms of the Stone space of the Boolean algebra $\Sigma/\nu$.
Key words and phrases:
Completeness; submeasures; additive property; lower semicontinuous submeasures; upper Banach density; exact Hahn decomposition; Stone space.
2010 Mathematics Subject Classification:
Primary: 11B05, 54E50. Secondary: 28A12, 54A20.
1. Introduction
Let $\Sigma\subseteq\mathcal{P}(X)$ be a field of subsets of a nonempty set $X$, that is, a nonempty family of subsets of $X$ that is stable under finite unions and complements. Denote also by $\omega$ the set of nonnegative integers and by $\overline{\mathbb{R}}:=[-\infty,\infty]$ the two-point compactification of the reals, endowed with its natural total order. A map $\nu:\Sigma\to\overline{\mathbb{R}}$ is said to be a capacity if it is monotone (that is, $\nu(A)\leq\nu(B)$ for all $A,B\in\Sigma$ with $A\subseteq B$) and satisfies $\nu(\emptyset)=0$. If, in addition, the capacity $\nu$ is also subadditive (that is, $\nu(A\cup B)\leq\nu(A)+\nu(B)$ for all $A,B\in\Sigma$), then it is said to be a submeasure. Lastly, a capacity (or a submeasure) $\nu$ is called bounded if $\nu(X)<\infty$ and normalized if $\nu(X)=1$. It is clear that submeasures are necessarily nonnegative. If there is no risk of confusion, we will write also $\nu:\Sigma\to S$, for some $S\subseteq\overline{\mathbb{R}}$. Given a submeasure $\nu:\Sigma\to\overline{\mathbb{R}}$, it is easy to check that the map $d_{\nu}:\Sigma^{2}\to\mathbb{R}$ defined by
$\forall A,B\in\Sigma,\quad d_{\nu}(A,B):=\min\{1,\nu(A\bigtriangleup B)\}.$
is a pseudometric on the field $\Sigma$. In the context of Boolean algebras, $d_{\nu}$ is also known as a Fréchet–Nikodym metric, see e.g. [13, Fact 2.3]. The aim of this work is to provide necessary and sufficient conditions on $\nu$ such that the pseudometric space $(\Sigma,d_{\nu})$ is complete, hence to show completeness or non-completeness for large classes of submeasures. It is worth noting that the completeness of the pseudometrics $d_{\nu}$ was important in studies of the famous Maharam’s problem [41, 47], cf. e.g. [3, p. 62] and [2, 31].
Examples of normalized submeasures include the nonnegative finitely additive probability measures $\mu:\Sigma\to\mathbb{R}$ (for instance, the characteristic functions of free ultrafilters on $\omega$ and, of course, all the $\sigma$-additive probability measures), the upper asymptotic density $\mathsf{d}^{\star}:\mathcal{P}(\omega)\to\mathbb{R}$ defined by
(1) $\forall A\subseteq\omega,\quad\mathsf{d}^{\star}(A):=\limsup_{n\to\infty}\frac{|A\cap n|}{n}$
(where, as usual, we identify each $n\in\omega$ with the set $\{0,1,\ldots,n-1\}$), and all upper densities on $\omega$ considered in [14, 24, 38], e.g. the upper analytic, upper Pólya, and upper Banach densities. In different contexts, we find also exhaustive submeasures [33, 47], nonpathological and lower semicontinuous submeasures [18, 45], and duals of exact games [42, 43].
$1.$ Literature results: additive setting
Throughout, we reserve the symbol $\mu$ for finitely additive maps. The following result characterizes the completeness of the pseudometric space $(\Sigma,d_{\mu})$ in the case where the domain $\Sigma$ is a $\sigma$-field.
Theorem $1.$
Let $\Sigma\subseteq\mathcal{P}(X)$ be a $\sigma$-field of subsets on a set $X$, and let $\mu:\Sigma\to\mathbb{R}$ be a nonnegative, finitely additive, bounded map. The following are equivalent:
  1. (i)
    The pseudometric space $(\Sigma,d_{\mu})$ is complete;
  2. (ii)
    For every increasing sequence $(A_{n}:n\in\omega)$ in $\Sigma$, there exists $A\in\Sigma$ such that $\mu(A_{n}\setminus A)=0$ for all $n\in\omega$ and $\mu(A)=\lim_{n}\mu(A_{n})$.
Proof.
See [28, Theorem 1] or [27, Lemma 2.2, Theorem 2.4, and Remark 2.5]. ∎
Additional equivalent conditions can be found in [28, Theorem 1], including completeness of the function space $L_{1}(\mu)$ or $L_{p}(\mu)$ for some $p\in[0,\infty)$ (see also [6, Theorem 3.4]). We refer the reader to the article [6] of Basile and Bhaskara Rao for a systematic study of completeness of $L_{p}(\mu)$ spaces and the related spaces $(\Sigma,d_{\mu})$. However, we will not consider these characterisations further here.
The property in item (ii) above is usually called AP(null) and provides a weak form of continuity from below for $\mu$, cf. e.g. [10]. The next result shows that, even in the case $\Sigma=\mathcal{P}(\omega)$, one cannot replace the condition $\mu(A_{n}\setminus A)=0$ for all $n\in\omega$ in the latter property with $|A_{n}\setminus A|<\infty$ for all $n\in\omega$. To this aim, recall that the asymptotic density $\mathsf{d}(A)$ of a set $A\subseteq\omega$ is defined as in (1) by replacing $\limsup_{n\to\infty}$ with $\lim_{n\to\infty}$, provided that latter limit exists (cf. Section 2.2).
Theorem 1.2.
There exists an additive extension $\mu:\mathcal{P}(\omega)\to[0,1]$ of the asymptotic density $\mathsf{d}$ with the following properties:
  1. (i)
    The pseudometric space $(\mathcal{P}(\omega),d_{\mu})$ is complete;
  2. (ii)
    There exists an increasing sequence $(A_{n}:n\in\omega)$ in $\mathcal{P}(\omega)$ such that, if $A\subseteq\omega$ satisfies $|A_{n}\setminus A|<\infty$ for all $n\in\omega$, then $\mu(A)\neq\lim_{n}\mu(A_{n})$.
Proof.
It follows by [10, Corollary 4 or Theorem 5(b)] and Theorem $1.$ ∎
In contrast, the same authors also proved the following:
Theorem 1.3.
There exists an additive extension $\mu:\mathcal{P}(\omega)\to[0,1]$ of the asymptotic density $\mathsf{d}$ such that the pseudometric space $(\mathcal{P}(\omega),d_{\mu})$ is not complete.
Proof.
It follows by [10, Theorem 6] and the argument following its proof, together with Theorem $1.$ ∎
In both results above from [10], the authors consider special additive extensions of the asymptotic density $\mathsf{d}$ which are also called density measures, namely, maps $\mu^{\mathscr{F}}:\mathcal{P}(\omega)\to[0,1]$ defined by
(2) $\forall A\subseteq\omega,\quad\mu^{\mathscr{F}}(A):=\mathscr{F}\text{-}\lim_{n\to\infty}\frac{|A\cap n|}{n},$
where $\mathscr{F}$ is a free ultrafilter on $\omega$. Kunisada proved in [35, Theorem 5.1] a necessary and sufficient condition for a density measure $\mu^{\mathscr{F}}$ to have AP(null).
1.2. Literature results: non-additive setting
So far, only some rather special cases have been proved in the literature, all in the case $\Sigma=\mathcal{P}(\omega)$.
The first example is due to Solecki in his seminal work [45]: a map $\varphi:\mathcal{P}(\omega)\to[0,\infty]$ is said to be a lower semicontinuous submeasure (in short, lscsm) if it is a submeasure such that $\varphi(F)<\infty$ for all finite $F\subseteq\omega$ and
$\forall A\subseteq\omega,\quad\varphi(A)=\sup\{\varphi(A\cap n):n\in\omega\}.$
Notice that the above property is precisely the lower semicontinuity of the submeasure $\varphi$, regarding its domain $\mathcal{P}(\omega)$ as the Cantor space $2^{\omega}$, that is, if $A_{n}\to A$ then $\liminf_{n}\varphi(A_{n})\geq\varphi(A)$. Examples of lscsms include $\varphi(A)=|A|$ or $\varphi(A)=\sum_{n\in A}1/(n+1)$ or $\varphi(A)=\sup_{n\geq 1}|A\cap n|/n$, cf. also [18, Chapter 1].
Theorem 1.4.
Let $\varphi:\mathcal{P}(\omega)\to[0,\infty]$ be a lscsm. Then the pseudometric space $(\mathcal{P}(\omega),d_{\varphi})$ is complete.
Proof.
See the proof of implication (iii) $\implies$ (i) in [45, Theorem 3.1]. ∎
Of course, the upper asymptotic density $\mathsf{d}^{\star}$ defined in (1) is not a lscsm (since $\mathsf{d}^{\star}(A)=0$ for every finite $A\subseteq\omega$). However, the first-named author recently proved that the analogous result holds for $\mathsf{d}^{\star}$:
Theorem 1.5.
The pseudometric space $(\mathcal{P}(\omega),d_{\mathsf{d}^{\star}})$ is complete.
Proof.
See [34, Theorem 1.1]. ∎
Just and Krawczyk claimed in [32, Lemma 3.1] that the analogue of Theorem 1.5 holds for all weighted upper densities, that is, maps $\nu_{f}:\mathcal{P}(\omega)\to\mathbb{R}$ defined by $\nu_{f}(A):=\limsup_{n}\sum_{i\in A\cap n}f(i)/\sum_{i\in n}f(i)$, where $f:\omega\to\mathbb{R}$ is an Erdős–Ulam function, that is, a nonnegative function such that $\lim_{n}\sum_{i\in n}f(i)=\infty$ and $\lim_{n}f(n)/\sum_{i\in n}f(i)=0$. It seems to us that their proof contains a gap that does not appear easily fixable. [^1] More precisely, the last centered upper bound of the distance $\rho_{h}(a_{n}/I_{h},a/I_{h})$ in the proof of [32, Lemma 3.1] is given by $2^{-n}+\sum_{k=1}^{\infty}2^{-(n+k)}$, which goes to 0 as $n\to\infty$. However, the latter should be replaced by $2^{-n}+k2^{-n}$, which does not necessarily go to 0 as $n\to\infty$ because in the construction there is no explicit dependence between $n$ and $k$. A further generalization was claimed by Farah in [18, Lemma 1.3.3(c)], which has been used also in [22, Proposition 2]. However, as confirmed by Farah in personal communications regarding the latter result, his argument also contains a gap. We show in Theorem 2.5 below, with a different proof, that their claims were, in fact, correct. (A revised and corrected version of [18, Lemma 1.3.3(c)] can be found in [17, Proposition 5.2.2].)
2. Main results
Our first main result characterizes the completeness of the pseudometric spaces $(\Sigma,d_{\nu})$, hence providing the analogue of Theorem 1.1 for submeasures.
To this aim, if $\Sigma$ is a field of subsets on a set $X$, we say then a $\Sigma$-partition $\langle A,A^{c}\rangle$ of $X$ is an exact Hahn decomposition of a map $\lambda:\Sigma\to\overline{\mathbb{R}}$ if $\lambda(B)\geq 0$ for all $B\in\Sigma$ with $B\subseteq A$, and $\lambda(B)\leq 0$ for all $B\in\Sigma$ with $B\subseteq A^{c}$. It is known that exact Hahn decompositions do not necessarily exist, even if $\lambda$ is finitely additive, see [9, Remark 2.6.3 and Example 11.4.7]. Characterizations of completeness in the finitely additive case through exact Hahn decompositions can be found in [1, Proposition 3] and [6, Proposition 6.2]. It is worth remarking that exact Hahn decompositions played a role to prove a Radon-Nikodym type theorem for certain submeasures, see [29].
Moreover, given a submeasure $\nu:\Sigma\to\overline{\mathbb{R}}$ and a $d_{\nu}$-Cauchy sequence $(A_{n}:n\in\omega)$ with values in $\Sigma$, we write
[…]
for all $B\in\Sigma$. Observe that both limits exist in $0,\infty]$: in fact, since $(A_{n}\cap B)\bigtriangleup(A_{m}\cap B)\subseteq A_{n}\bigtriangleup A_{m}$ for all $n,m\in\omega$ then the sequence $(\nu(A_{n}\cap B):n\in\omega)$ is Cauchy in $0,\infty]$, hence convergent to $\nu^{+}_{(A_{n})}(B)\in[0,\infty]$ (the case of $\nu^{-}_{(A_{n})}(B)$ is analogous). In the following, we assume the convention $\infty-\infty:=0$.
Theorem 2.1.
Let $\Sigma\subseteq\mathcal{P}(X)$ be a field of subsets on a set $X$, and let $\nu:\Sigma\to\overline{\mathbb{R}}$ be a submeasure. Then the following are equivalent:
  1. (i)
    The pseudometric space $(\Sigma,d_{\nu})$ is complete;
  2. (ii)
    For every $d_{\nu}$-Cauchy sequence $(A_{n}:n\in\omega)$ in $\Sigma$, the map $\nu^{+}_{(A_{n})}-\nu^{-}_{(A_{n})}$ admits an exact Hahn decomposition;
  3. (iii)
    For every increasing sequence $(A_{n}:n\in\omega)$ in $\Sigma$ such that $\sum_{n}\nu(A_{n+1}\setminus A_{n})<\infty$, there exists $A\in\Sigma$ such that $\nu(A_{n}\setminus A)=0$ for all $n\in\omega$ and $\lim_{n}\nu(A\setminus A_{n})=0$.
Condition (iii) above is the analogue of AP(null) for submeasures. Note that, even if $\nu(F)=0$ for all finite $F\in\Sigma$, the condition $\nu(A_{n}\setminus A)=0$ for all $n\in\omega$”  cannot be replaced with $|A_{n}\setminus A|<\infty$ for all $n\in\omega$,” (see Remark 3.1 below).
As a first immediate application, we show that Gangopadhyay’s characterization (namely, Theorem 1.1) holds, more generally, for fields of sets.
Corollary $2.$
Let $\Sigma\subseteq\mathcal{P}(X)$ be a field of subsets on a set $X$, and let $\mu:\Sigma\to\mathbb{R}$ be a nonnegative, finitely additive, bounded map. The following are equivalent:
  1. (i)
    The pseudometric space $(\Sigma,d_{\mu})$ is complete;
  2. (ii)
    For every increasing sequence $(A_{n}:n\in\omega)$ in $\Sigma$, there exists $A\in\Sigma$ such that $\mu(A_{n}\setminus A)=0$ for all $n\in\omega$ and $\mu(A)=\lim_{n}\mu(A_{n})$.
Quite surprisingly, it seems that the above characterization cannot be found explicitly in the literature. As remarked by K. P. S. Bhaskara Rao in personal communications, [8, Theorem 3.3] is a relevant result in the finitely additive setting and may provide an alternative route to a direct proof of Corollary $2.$
As another immediate application, we recover [34, Proposition 3.4]:
Corollary 2.3.
Let $\Sigma\subseteq\mathcal{P}(X)$ be a $\sigma$-field of subsets on a set $X$, and let $\nu:X\to\overline{\mathbb{R}}$ be a submeasure which is $\sigma$-subadditive. Then $(\Sigma,d_{\nu})$ is complete.
In Section 2.1, we provide a large class of submeasures $\nu:\mathcal{P}(\omega)\to[0,\infty]$ such that $(\mathcal{P}(\omega),d_{\nu})$ is complete (see Theorem 2.5), which will give a positive proof to the claims in [32, Lemma 3.1] and [18, Lemma 1.3.3(c)]. In Section 2.2, we provide another large class of submeasures $\nu:\mathcal{P}(\omega)\to[0,\infty]$ such that $(\mathcal{P}(\omega),d_{\nu})$ is not complete (see Theorem 2.7). Remarkably, this class includes the upper Banach density $\mathsf{bd}^{\star}$ and the upper Buck density $\mathfrak{b}^{\star}$ defined in (3) and (4) below, resp.; in both cases, some preliminaries will be necessary. Lastly, we will provide in Section 2.3 further relationships between the completeness of the pseudometric spaces $(\Sigma,d_{\nu})$ and a certain submeasure defined on the power set of the Stone space of the quotient $\Sigma/\nu$. Proofs of our results follow in Section 3.
2.1. Positive results: ideals and lscsms
Let $\mathcal{I}$ be an ideal on the nonnegative integers $\omega$, that is, a family of subsets of $\omega$ which is closed under subsets and finite unions. Unless otherwise stated, it is also assumed that $\mathcal{I}$ contains the family $\mathrm{Fin}$ of finite subsets of $\omega$, and that $\omega\notin\mathcal{I}$. Set also $\mathcal{I}^{+}:=\mathcal{P}(\omega)\setminus\mathcal{I}$.
Proposition 2.4.
Let $\mathcal{I}$ be an ideal on $\omega$. Then $(\mathcal{P}(\omega),d_{\bm{1}_{\mathcal{I}^{+}}})$ is complete.
Regarding ideals as subsets of the Cantor space $2^{\omega}$, we can speak about their topological complexity: for instance, $\mathcal{I}$ can be an analytic, Borel, or a $F_{\sigma}$-subset of $\mathcal{P}(\omega)$. In addition, we say that $\mathcal{I}$ is a $P$-ideal if it is $\sigma$-directed modulo finite sets, that is, for every sequence $(A_{n})$ of sets in $\mathcal{I}$ there exists $A\in\mathcal{I}$ such that $A_{n}\setminus A$ is finite for all $n\in\omega$. For instance, $\mathrm{Fin}$ is a $F_{\sigma}$ $P$-ideal, and the family of asymptotic density zero sets
$\mathcal{Z}:=\left\{A\subseteq\omega:\mathsf{d}^{\star}(A)=0\right\}$
is an analytic $P$-ideal which is not $F_{\sigma}$. Also, maximal ideals (i.e., the complements of free ultrafilters on $\omega$) do not have the Baire property, hence they are not analytic. We refer to [18] for an excellent textbook on the theory of ideals.
Given a lscsm $\varphi:\mathcal{P}(\omega)\to[0,\infty]$, define the family
$\mathrm{Exh}(\varphi):=\{S\subseteq\omega:\|S\|_{\varphi}=0\},\,\,\,\,\text{ where }\,\,\|S\|_{\varphi}:=\lim_{n\to\infty}\varphi(S\setminus n).$
Informally, $\|S\|_{\varphi}$ stands for the $\varphi$-mass at infinity of the set $S$. A classical result of Solecki [45, Theorem 3.1] states that an ideal $\mathcal{I}$ on $\omega$ is an analytic $P$-ideal if and only if there exists a lscsm $\varphi$ such that
$\mathcal{I}=\mathrm{Exh}(\varphi)\quad\text{ and }\quad\varphi(\omega)<\infty.$
We remark that the family of analytic $P$-ideals is large and includes, among others, all Erdős–Ulam ideals introduced by Just and Krawczyk in [32], ideals generated by nonnegative regular matrices [23, 25], the Fubini products $\emptyset\times\mathrm{Fin}$, which can be defined as $\{A\subseteq\omega:\forall n\in\omega,A\cap I_{n}\in\mathrm{Fin}\},$ where $(I_{n})$ is a given partition of $\omega$ into infinite sets, certain ideals used by Louveau and Velic̆ković [40], and, more generally, density-like ideals and generalized density ideals [12, 36]. Additional pathological examples can be found in [46]. It has been suggested in [11, 12] that the theory of analytic $P$-ideals may have some relevant yet unexploited potential for the study of the geometry of Banach spaces.
It is easy to check that, for each finite lscsm $\varphi$, the map $\|\cdot\|_{\varphi}$ is monotone, subadditive, and invariant under finite modifications, cf. [18, Lemma 1.3.3(a)-(b)]. In addition, the lscsm $\varphi$ in such a representation is not necessarily unique, cf. Example 3.3 below for the case of the ideal $\mathcal{Z}$. On a similar note, we will show in Proposition 3.5 that, if $\varphi_{1},\varphi_{2}$ are two lscsms with $\mathrm{Exh}(\varphi_{1})=\mathrm{Exh}(\varphi_{2})$, then the pseudometrics $d_{\|\cdot\|_{\varphi_{1}}}$ and $d_{\|\cdot\|_{\varphi_{2}}}$ are topologically equivalent, though not necessarily metrically equivalent (see Proposition 3.4).
With the above context, our main positive result follows:
Theorem 2.5.
Let $\varphi:\mathcal{P}(\omega)\to[0,\infty]$ be a lscsm. Then $(\mathcal{P}(\omega),d_{\|\cdot\|_{\varphi}})$ is complete.
Since $\mathcal{Z}$ is an analytic $P$-ideal, Theorem 2.5 provides a generalization of Theorem 1.5, cf. also Example 3.3 below. In addition, a special case of Theorem 2.5 follows by a result of Solecki in [45, Theorem 3.4], where he shows that the topology induced by $d_{\|\cdot\|_{\varphi}}$ is discrete if and only if $\mathrm{Exh}(\varphi)$ is a $F_{\sigma}$-ideal. It seems plausible that an alternative proof of Theorem 2.5 could be obtained using the notion of algebraic convergence studied in [3, Section 2], cf. also [2, 31].
As remarked by Farah in personal communications, an instance of Theorem 2.5 can also be obtained as follows: suppose that $\varphi$ is a lscsm such that $\mathrm{Exh}(\varphi)$ is a generalized density ideal. Then the Boolean algebra $\mathcal{P}(\omega)/\mathrm{Exh}(\varphi)$ endowed with the metric induced by $d_{\|\cdot\|_{\varphi}}$ is countably saturated, see [21, Proposition 2.8] and cf. also [19, Theorem 16.5.1] and the argument preceding [20, Corollary 6.5]; thus, completeness is a consequence of the latter property. However, it is an open question whether the analogous argument holds for all analytic $P$-ideals; see the last paragraph of [21, Section 2].
The density measures $\mu^{\mathscr{F}}$ in (2) used by Blass et al. [10] do not fall into the realm of Theorem 2.5: in fact, for each free ultrafilter $\mathscr{F}$ on $\omega$, the family $\{A\subseteq\omega:\mu^{\mathscr{F}}(A)=0\}$ is an ideal which is not analytic by the proof of [5, Theorem 2.10], hence it cannot coincide with $\mathrm{Exh}(\varphi)$ for any lscsm $\varphi$.
$2.$ Negative results: upper densities
In this section, complementing Proposition 2.4, we provide a simple class of pseudometric spaces $(\mathcal{P}(\omega),d_{\nu})$ which are not complete.
Proposition 2.6.
Let $\mathcal{I}$ be an ideal on $\omega$. Let also $(a_{n}:n\in\omega)$ be a sequence of strictly positive reals such that $\sum_{n}a_{n}=1$. Define the submeasure $\nu$ by
$\forall A\subseteq\omega,\quad\nu(A):=\bm{1}_{\mathcal{I}^{+}}(A)+\sum_{n\in A}a_{n}.$
Then $(\mathcal{P}(\omega),d_{\nu})$ is not complete.
Then, we study natural instances of $\nu$ such as upper Banach density. To this aim, following [38, Definition 1], we say that a map $\mu^{\star}:\mathcal{P}(\omega)\to\mathbb{R}$ is an upper density on $\omega$ if it satisfies the following properties:
  1. (f1)
    $\mu^{\star}(\omega)=1$;
  2. (f2)
    $\mu^{\star}(A)\leq\mu^{\star}(B)$ for all $A\subseteq B\subseteq\omega$;
  3. (f3)
    $\mu^{\star}(A\cup B)\leq\mu^{\star}(A)+\mu^{\star}(B)$ for all $A,B\subseteq\omega$;
  4. (f4)
    $\mu^{\star}(k\cdot A)=\mu^{\star}(A)/k$ for all $A\subseteq\omega$ and all nonzero $k\in\omega$;
  5. (f5)
    $\mu^{\star}(A+h)=\mu^{\star}(A)$ for all $A\subseteq\omega$ and $h\in\omega$.
It is readily seen that, if $\mu^{\star}$ is an upper density, then $0\leq\mu^{\star}(A)\leq 1$ for all $A\subseteq\omega$, and $\mu^{\star}(F)=0$ for every finite $F\subseteq\omega$, see [38, Proposition 2 and Proposition 6]; in particular, each $\mu^{\star}$ is a normalized submeasure.
Examples of upper densities include the upper asymptotic density $\mathsf{d}^{\star}$ defined in (1), the upper $\alpha$-densities with $\alpha\geq-1$ [38, Example 4], the upper Banach density $\mathsf{bd}^{\star}$ defined by
(3) $\forall A\subseteq\omega,\quad\mathsf{bd}^{\star}(A):=\lim_{n\to\infty}\,\max_{k\in\omega}\,\frac{|A\cap[k,k+n)|}{n},$
the upper analytic density [38, Example 6], the upper Polya density [38, Example 8], and the upper Buck density $\mathfrak{b}^{\star}$ defined by
(4) $\forall A\subseteq\omega,\quad\mathfrak{b}^{\star}(A):=\inf_{X\in\mathscr{A},A\subseteq X}\,\mathsf{d}^{\star}(A),$
where $\mathscr{A}$ stands for the family of finite unions of infinite arithmetic progressions $k\cdot\omega+h$ (with $k,h\in\omega$ and $k\neq 0$). With these premises, recall that
(5) $\mathsf{d}^{\star}(A)\leq\mathsf{bd}^{\star}(A)\quad\text{ and }\quad\mu^{\star}(A)\leq\mathfrak{b}^{\star}(A)$
for all $A\subseteq\omega$ and all upper densities $\mu^{\star}$ on $\omega$, see [38, Theorem 3].
For any upper density $\mu^{\star}$ on $\omega$, we denote by $\mu_{\star}$ its lower dual, that is, the map defined by $\mu_{\star}(A):=1-\mu^{\star}(\omega\setminus A)$ for all $A\subseteq\omega$. We denote by $\mu$ the density induced by the pair $(\mu^{\star},\mu_{\star})$, that is, the restriction of $\mu^{\star}$ to the family $\mathrm{dom}(\mu):=\{A\subseteq\omega:\mu^{\star}(A)=\mu_{\star}(A)\}.$ For instance, if $\mu^{\star}$ is the upper asymptotic density $\mathsf{d}^{\star}$, then $\mathsf{d}_{\star}$ is the lower asymptotic density, $\mathsf{d}$ is the asymptotic density, and $\mathrm{dom}(\mathsf{d})$ is the family of sets $A\subseteq\omega$ which admits asymptotic density (that is, $\mathsf{d}^{\star}(A)=\mathsf{d}_{\star}(A)$).
Our main negative result follows:
Theorem 2.7.
Let $\mu^{\star}$ be an upper density on $\omega$ such that $\mathsf{bd}^{\star}(A)\leq\mu^{\star}(A)$ for all $A\subseteq\omega$. Then $(\mathcal{P}(\omega),d_{\mu^{\star}})$ is not complete.
Thanks to (5), the above result applies to the upper Banach density $\mathsf{bd}^{\star}$ and the upper Buck density $\mathfrak{b}^{\star}$. In addition, thanks to [38, Proposition 10], it applies also to the upper density $(\alpha(\textsf{bd}^{\star})^{q}+(1-\alpha)(\mathfrak{b}^{\star})^{q})^{1/q}$ for every $q\in[1,\infty)$ and every $\alpha\in[0,1]$ (more generally, the set of upper densities satisfying the hypothesis of Theorem 2.7 is countably $q$-convex for every $q\in[1,\infty))$.
2.3. Relationships with Stone spaces
Let $\Sigma$ and $\nu$ be as in the statement of Theorem 2.1, and let $\nu^{\star}:\mathcal{P}(X)\to\overline{\mathbb{R}}$ be the map defined by
$\forall A\subseteq X,\quad\nu^{\star}(A):=\inf\left\{\nu(C):C\in\Sigma\text{ and }A\subseteq C\right\}.$
Of course, $\nu^{\star}$ is a submeasure which extends $\nu$. We write $\overline{\Sigma}$ for the closure of $\Sigma$ in the space $(\mathcal{P}(X),d_{\nu^{\star}})$. As it follows by Remark 3.8, this is precisely the classical Peano–Jordan completion in the finitely additive case. With a small abuse of notation, given a field $\Sigma^{\prime}\subseteq\mathcal{P}(X)$, we write $(\Sigma^{\prime},d_{\nu^{\star}})$ for the pseudometric space $(\Sigma^{\prime},d_{\psi})$, where $\psi$ stands for the restriction of $\nu^{\star}$ on $\Sigma^{\prime}$.
Lemma 2.8.
Let $\Sigma\subseteq\mathcal{P}(X)$ be a field of subsets on a set $X$, and let $\nu:\Sigma\to\overline{\mathbb{R}}$ be a submeasure. Then $(\overline{\Sigma},d_{\nu^{\star}})$ is complete whenever $(\Sigma,d_{\nu})$ is complete.
It is worth noting that, as it follows by [6, Proposition 5.1 and Example 5.2], the converse implication of Lemma 2.8 fails even if $\nu$ is finitely additive.
Several equivalent conditions for the completeness of $(\overline{\Sigma},d_{\nu^{\star}})$ will be given in Theorem 2.9 below. To this aim, denote by $\Sigma/\nu$ the quotient space of $\Sigma$ with the ideal $\{A\in\Sigma:\nu(A)=0\}$. Equip $\Sigma/\nu$ with the quotient topology, and denote the canonical projection by
$\pi:\Sigma\to\Sigma/\nu.$
Thus, a generic element of $\Sigma/\nu$ is an equivalence class of the type $\pi(A)=\{C\in\Sigma:\nu(A\bigtriangleup C)=0\}$ for some set $A\in\Sigma$. Now, let $S$ be the Stone space of the Boolean algebra $\Sigma/\nu$, that is, the set of ultrafilters on $\Sigma/\nu$, equipped with its usual topology. Recall that $S$ is compact and totally disconnected. Denote by $\mathcal{C}$ the field of all clopen (that is, simultaneously open and closed) subsets of $S$, so that $\mathcal{C}\subseteq\mathcal{P}(S)$. We denote by $\sigma(\mathcal{C})$ the $\sigma$-field generated by $\mathcal{C}$. It is known that $\sigma(\mathcal{C})$ coincides with the Baire $\sigma$-field of $S$, that is, the smallest $\sigma$-field containing the compact $G_{\delta}$ sets (the short argument is essentially contained in Lemma 3.14(vi) below. By Stone’s representation theorem, there exists a Boolean isomorphism
$\phi:\Sigma/\nu\to\mathcal{C},$
see e.g. [44, Chapter 1, Section 8] for a textbook exposition on Boolean algebras. For each $B\subseteq S$ let $\Delta(B)$ be the family of all sequences $(A_{n}:n\in\omega)$ with values in $\Sigma$ such that $B\subseteq\bigcup_{n}\phi(\pi(A_{n}))$ (of course, each $\Delta(B)$ is nonempty since $\phi(\pi(X))=S$). Thus, define the outer measure $\hat{\nu}:\mathcal{P}(S)\to\overline{\mathbb{R}}$ by
$\forall B\subseteq S,\quad\hat{\nu}(B):=\inf\left\{\sum_{n\in\omega}\nu(A_{n}):(A_{n})\in\Delta(B)\right\},$
and observe that $\hat{\nu}$ is $\sigma$-subadditive by the very same argument used in the first part of the proof of [15, Lemma III.5.5]. Define also the map $\tilde{\nu}:\mathcal{P}(S)\to\overline{\mathbb{R}}$ by
$\forall B\subseteq S,\quad\tilde{\nu}(B):=\inf\left\{\nu(A):A\in\Sigma\text{ and }B\subseteq\phi(\pi(A))\right\}.$
Of course, both $\tilde{\nu}$ and $\tilde{\nu}$ are submeasures, and $\hat{\nu}\leq\tilde{\nu}$. Let also $\partial B$ be the set of boundary points of $B\subseteq S$. Finally, let $\mathrm{Cl}_{\hat{\nu}}(\mathcal{C})$ be the closure of $\mathcal{C}$ with respect to $d_{\hat{\nu}}$, and $\mathcal{N}_{\hat{\nu}}$ denote the family of $\hat{\nu}-$null sets. Several properties of $\mathrm{Cl}_{\hat{\nu}}(\mathcal{C})$ can be found in Lemma 3.9 and the subsequent results. Analogous notations are used for $\tilde{\nu}$. Lastly, let $\mathrm{nwd}(S)$ denote the family of the nowhere dense sets in $S$.
With the above premises, we have the following characterization, in the same spirit of [6, Theorem 4.2] for the finitely additive case, cf. also [7, 30] and the related comments in [6, Section 4].
Theorem 2.9.
Let $\Sigma\subseteq\mathcal{P}(X)$ be a field of subsets on a set $X$, and let $\nu:\Sigma\to\overline{\mathbb{R}}$ be a submeasure. Then the following are equivalent:
  1. (a1)
    The pseudometric space $(\overline{\Sigma},d_{\nu^{\star}})$ is complete;
  2. (a2)
    For every increasing sequence $(A_{n}:n\in\omega)$ in $\Sigma$ such that $\sum_{n}\nu(A_{n+1}\setminus A_{n})<\infty$, there exists a decreasing sequence $(D_{m}:m\in\omega)$ in $\Sigma$ such that $\nu(A_{n}\setminus D_{m})=0$ for all $n,m\in\omega$ and $\lim_{n}\nu(D_{n}\setminus A_{n})=0$;
  3. (a3)
    For every increasing sequence $(A_{n}:n\in\omega)$ in $\Sigma$ such that $\sum_{n}\nu(A_{n+1}\setminus A_{n})<\infty$, there exists $C\subseteq X$ such that $\nu^{\star}(A_{n}\setminus C)=0$ for all $n\in\omega$ and $\lim_{n}\nu^{\star}(C\setminus A_{n})=0$;
  1. (b1)
    $\hat{\nu}(U)=\hat{\nu}(\overline{U})$ for every set $U\subseteq S$;
  2. (b2)
    $\hat{\nu}(U)=\hat{\nu}(\overline{U})$ for every set $U\subseteq S$ which is a countable union of clopen sets;
  3. (b3)
    For every increasing sequence $(A_{n}:n\in\omega)$ in $\Sigma$ such that $\sum_{n}\nu(A_{n+1}\setminus A_{n})<\infty$, there exists a decreasing sequence $(D_{m}:m\in\omega)$ in $\Sigma$ such that $\nu(A_{n}\setminus D_{m})=0$ for all $n,m\in\omega$ and $\sup_{n}\nu(A_{n})=\inf_{m}\nu(D_{m})$;
  4. (b4)
    For every increasing sequence $(A_{n}:n\in\omega)$ in $\Sigma$ such that $\sum_{n}\nu(A_{n+1}\setminus A_{n})<\infty$, there exists $C\subseteq X$ such that $\nu^{\star}(A_{n}\setminus C)=0$ for all $n\in\omega$ and $\sup_{n}\nu(A_{n})=\nu^{\star}(C)$;
  1. (c1)
    The pseudometric space $(\mathcal{P}(X),d_{\nu^{\star}})$ is complete;
  2. (c2)
    The pseudometric space $(\Sigma^{\prime},d_{\nu^{\star}})$ is complete for every closed field $\Sigma^{\prime}\subseteq\mathcal{P}(X)$ which contains $\Sigma$;
  3. (c3)
    The pseudometric space $(\Sigma^{\prime},d_{\nu^{\star}})$ is complete for some closed field $\Sigma^{\prime}\subseteq\mathcal{P}(X)$ which contains $\Sigma$;
  1. (d1)
    $\hat{\nu}=\tilde{\nu}$;
  2. (d2)
    $\tilde{\nu}$ is $\sigma$-subadditive;
  3. (d3)
    For every $B\in\mathrm{Cl}_{\hat{\nu}}(\mathcal{C})$ and $\varepsilon>0$, there exist $A,C\in\mathcal{C}$ such that $A\subseteq B\subseteq C$ and $\tilde{\nu}(C\setminus A)<\varepsilon$;
  4. (d4)
    For every $B\in\mathrm{Cl}_{\hat{\nu}}(\mathcal{C})$, there exist sets $A,C\subseteq S$ such that AA is open, $C$ is closed, $A\subseteq B\subseteq C$, and $\hat{\nu}(C\setminus A)=0$;
  5. (d5)
    $\mathrm{Cl}_{\hat{\nu}}(\mathcal{C})=\mathrm{Cl}_{\tilde{\nu}}(\mathcal{C})$;
  6. (d6)
    A subset of $S$ is the boundary of some element of $\mathrm{Cl}_{\hat{\nu}}(\mathcal{C})$ if and only if it is closed and $\hat{\nu}-$null;
  7. (d7)
    $\mathcal{N}_{\hat{\nu}}=\mathrm{nwd}(S)\cap\mathrm{Cl}_{\hat{\nu}}(\mathcal{C})$ and $\partial B\in\mathrm{Cl}_{\hat{\nu}}(\mathcal{C})$ for every $B\in\mathrm{Cl}_{\hat{\nu}}(\mathcal{C})$;
  1. (e1)
    $\tilde{\nu}(\partial B)=0$ for every set $B\in\mathrm{Cl}_{\hat{\nu}}(\mathcal{C})$;
  2. (e2)
    $\hat{\nu}(\partial B)=0$ for every set $B\in\mathrm{Cl}_{\hat{\nu}}(\mathcal{C})$;
  3. (e3)
    $\hat{\nu}(\partial B)=0$ for every set $B\in\mathrm{Cl}_{\hat{\nu}}(\mathcal{C})$ which is a countable union of clopen sets.
As an application in the case where $\Sigma$ is closed, we provide some additional equivalences for the completeness of $(\Sigma,d_{\nu})$ in terms of properties of the Stone space $S$.
Corollary 2.10.
Let $\Sigma\subseteq\mathcal{P}(X)$ be a field of subsets on a set $X$ such that $\Sigma=\overline{\Sigma}$, and let $\nu:\Sigma\to\overline{\mathbb{R}}$ be a submeasure. Then the following are equivalent:
  1. (i)
    The pseudometric space $(\Sigma,d_{\nu})$ is complete;
  2. (ii)
    For every $B\in\mathrm{Cl}_{\hat{\nu}}(\mathcal{C})$ there exists $A\in\Sigma$ such that $\tilde{\nu}(\phi(\pi(A))\bigtriangleup B)=0$;
  3. (iii)
    For every $B\in\mathrm{Cl}_{\hat{\nu}}(\mathcal{C})$ there exists $A\in\Sigma$ such that $\hat{\nu}(\phi(\pi(A))\bigtriangleup B)=0$;
  4. (iv)
    For every $B\in\mathrm{Cl}_{\hat{\nu}}(\mathcal{C})$ which is a countable union of clopen sets, there exists $A\in\Sigma$ such that $\hat{\nu}(\phi(\pi(A))\bigtriangleup B)=0$.
To conclude, we remark that, in the case where $\nu$ is finitely additive, an equivalence in the same spirit of (a1) $\Longleftrightarrow$ (b1) can be obtained as a consequence of [6, Theorem 4.2 and Proposition 5.1] (however, in the latter work the definition of $\hat{\nu}$ is different, see [6, Section 1.15] for details).
3. Proofs of the main results in Section 2
Let us start with the proof of our main characterization.
Proof of Theorem 2.1.
(i) $\implies$ (ii). Let $(A_{n}:n\in\omega)$ be a $d_{\nu}$-Cauchy sequence in $\Sigma$ and let $A\in\Sigma$ be a $d_{\nu}$-limit. Then, for every $B\in\Sigma$, since $(A_{n}\cap B)\triangle(A\cap B)\subseteq A_{n}\triangle A$ for all $n\in\omega$, we have
$\nu^{+}_{(A_{n})}(B)=\lim_{n\to\infty}\nu(A_{n}\cap B)=\nu(A\cap B),$
and similarly $\nu^{-}_{(A_{n})}(B)=\nu(B\setminus A)$. Now, define $\lambda:=\nu^{+}_{(A_{n})}-\nu^{-}_{(A_{n})}$. It follows that $\lambda_{(A_{n})}(B)\geq 0$ for all $B\in\Sigma$ with $B\subseteq A$, and $\lambda_{(A_{n})}(B)\leq 0$ for all $B\in\Sigma$ with $B\subseteq A^{c}$. Hence $\langle A,A^{c}\rangle$ is an exact Hahn decomposition of $\lambda_{(A_{n})}$.
(ii) $\implies$ (iii). Pick an increasing sequence $(A_{n}:n\in\omega)$ in $\Sigma$ such that $\sum_{n}\nu(A_{n+1}\setminus A_{n})<\infty$. Observe that, for all $n,m\in\omega$ with $m\geq n$, we have $d_{\nu}(A_{n},A_{m})\leq\nu(A_{m}\setminus A_{n})\leq\sum_{k\geq n}\nu(A_{k+1}\setminus A_{k})$, hence $(A_{n}:n\in\omega)$ is $d_{\nu}$-Cauchy. It follows by item (ii) that the map $\lambda:=\nu^{+}_{(A_{n})}-\nu^{-}_{(A_{n})}$ admits an exact Hahn decomposition $\langle A,A^{c}\rangle$, for some $A\in\Sigma$. We claim that AA is a set witnessing (iii).
To this aim, fix $k\in\omega$. Since $A_{k}\setminus A\subseteq A^{c}$, then $\lambda(A_{k}\setminus A)\leq 0$. But $\nu^{-}_{(A_{n})}(A_{k}\setminus A)=\lim_{n}\nu((A_{k}\setminus A)\setminus A_{n})=0$. It follows that $\nu^{+}_{(A_{n})}(A_{k}\setminus A)=\lambda(A_{k}\setminus A)\leq 0$, hence $0=\nu^{+}_{(A_{n})}(A_{k}\setminus A)=\lim_{n}\nu(A_{n}\cap(A_{k}\setminus A))=\nu(A_{k}\setminus A)$.
Similarly, $A\setminus A_{k}\subseteq A$, hence $\lambda(A\setminus A_{k})\geq 0$, so $\nu^{+}_{(A_{n})}(A\setminus A_{k})\geq\nu^{-}_{(A_{n})}(A\setminus A_{k})$. Since $(A_{n}:n\in\omega)$ is increasing, we have also
$\nu^{-}_{(A_{n})}(A\setminus A_{k})=\lim_{n\to\infty}\nu((A\setminus A_{k})\setminus A_{n})=\lim_{n\to\infty}\nu(A\setminus A_{n}).$
Putting everything together, for all $k\in\omega$ it follows that
$\begin{split}\lim_{n\to\infty}\nu(A\setminus A_{n})&=\nu^{-}_{(A_{n})}(A\setminus A_{k})\leq\nu^{+}_{(A_{n})}(A\setminus A_{k})\\ &=\lim_{n\to\infty}\nu(A_{n}\cap(A\setminus A_{k}))\leq\lim_{n\to\infty}\nu(A_{n}\setminus A_{k})\\ &\leq\lim_{n\to\infty}\sum_{i=k}^{n-1}\nu(A_{i+1}\setminus A_{i})=\sum_{i\geq k}\nu(A_{i+1}\setminus A_{i}).\end{split}$
By the arbitrariness of $k$, we conclude that $\lim_{n}\nu(A\setminus A_{n})=0$.
(iii) $\implies$ (i). Let $(A_{n}:n\in\omega)$ be a Cauchy sequence in the pseudometric space $(\Sigma,d_{\nu})$. Without loss of generality, we can assume that $\nu(A_{i}\triangle A_{j})<2^{-i}$ for all $i,j\in\omega$ with $i\leq j$, since there is a subsequence with this property and a Cauchy sequence converges to the same limits as any subsequence. For all $i,j\in\omega$ with $i\leq j$, define the $\Sigma$-measurable set
$B_{i,j}:=\bigcap_{k=i}^{j}A_{k}$
and note that
(6) $\nu(A_{i}\triangle B_{i,j})\leq\sum_{k=i}^{j-1}\nu(A_{k}\setminus A_{k+1})<2^{1-i}.$
In addition, for each $i\in\omega$, $(B_{i,j}^{c}:j\geq i)$ is an increasing sequence with
$\sum_{j\geq i}\nu(B_{i,j+1}^{c}\setminus B_{i,j}^{c})\leq\sum_{j\geq i}\nu(A_{j}\setminus A_{j+1})<2^{1-i}.$
Hence, by assumption, for each $i\in\omega$ there exists $B_{i}\in\Sigma$ such that $\nu(B_{i}\setminus B_{i,j})=\nu(B_{i,j}^{c}\setminus B_{i}^{c})=0$ for all $j\geq i$, and $\lim_{j}\nu(B_{i,j}\setminus B_{i})=\lim_{j}\nu(B_{i}^{c}\setminus B_{i,j}^{c})=0$. Thus
(7) $\lim_{j\to\infty}\nu(B_{i,j}\bigtriangleup B_{i})=0.$
At this point, for all $i\in\omega$, define the $\Sigma$-measurable set
$C_{i}:=\bigcup_{k=0}^{i}B_{k}.$
Observe that, for all integers $i,k\in\omega$ with $k\geq i$, we have
$\begin{split}\nu(B_{i}\triangle C_{i})&\leq\sum_{j=0}^{i-1}\nu(B_{j}\setminus B_{j+1})\\ &\leq\sum_{j=0}^{i-1}(\nu(B_{j}\setminus B_{j,k})+\nu(B_{j,k}\setminus B_{j+1,k})+\nu(B_{j+1,k}\setminus B_{j+1}))\\ &=\sum_{j=0}^{i-1}\nu(B_{j+1,k}\setminus B_{j+1}).\end{split}$
By the hypothesis as applied to each $B_{j}$ and letting $k\to\infty$, this implies that
(8) $\forall i\in\omega,\quad\nu(B_{i}\bigtriangleup C_{i})=0.$
Moreover, $(C_{i}:i\in\omega)$ is an increasing sequence and, for all $i,j\in\omega$ with $i\leq j$,
$C_{i+1}\setminus C_{i}$ $\subseteq$ $B_{i+1}\setminus B_{i}$
$\subseteq$ $(B_{i+1}\setminus B_{i+1,j})\cup(B_{i+1,j}\setminus B_{i,j})\cup(B_{i,j}\setminus B_{i})$
$\subseteq$ $(B_{i+1}\setminus B_{i+1,j})\cup(A_{i+1}\setminus A_{i})\cup(B_{i,j}\setminus B_{i}).$
Since $\nu$ is a submeasure then
$\nu(C_{i+1}\setminus C_{i})\leq\nu(B_{i+1}\setminus B_{i+1,j})+\nu(A_{i+1}\setminus A_{i})+\nu(B_{i,j}\setminus B_{i})<2^{-i}$
for sufficiently large $j\in\omega$ and thus $\sum_{i}\nu(C_{i+1}\setminus C_{i})<\infty.$ By the standing assumption, there exists $A\in\Sigma$ such that $\nu(C_{i}\setminus A)=0$ for all $i\in\omega$, and $\lim_{i}\nu(A\setminus C_{i})=0$. In particular,
(9) $\lim_{i\to\infty}\nu(C_{i}\bigtriangleup A)=0.$
To conclude, we claim that $(A_{n})$ is $d_{\nu}$-convergent to AA. In fact, observe that for all $i,j\in\omega$ with $i\leq j$, we have
$A_{i}\triangle A\subseteq(A_{i}\triangle B_{i,j})\cup(B_{i,j}\triangle B_{i})\cup(B_{i}\triangle C_{i})\cup(C_{i}\triangle A).$
Since $\nu$ is a submeasure, it follows by inequalities (6), (7), and (8) that
$\forall i\in\omega,\quad\nu(A_{i}\triangle A)<2^{1-i}+\nu(C_{i}\triangle A).$
Finally, it follows by (9) that $\lim_{i}\nu(A_{i}\triangle A)=0$. Hence $(\Sigma,d_{\nu})$ is complete. ∎
We continue with the proofs of its applications.
Proof of Corollary $2.$
Pick an increasing sequence $(A_{n})$ in $\Sigma$. Observe that $\sum_{n}\mu(A_{n+1}\setminus A_{n})<\infty$: otherwise there would exist $n_{0}\in\omega$ such that $\mu(X)<\sum_{n\leq n_{0}}\mu(A_{n+1}\setminus A_{n})=\mu(A_{n_{0}+1}\setminus A_{0})\leq\mu(A_{n_{0}+1})$, which is impossible. Suppose that there exists $A\in\Sigma$ such that $\mu(A_{n}\setminus A)=0$ for all $n\in\omega$ and $\lim_{n}\mu(A\setminus A_{n})=0$. Since $|\mu(A)-\mu(A_{n})|\leq\mu(A_{n}\setminus A)+\mu(A\setminus A_{n})$ for all $n\in\omega$, then $\lim_{n}\mu(A_{n})=\mu(A)$. Conversely, suppose that there exists $A\in\Sigma$ such that $\mu(A_{n}\setminus A)=0$ for all $n\in\omega$ and $\lim_{n}\mu(A_{n})=\mu(A)$. Since $\mu(A\setminus A_{n})=\mu(A)-\mu(A_{n})+\mu(A_{n}\setminus A)$, then $\lim_{n}\mu(A\setminus A_{n})=0$. This proves that item (ii) of Corollary 2.2 is a rewriting of item (iii) of Theorem 2.1 for nonnegative finitely additive bounded maps. The conclusion follows by Theorem 2.1. ∎
Proof of Corollary 2.3.
Pick an increasing sequence $(A_{n})$ in $\Sigma$ such that $\sum_{n}\nu(A_{n+1}\setminus A_{n})<\infty$ and define $A:=\bigcup_{k}A_{k}\in\Sigma$. Then $A_{n}\setminus A=\emptyset$ for all $n\in\omega$. In addition, since $A\setminus A_{n}=\bigcup_{k\geq n}(A_{k+1}\setminus A_{k})$ for all $n\in\omega$, we obtain
$\limsup_{n\to\infty}\nu(A\setminus A_{n})\leq\limsup_{n\to\infty}\sum_{j\geq n}\nu(A_{j+1}\setminus A_{j})=0.$
The conclusion follows by Theorem 2.1. ∎
Remark 3.1.
Item (iii) in Theorem 2.1 cannot be replaced by:
  1. (ii)
    For every increasing sequence $(A_{n}:n\in\omega)$ in $\Sigma$ such that $\sum_{n}\nu(A_{n+1}\setminus A_{n})<\infty$, there exists $A\in\Sigma$ such that $A_{n}\setminus A\in\mathrm{Fin}$ for all $n\in\omega$ and $\lim_{n}\nu(A\setminus A_{n})=0$.
In fact, choose $\Sigma=\mathcal{P}(\omega)$ and let $\mu$ be as in the statement of Theorem 1.2. Then $(\mathcal{P}(\omega),d_{\mu})$ is complete, i.e., item (i) in Theorem 2.1 holds. On the other hand, there exists an increasing sequence $(A_{n}:n\in\omega)$ in $\mathcal{P}(\omega)$ such that, if $A\subseteq\omega$ satisfies $A_{n}\setminus A\in\mathrm{Fin}$ for all $n\in\omega$, then $\mu(A)\neq\lim_{n}\mu(A_{n})$. Since $\mu$ is invariant modulo finite sets then there exists $\kappa>0$ such that $\mu(A)\geq\mu(A_{n})+\kappa$ for all $n\in\omega$. Since $\mu(A\setminus A_{n})=\mu(A)-\mu(A_{n})+\mu(A_{n}\setminus A)$ for all $n\in\omega$, we obtain $\liminf_{n}\mu(A\setminus A_{n})\geq\kappa$. In addition, $\sum_{n}\mu(A_{n+1}\setminus A_{n})<\infty$, by the same argument as in the proof of Corollary $2.$ Therefore item (ii) above fails.
Remark 3.2.
In the same vein as Remark 3.1, there exists a finitely additive map $\mu:\mathcal{P}(\omega)\to[0,1]$ satisfying property AP(null) and such that its zero set $\mathcal{I}:=\{A\subseteq\omega:\mu(A)=0\}$ is not a $P$-ideal.
In fact, let $\mu$ be an additive extension of the asymptotic density $\mathsf{d}$ satisfying the claim of Theorem 1.2. Then by item (i) the induced pseudometric $d_{\mu}$ is complete, hence AP(null) holds by Corollary $2.$ In addition, by item (ii), it is possible to pick an increasing sequence $(A_{n}:n\in\omega)$ in $\mathcal{P}(\omega)$ such that, if $B\subseteq\omega$ satisfies $A_{n}\setminus B\in\mathrm{Fin}$ for all $n\in\omega$, then $\mu(B)\neq\lim_{n}\mu(A_{n})$. Thanks to AP(null), there exists $A\subseteq\omega$ such that $C_{n}:=A_{n}\setminus A\in\mathcal{I}$ for all $n\in\omega$ and $\mu(A)=\lim_{n}\mu(A_{n})$. Suppose for the sake of contradiction that $\mathcal{I}$ is a $P$-ideal. Then there exists $C\in\mathcal{I}$ such that $C_{n}\setminus C\in\mathrm{Fin}$ for all $n\in\omega$. Lastly, define $B:=A\cup C$. Then $A_{n}\setminus B=C_{n}\setminus C\in\mathrm{Fin}$ for all $n\in\omega$, and $\mu(B)=\mu(A)=\lim_{n}\mu(A_{n})$. This provides the claimed contradiction.
It is natural to investigate whether the topological and uniform properties of the pseudometric $d_{\|\cdot\|_{\varphi}}$ induced by an lscsm $\varphi$ can be deduced from some property of the exhaustive ideal $\mathrm{Exh}(\varphi)$. For example, below we identify two lscsms with the same exhaustive ideal that also induce metrically equivalent pseudometrics.
Example $3.$
Define $I_{n}:=\omega\cap[2^{n},2^{n+1})$ for all $n\in\omega$, and let $\varphi:\mathcal{P}(\omega)\to[0,\infty]$ and $\psi:\mathcal{P}(\omega)\to[0,\infty]$ be the lscsms given by
$\varphi(A):=\sup_{n\geq 1}\,\frac{|A\cap n|}{n}\quad\text{ and }\quad\psi(A):=\sup_{n\in\omega}\,\frac{|A\cap I_{n}|}{|I_{n}|}$
for all $A\subseteq\omega$. It is clear that $\|A\|_{\varphi}=\mathsf{d}^{\star}(A)$ and $\|A\|_{\psi}=\limsup_{n}\,|A\cap I_{n}|/|I_{n}|$ for each $A\subseteq\omega$. Note
$\mathcal{Z}=\mathrm{Exh}(\varphi)=\mathrm{Exh}(\psi)$
(see [4, Lemma 3.1] or [37, Theorem 2]; cf. also the proof of [18, Theorem 1.13.3(a)], which shows that $\mathcal{Z}$ is a density-ideal).
In addition, $d_{\|\cdot\|_{\psi}}$ and $d_{\mathsf{d}^{\star}}$ are metrically equivalent or, more precisely,
$\forall A\subseteq\omega,\quad\frac{\|A\|_{\psi}}{2}\leq\mathsf{d}^{\star}(A)\leq 16\|A\|_{\psi}.$
For, fix $A\subseteq\omega$ such that $\alpha:=\mathsf{d}^{\star}(A)\in(0,1]$. On the one hand, for each $\varepsilon>0$ there exist infinitely many $n\in\omega$ such that $|A\cap I_{n}|\geq 2^{n}(1-\varepsilon)\|A\|_{\psi}$. Since $A\cap I_{n}\subseteq A\cap 2^{n+1}$, it follows that $|A\cap 2^{n+1}|\geq 2^{n}(1-\varepsilon)\|A\|_{\psi}$. Since $\varepsilon>0$ is arbitrary, we obtain $\|A\|_{\psi}\leq 2\mathsf{d}^{\star}(A)$. On the other hand, observe that there exists $n_{0}\in\omega$ such that $|A\cap n|\leq\frac{5}{4}\alpha n$ for all $n\geq n_{0}$. Also, there exists an infinite set $S\subseteq\omega$ such that $\min S\geq 2n_{0}$ and $|A\cap n|\geq\frac{3}{4}\alpha n$ for all $n\in S$. Now, pick $n\in S$ and $m\in\omega$ such that $n\in I_{m}$. Since $\frac{n}{2}\geq n_{0}$ it follows that $|A\cap\frac{n}{2}|\leq\frac{5}{8}\alpha n$, hence $|A\cap[\frac{n}{2},n)|\geq\frac{1}{8}\alpha n$. Thus $\max\{|A\cap I_{m-1}|,|A\cap I_{m}|\}\geq\frac{1}{16}\alpha n\geq\frac{1}{16}\,2^{m}\alpha$. Hence there exist infinitely many $k\in\omega$ such that $|A\cap I_{k}|\geq\frac{1}{16}\alpha|I_{k}|$, so that $\|A\|_{\psi}\geq\frac{1}{16}\mathsf{d}^{\star}(A)$. Consequently, $d_{\|\cdot\|_{\psi}}$ is complete if and only if $d_{\mathsf{d}^{\star}}$ is complete (cf. Theorem 1.5).
However, it is not always the case that lscsms with the same exhaustive ideal induce metrically equivalent pseudometrics.
Proposition 3.4.
There exists a lscsm $\varphi:\mathcal{P}(\omega)\to[0,\infty]$ such that:
  1. (i)
    $\mathrm{Exh}(\varphi)=\mathcal{Z}$;
  2. (ii)
    $\|\cdot\|_{\varphi}$ is an upper density;
  3. (iii)
    $d_{\|\cdot\|_{\varphi}}$ is not metrically equivalent to $d_{\mathsf{d}^{\star}}$.
Proof.
For each real $\alpha>-1$, let $\varphi_{\alpha}:\mathcal{P}(\omega)\to[0,\infty]$ be the lscsm defined by
$\forall A\subseteq\omega,\quad\varphi_{\alpha}(A):=\sup_{n\geq 1}\frac{\sum_{i\in A\cap[1,n]}i^{\alpha}}{\sum_{i=1}^{n}i^{\alpha}}.$
Then $\varphi_{0}$ coincides with the lscsm $\varphi$ used in Example 3.3 and $\|\cdot\|_{\varphi_{\alpha}}$ is the classical upper $\alpha$-density, cf. [38, Example 4]. In addition, it follows by [39, Theorem 4.1 and Theorem 4.3] that $\mathcal{Z}=\mathrm{Exh}({\varphi_{\alpha}})$ for each $\alpha>-1$. Thus, define the map $\varphi_{\infty}:\mathcal{P}(\omega)\to[0,\infty]$ by
$\forall A\subseteq\omega,\quad\varphi_{\infty}(A):=\sum_{\alpha\in\omega}\frac{\varphi_{2^{\alpha}}(A)}{2^{\alpha}}.$
Then $\varphi_{\infty}$ is a lscsm that satisfies our claim.
It is immediate to see that $\varphi_{\infty}$ is monotone, subadditive, and $\varphi_{\infty}(\omega)=1$. Moreover, for each $\varepsilon>0$ and $\alpha>-1$ there exists $n_{\alpha,\varepsilon}\in\omega$ such that $\varphi_{2^{\alpha}}(A)\leq\varphi_{2^{\alpha}}(A\cap n_{\alpha,\varepsilon})+\varepsilon$ for $n\geq n_{\alpha,\varepsilon}$ (since $\varphi_{2^{\alpha}}$ is an lscsm). Now, fix $\varepsilon>0$, pick $\alpha_{0}\in\omega$ such that $2^{-\alpha_{0}}\leq\varepsilon/3$, and define $n_{\varepsilon}:=\max\{n_{\alpha,\varepsilon/3}:\alpha\in\alpha_{0}+1\}$. It follows that
$\begin{split}\varphi_{\infty}(A)&\leq\sum_{\alpha=0}^{\alpha_{0}}\frac{\varphi_{2^{\alpha}}(A)}{2^{\alpha}}+\sum_{\alpha>\alpha_{0}}\frac{1}{2^{\alpha}}=\sum_{\alpha=0}^{\alpha_{0}}\frac{\varphi_{2^{\alpha}}(A)}{2^{\alpha}}+\frac{1}{2^{\alpha_{0}}}\\ &\leq\sum_{\alpha=0}^{\alpha_{0}}\frac{\varphi_{2^{\alpha}}(A\cap n_{\alpha,\varepsilon/3})+\varepsilon/3}{2^{\alpha}}+\frac{\varepsilon}{3}\\ &\leq\sum_{\alpha=0}^{\alpha_{0}}\frac{\varphi_{2^{\alpha}}(A\cap n_{\varepsilon})}{2^{\alpha}}+\frac{\varepsilon}{3}\left(1+\sum_{\alpha=0}^{\alpha_{0}}\frac{1}{2^{\alpha}}\right)\leq\varphi_{\infty}(A\cap n_{\varepsilon})+\varepsilon.\end{split}$
This proves that $\varphi_{\infty}$ is a (bounded) lscsm.
By the Dominated Convergence theorem, we have that
(10) $\forall A\subseteq\omega,\quad\|A\|_{\varphi_{\infty}}=\lim_{n\to\infty}\sum_{\alpha\in\omega}\frac{\varphi_{2^{\alpha}}(A\setminus n)}{2^{\alpha}}=\sum_{\alpha\in\omega}\frac{\|A\|_{\varphi_{2^{\alpha}}}}{2^{\alpha}}.$
Since each $\|\cdot\|_{\varphi_{\alpha}}$ is an upper density and the set of upper densities is countably convex by [38, Proposition 10], then $\|\cdot\|_{\varphi_{\infty}}$ is an upper density as well, i.e., item (ii) holds. Moreover, by the above observations, identity (10) implies that $\mathcal{Z}=\mathrm{Exh}(\varphi_{\infty})$, i.e., item (i) holds.
Lastly, we claim that there is no constant $C>0$ such that $\|A\|_{\varphi_{\infty}}\leq C\|A\|_{\psi}$ for all $A\subseteq\omega$, where $\psi$ is the lscsm used in Example $3.$ For, define $A_{n}:=\omega\cap\bigcup_{i\in\omega}[2^{i},2^{i}(1+2^{-n}))$ for each $n\in\omega$, and observe that $\|A_{n}\|_{\psi}=2^{-n}$. Then, for each $n\in\omega$ and each $\alpha>-1$, we obtain
$\begin{split}\|A_{n}\|_{\varphi_{\alpha}}&=\limsup_{k\to\infty}\frac{\sum_{i\in A_{n}\cap[1,2^{k}(1+2^{-n})]}i^{\alpha}}{\sum_{i=1}^{2^{k}(1+2^{-n})}i^{\alpha}}\\ &\geq\limsup_{k\to\infty}\frac{\sum_{i=2^{k}}^{2^{k}(1+2^{-n})}i^{\alpha}}{\sum_{i=1}^{2^{k}(1+2^{-n})}i^{\alpha}}=1-(1+2^{-n})^{-{\alpha}-1}.\end{split}$
Putting it all together, it follows that
(11) $\forall\alpha>-1,\quad\liminf_{n\to\infty}\frac{\|A_{n}\|_{\varphi_{\alpha}}}{\|A_{n}\|_{\psi}}\geq\lim_{n\to\infty}\frac{1-(1+2^{-n})^{-\alpha-1}}{2^{-n}}=\alpha+1.$
Now, fix $C\in\omega$. Thanks to (11), there exists a sufficiently large $n_{C}\in\omega$ such that
$\forall\alpha\in 2C,\quad\frac{\|A_{n_{C}}\|_{\varphi_{2^{\alpha}}}}{\|A_{n_{C}}\|_{\psi}}\geq\frac{2^{\alpha}+1}{2}.$
Since $\frac{2^{\alpha}+1}{2}>2^{\alpha-1}$, it follows by (10) that
$\|A_{n_{C}}\|_{\varphi_{\infty}}\geq\sum_{\alpha\in 2C}\frac{\|A_{n_{C}}\|_{\varphi_{2^{\alpha}}}}{2^{\alpha}}>\|A_{n_{C}}\|_{\psi}\sum_{\alpha\in 2C}\frac{2^{\alpha-1}}{2^{\alpha}}=C\|A_{n_{C}}\|_{\psi}.$
Therefore $d_{\|\cdot\|_{\varphi_{\infty}}}$ and $d_{\|\cdot\|_{\psi}}$ are not metrically equivalent. Since $d_{\|\cdot\|_{\psi}}$ and $d_{\mathsf{d}^{\star}}$ are metrically equivalent by Example 3.3, we conclude that item (iii) holds. ∎
It is true that lscsms with the same exhaustive ideal induce pseudometrics that are topologically equivalent.
Proposition 3.5.
Let $\varphi_{1},\varphi_{2}$ be two lscsms such that $\mathrm{Exh}(\varphi_{1})=\mathrm{Exh}(\varphi_{2})$. Then the pseudometrics $d_{\|\cdot\|_{\varphi_{1}}}$ and $d_{\|\cdot\|_{\varphi_{2}}}$ are topologically equivalent.
Proof.
It is easy to check that $\mathrm{Exh}(\varphi_{1})=\mathrm{Exh}(\varphi_{2})$ is equivalent to the fact that, for every disjoint sequence $(F_{n})$ of finite sets of $\omega$, we have $\lim_{n}\varphi_{1}(F_{n})=0$ if and only if $\lim_{n}\varphi_{2}(F_{n})=0$, cf. [12, Remark 2.3]. Now, pick a sequence of sets $(A_{n})$ and $A\subseteq\omega$. Since $\varphi_{1}$ and $\varphi_{2}$ are two lscsms, there exists a partition $(P_{n}:n\in\omega)$ of $\omega$ into nonempty finite sets such that
$|\varphi_{i}((A_{n}\bigtriangleup A)\cap P_{n})-\|A_{n}\bigtriangleup A\|_{\varphi_{i}}|\leq 2^{-n}$
for each $i\in\{1,2\}$ and for all sufficiently large $n$. It follows that $\lim_{n}\|A\bigtriangleup A_{n}\|_{\varphi_{1}}=0$ if and only if $\lim_{n}\|A\bigtriangleup A_{n}\|_{\varphi_{2}}=0$, concluding the proof. ∎
Next we prove Proposition 2.4 and Theorem 2.5 using Theorem 2.1.
Proof of Proposition 2.4.
Let $(A_{n}:n\in\omega)$ be an increasing sequence in $\mathcal{P}(\omega)$ such that $\sum_{n}\bm{1}_{\mathcal{I}^{+}}(A_{n+1}\setminus A_{n})<\infty$. Define $S:=\{n\in\omega:A_{n+1}\setminus A_{n}\in\mathcal{I}^{+}\}$, so that $S$ is necessarily finite by the standing hypothesis. Define $A:=A_{0}$ if $S=\emptyset$, otherwise $A:=A_{1+\max S}$. It easily follows that $\nu(A_{n}\setminus A)=0$ for all $n\in\omega$, and that the real sequence $(\nu(A\setminus A_{n}):n\in\omega)$ is eventually 0. Therefore item (iii) holds. The claim follows by Theorem 2.1. ∎
Proof of Theorem 2.5.
Let $(A_{n}:n\in\omega)$ be an increasing sequence of sets such that $\sum_{n}\|A_{n+1}\setminus A_{n}\|_{\varphi}<\infty$. Thanks to Theorem 2.1, it will be enough to show that there exists $A\subseteq\omega$ such that $A_{n}\setminus A$ is finite for all $n\in\omega$ and $\lim_{n}\|A\setminus A_{n}\|_{\varphi}=0$ (note that this is stronger than item (iii) in Theorem 2.1).
To this aim, define $B_{j}:=A_{j+1}\setminus A_{j}$ for all $j\in\omega$. Let $(n_{j}:j\in\omega)$ be a strictly increasing sequence in $\omega$ such that
(12) $\forall j\in\omega,\quad\varphi(B_{j}\setminus n_{j})\leq 2\|B_{j}\|_{\varphi}.$
We claim that $A:=A_{0}\cup\bigcup_{j\in\omega}(B_{j}\setminus n_{j})$ satisfies our claim. For, note that $A_{k}\setminus A\subseteq n_{k}$ for all $k\in\omega$. To complete the proof, we need to show that $\lim_{n}\|A\setminus A_{n}\|_{\varphi}=0$. To this aim, fix $\varepsilon>0$ and pick $k_{0}\in\omega$ such that $\sum_{j\geq k_{0}}\|B_{j}\|_{\varphi}<\varepsilon/4$. Pick an integer $k\geq k_{0}$. Taking into account that $\|\cdot\|_{\varphi}$ is invariant under finite modifications and that the elements of $(B_{j})$ are pairwise disjoint, it follows that
(13) $\begin{split}\|A\bigtriangleup A_{k}\|_{\varphi}\leq\left\|\,n_{k}\cup\bigcup_{j\geq k}((B_{j}\setminus n_{j})\setminus A_{k})\right\|_{\varphi}=\left\|\,\bigcup_{j\geq k}(B_{j}\setminus n_{j})\right\|_{\varphi}.\end{split}$
Since $\varphi$ is a lscsm, there exists an integer $m_{k}>n_{k+1}$ such that
(14) $\left\|\,\bigcup_{j\geq k}(B_{j}\setminus n_{j})\right\|_{\varphi}\leq 2\varphi\left(m_{k}\cap\bigcup_{j\geq k}(B_{j}\setminus n_{j})\right).$
Fix an integer $i_{k}\in\omega$ such that $n_{i_{k}}>m_{k}$. Since $\min(B_{j}\setminus n_{j})\geq n_{j}>m_{k}$ for all $j\geq i_{k}$, it follows by (12), (13), (14), and the above observations that
$\begin{split}\|A\bigtriangleup A_{k}\|_{\varphi}&\leq 2\varphi\left(\bigcup_{k\leq j<i_{k}}(B_{j}\setminus n_{j})\right)\\ &\leq 2\sum_{k\leq j<i_{k}}\varphi(B_{j}\setminus n_{j})\leq 2\sum_{j\geq k}\varphi(B_{j}\setminus n_{j})\\ &\leq 4\sum_{j\geq k}\|B_{j}\|_{\varphi}\leq 4\sum_{j\geq k_{0}}\|B_{j}\|_{\varphi}<\varepsilon.\end{split}$
Therefore $\lim_{n}\|A\bigtriangleup A_{n}\|_{\varphi}=0$. ∎
Remark 3.6.
Solecki proved also in [45, Theorem 3.1] that, if $\varphi:\mathcal{P}(\omega)\to[0,\infty]$ is a lscsm, then $\mathrm{Exh}(\varphi)$ is $d_{\varphi}$-closed and separable. Since $d_{\|\cdot\|_{\varphi}}(A,B)\leq d_{\varphi}(A,B)$ for all $A,B\subseteq\omega$, then $\mathrm{Exh}(\varphi)$ is $d_{\|\cdot\|_{\varphi}}$-closed and separable as well. Alternatively, this can be seen directly: since $\|\cdot\|_{\varphi}$ is $d_{\|\cdot\|_{\varphi}}$-continuous, then the $\|\cdot\|_{\varphi}$-preimage of $\{0\}$, that is, $\mathrm{Exh}(\varphi)$, is closed (in this case, the separability is obvious).
We proceed with the proofs of our main results.
Proof of Proposition 2.6.
Let us suppose for the sake of contradiction that the space $(\mathcal{P}(\omega),d_{\nu})$ is complete. Set $A_{n}:=\{0,1,\ldots,n\}$ for each $n\in\omega$, and observe that $\sum_{n}\nu(A_{n+1}\setminus A_{n})\leq\sum_{n}a_{n}<\infty$. It follows by Theorem 2.1 that there exists $A\subseteq\omega$ such that $\nu(A_{n}\setminus A)=0$ for all $n\in\omega$ and $\lim_{n}\nu(A\setminus A_{n})=0$. By the definition of $\nu$, the former condition implies that $A_{n}\subseteq A$ for all $n\in\omega$. Hence necessarily $A=\omega$. On the other hand, we have $\nu(A\setminus A_{n})\geq\bm{1}_{\mathcal{I}^{+}}(\omega\setminus n+1)=1$ for all $n\in\omega$. This provides the claimed contradiction. ∎
Proof of Theorem 2.7.
Let $\mu^{\star}$ be an upper density on $\omega$ such that $\mathsf{bd}^{\star}(A)\leq\mu^{\star}(A)$ for all $A\subseteq\omega$. Thanks to Theorem 2.1, it will be enough to show that for each $\kappa\in(0,1)$ there exists an increasing sequence $(A_{n}:n\in\omega)$ of subsets of $\omega$ such that $\sum_{n}\mu^{\star}(A_{n+1}\setminus A_{n})<\infty$ and, if a set $A\subseteq\omega$ satisfies $\mu^{\star}(A_{n}\setminus A)=0$ for all $n\in\omega$, then $\mu^{\star}(A\setminus A_{n})\geq\kappa$ for all $n\in\omega$.
To this aim, fix $\kappa\in(0,1)$ and a positive integer $N>e^{2}$ such that
(15) $\frac{\log^{2}(N)}{N}<\frac{1-\kappa}{2}.$
Define the positive integer $C:=\lceil 1/\kappa\rceil$ and let $(a_{n}:n\in\omega)$ be a strictly increasing sequence of positive integers with the property that
(16) $\sum_{j\in\omega}\frac{j}{a_{j}!}<\frac{1-\kappa}{2}\quad\text{ and }\quad a_{n}!>\max\left\{\frac{n}{\kappa},\,Ce^{2(n+1)},\,(N+1)^{2}\right\}$
for all $n\in\omega$. Thus, define recursively three sequences of sets of nonnegative integers $(A_{n}:n\in\omega)$, $(B_{n}:n\in\omega)$ and $(H_{n}:n\in\omega)$ as follows:
  1. (i)
    Set $H_{0}:=B_{0}:=\emptyset$.
  2. (ii)
    Set $H_{1}:=\{0\}$, $A_{0}:=B_{1}:=a_{1}!\cdot(\omega\setminus\{0\})$.
    Observe that $A_{0}=a_{2}!\cdot\omega+\{a_{1}!j:j\in\{1,2,\ldots,a_{2}!/a_{1}!\}\}$ and $\mathsf{d}(A_{0})=\mathsf{d}(B_{1})=1/a_{1}!$; in addition, we have $|A_{0}\cap n|/n<1/a_{1}!$ for all $n\geq 1$, hence choosing $n=a_{2}!$ we get
    $|a_{2}!\setminus A_{0}|>a_{2}!(1-1/a_{1}!)\geq a_{2}>2.$
  3. (iii)
    Given a positive integer $n\geq 1$, suppose that all sets $(A_{i}:i\in n)$, $(B_{i}:i\in n+1)$, and $(H_{i}:i\in n+1)$ have been defined satisfying the following properties:
    1. (a)
      all sets $B_{i}$ are pairwise disjoint;
    2. (b)
      $H_{i}\subseteq a_{i}!$ and $|H_{i}|=i$ for all $i\in n+1$;
    3. (c)
      $i\geq\kappa\ell_{i}$ for all nonzero $i\in n+1$, where
      $\ell_{i}:=1+\max H_{i}-\min H_{i};$
    4. (d)
      $B_{i}=a_{i}!\cdot(\omega\setminus\{0\})+H_{i}$ for all $i\in n+1$;
    5. (e)
      $A_{i}=B_{0}\cup\cdots\cup B_{i+1}$ for all $i\in n$;
    6. (f)
      $|A_{i}\cap m|/m<\mathsf{d}(A_{i})=\sum_{j\in i+2}j/a_{j}!$ for all $m\geq 1$ and $i\in n$;
    (In particular, such conditions hold for $n=1$, thanks to item (ii).)
  4. (iv)
    It follows by item (iii) that there exists a unique $K_{n+1}\subseteq a_{n+1}!$ such that
    $A_{n-1}\setminus a_{n+1}!=\bigcup_{i\in n+1}B_{i}\setminus a_{n+1}!=a_{n+1}!\cdot(\omega\setminus\{0\})+K_{n+1}$
    and, by hypothesis (16), also
    $a_{n+1}!-|K_{n+1}|=a_{n+1}!\left(1-\sum_{i\in n+1}\frac{i}{a_{i}!}\right)\geq\kappa a_{n+1}!>n+1.$
    Hence it is possible to pick a set $H_{n+1}\subseteq a_{n+1}!\setminus K_{n+1}$ such that $|H_{n+1}|=n+1$ and $\max H_{n+1}$ is minimized.
  5. (v)
    Lastly, define
    (17) $B_{n+1}:=a_{n+1}!\cdot(\omega\setminus\{0\})+H_{n+1}\,\,\text{ and }\,\,A_{n}:=\bigcup_{i\in n+2}B_{i}.$
    To complete the induction, we have to show that the conditions in item (iii) hold for $n+1$. For, observe by item (iv) that $B_{n+1}\cap A_{n-1}=\emptyset$, hence $B_{n+1}\cap B_{i}=\emptyset$ for all $i\in n+1$. Together with the inductive step, this proves condition (a). Condition (b) follows by the construction in item (iv), while conditions (d) and (e) follow by the above definitions of $B_{n+1}$ and $A_{n}$ in (17). Since $B_{n+1}$ is a finite union of infinite arithmetic progressions (i.e., $B_{n+1}\in\mathscr{A}$) and it is disjoint from $A_{n-1}$, we obtain
    $\mathsf{d}(A_{n})=\mathsf{d}(A_{n-1})+\mathsf{d}(B_{n+1})=\sum_{j\in n+1}\frac{j}{a_{j}!}+\frac{|H_{n+1}|}{a_{n+1}!}=\sum_{j\in n+2}\frac{j}{a_{j}!},$
    cf. also [38, Proposition 7 and Proposition 8]. In addition, by the definitions in (17), it is not difficult to observe (since $B_{i}\cap a_{i}!=\emptyset$ for all ii) that $|A_{n}\cap m|/m<\mathsf{d}(A_{n})$ for all $m\geq 1$, proving also condition (f). Now, note that for all nonzero $i\in n+2$ we have by (16) that
    $\begin{split}\ell_{i}&\leq\max H_{i}\leq\sum_{j=0}^{i}j\left\lceil\frac{\ell_{i}}{a_{j}!}\right\rceil\\ &\leq\sum_{j=0}^{i}j+\ell_{i}\sum_{j\in\omega}\frac{j}{a_{j}!}\leq i^{2}+\ell_{i}(1-\kappa),\end{split}$
    which implies that
    $\ell_{i}\leq\frac{i^{2}}{\kappa}\quad\text{ and }\quad\max H_{i}\leq i^{2}+\ell_{i}(1-\kappa)\leq\frac{i^{2}}{\kappa}.$
    Recalling that $C=\lceil 1/\kappa\rceil$, it follows that
    (18) $\forall i\in n+2,\quad B_{i}\subseteq a_{i}!\cdot\omega+\{0,1,\ldots,Ci^{2}\}.$
    In addition, since $a_{0}!>(N+1)^{2}$ by (16), it follows by the construction of the sets $H_{i}$ that
    $\forall i\in N+2,\quad B_{i}=a_{i}!\cdot(\omega\setminus 0)+\left\{\binom{i}{2}+j:j\in i\right\}.$
    The latter identity implies that, if $n\in\{1,\ldots,N\}$, then $A_{n-1}\subseteq a_{0}!\cdot\omega+[0,1,\ldots,\max H_{n}]$ and $(a_{n+1}!+[\min H_{n+1},\max H_{n+1}])\subseteq a_{0}!\cdot\omega+[\min H_{n+1},\max H_{n+1}]$. Considering also that $\max H_{n}<\min H_{n+1}$ by construction, we get
    $\forall n\in\{1,\ldots,N\},\quad A_{n-1}\cap(a_{n+1}!+[\min H_{n+1},\max H_{n+1}])=\emptyset.$
    In particular, if $n\in\{1,\ldots,N\}$, then $\ell_{n+1}=n+1$. Suppose now that $n\geq N+1$ and note by (15) that $\log^{2}(n)/n\leq\log^{2}(N)/N<(1-\kappa)/2$. Observe that, if $i>\log(n)$ then $i>\log((n+1)/e)$, which is equivalent to $n+1<e^{i+1}$. Hence it follows by (16) and (18) that, if $i>\log(n)$, then
    $\max H_{n+1}\leq C(n+1)^{2}<Ce^{2(i+1)}\leq a_{i}!.$
    This implies that
    $\begin{split}|A_{n-1}\cap(a_{n+1}!&+[\min H_{n+1},\max H_{n+1}])|\\ &\leq\sum_{i=0}^{n}|B_{i}\cap(a_{n+1}!+[\min H_{n+1},\max H_{n+1}])|\\ &\leq\ell_{n+1}\left(\sum_{i=0}^{n}\frac{i}{a_{i}!}\right)+\sum_{i=0}^{\lfloor\log(n)\rfloor}i\\ &\leq\ell_{n+1}\left(\sum_{i\in\omega}\frac{i}{a_{i}!}+\frac{\log^{2}(n)}{\ell_{n+1}}\right)\\ &\leq\ell_{n+1}\left(\frac{1-\kappa}{2}+\frac{\log^{2}(n)}{n}\right)\\ &\leq\ell_{n+1}(1-\kappa).\end{split}$
    Hence, in both cases (i.e., if $n\leq N$ or if $n\geq N+1$), we get that $|A_{n-1}\cap(a_{n+1}!+[\min H_{n+1},\max H_{n+1}])|\leq\ell_{n+1}(1-\kappa)$. Therefore
    $\begin{split}n+1&=|H_{n+1}|=|B_{n+1}\cap(a_{n+1}!+[\min H_{n+1},\max H_{n+1}])|\\ &=\ell_{n+1}-|A_{n-1}\cap(a_{n+1}!+[\min H_{n+1},\max H_{n+1}])|\geq\kappa\ell_{n+1}.\end{split}$
    This proves condition (c) and completes the induction.
To complete the proof, we claim that the sequence of sets $(A_{n}:n\in\omega)$ in the above construction satisfies our claim. For, we have by construction that $A_{n+1}\setminus A_{n}=B_{n+2}\in\mathscr{A}$ for all $n\in\omega$. Since $\mu^{\star}(B_{n})=|H_{n}|/a_{n}!$ by [38, Proposition 7], it follows by (16) and the conditions in item (v) that
$\sum_{n\in\omega}\mu^{\star}(A_{n+1}\setminus A_{n})\leq\sum_{n\in\omega}\mu^{\star}(B_{n})=\sum_{n\in\omega}\frac{n}{a_{n}!}<\infty.$
Now, pick a set $A\subseteq\omega$ such that $\mu^{\star}(A_{n}\setminus A)=0$ for all $n\in\omega$. Define also
$\forall n\in\omega,\quad C_{n}:=B_{n}\setminus A\quad\text{ and }\quad D_{n}:=B_{n}\setminus C_{n},$
so that $\{C_{n},D_{n}\}$ is a partition of $B_{n}$, and $D_{n}\subseteq A$ for all $n\in\omega$. In addition, we get by monotonicity of $\mu^{\star}$ that $\mu^{\star}(C_{n})\leq\mu^{\star}(A_{n}\setminus A)=0$ for all $n\in\omega$, hence $\mu^{\star}(C_{n})=0$. Thanks to (5), we get $\mathsf{d}^{\star}(C_{n})=0$, i.e.,
(19) $\forall n\in\omega,\forall\varepsilon>0,\exists N_{n,\varepsilon}\geq 1,\forall m\geq N_{i,\varepsilon},\quad\frac{|C_{n}\cap m|}{m}<\varepsilon.$
It follows that for any positive integer $n$ and sufficiently large $m$ we have $|C_{n}\cap m|/m<1/a_{n}!$. This implies there exists $j_{n}\in\omega$ such that
$C_{n}\cap(a_{n}!j_{n}+[\min H_{n},\max H_{n}])=\emptyset.$
Hence, by the condition (c) in item (iii), we get
$\frac{|D_{n}\cap(a_{n}!j_{n}+[\min H_{n},\max H_{n}])|}{\ell_{n}}=\frac{|H_{n}|}{\ell_{n}}\geq\kappa.$
It follows by the definition of the upper Banach density that
(20) $\forall n\geq 1,\quad\mathrm{bd}^{\star}\left(\bigcup_{i\geq n}D_{i}\right)\geq\kappa.$
Therefore, for each $n\in\omega$, we get
$\mu^{\star}(A\setminus A_{n})\geq\mu^{\star}\left(\bigcup_{i\geq n+2}D_{i}\right)\geq\mathrm{bd}^{\star}\left(\bigcup_{i\geq n+2}D_{i}\right)\geq\kappa,$
which completes the proof. ∎
Remark 3.7.
The argument provided in the proof of Theorem 2.7 shows that, if $\mathscr{A}_{\infty}$ stands for the family of countably infinite unions of infinite arithmetic progressions $k\cdot\omega+h$, then
$\mathscr{A}_{\infty}\setminus\mathsf{dom}(\mathsf{bd})\neq\emptyset.$
In fact, suppose that $\kappa=1/2$, so that $\sum_{n}n/a_{n}!\leq 1/4$ by (16). Then the set $B:=\bigcup_{n}B_{n}\in\mathscr{A}_{\infty}$ contains $\bigcup_{i\geq 1}D_{n}$, hence $\mathsf{bd}^{\star}(B)\geq 1/2$ by (20). On the other hand, for each $m\geq 1$, there exists $k_{m}\geq 1$ for which $B\cap m=\bigcup_{i\in k_{m}}B_{i}\cap m$, so that by the definition of the sets $B_{i}$ we get
$\frac{|B\cap m|}{m}\leq\sum_{i\in k_{m}}\frac{|B_{i}\cap m|}{m}\leq\sum_{i\in k_{m}}\frac{i}{a_{i}!}\leq\frac{1}{4}.$
This implies that $\mathsf{bd}_{\star}(B)\leq\mathsf{d}_{\star}(B)\leq 1/4$, cf. (5). Therefore $B\notin\mathrm{dom}(\mathsf{bd})$.
This example also shows that the restriction of the upper Banach density $\mathsf{bd}^{\star}$ to $\mathcal{P}(\omega)\setminus\{A\subseteq\omega:\mathsf{bd}^{\star}(A)=0\}$ is not $\sigma$-subadditive. This comment applies to all upper densities $\mu^{\star}$ for which $\mu^{\star}(A)\geq\mathsf{bd}^{\star}(A)$ for all $A\subseteq\omega$.
Proof of Lemma 2.8.
Suppose that $(\Sigma,d_{\nu})$ is complete, and let $(A_{n}:n\in\omega)$ be a Cauchy sequence in $(\overline{\Sigma},d_{\nu^{\star}})$. Passing if needed to a suitable subsequence, we can suppose without loss of generality that $d_{\nu^{\star}}(A_{n},A_{m})\leq 2^{-\min\{n,m\}}$ for all $n,m\in\omega$. Now, for each $n\in\omega$, the set $A_{n}$ belongs to $\overline{\Sigma}$, hence there exists $C_{n}\in\Sigma$ such that $d_{\nu^{\star}}(A_{n},C_{n})\leq 2^{-n}$. Observe that the sequence $(C_{n}:n\in\omega)$ is $d_{\nu}$-Cauchy since, for all $n,m\in\omega$, we have
$d_{\nu}(C_{n},C_{m})\leq d_{\nu^{\star}}(C_{n},A_{n})+d_{\nu^{\star}}(A_{n},A_{m})+d_{\nu^{\star}}(A_{m},C_{m})<2^{2-\min\{n,m\}}.$
Since $(\Sigma,d_{\nu})$ is complete, the sequence $(C_{n}:n\in\omega)$ is $d_{\nu}$-convergent to some $C\in\Sigma$. To complete the proof, it is enough to note that $d_{\nu^{\star}}(A_{n},C)\leq d_{\nu^{\star}}(A_{n},C_{n})+d_{\nu^{\star}}(C_{n},C)$ for all $n\in\omega$, hence $(A_{n}:n\in\omega)$ is $d_{\nu^{\star}}$-convergent to $C$ as well. Therefore $(\overline{\Sigma},d_{\nu^{\star}})$ is complete. ∎
Remark 3.8.
Let $\Sigma$ and $\nu$ be as in the statement of Theorem 2.1, and let $\overline{\Sigma}$ be the closure of $\Sigma$ in $(\mathcal{P}(X),d_{\nu^{\star}})$, as in Section 2.3. Then
$\overline{\Sigma}=\left\{B\subseteq X:\forall\varepsilon>0,\exists A,C\in\Sigma,\quad A\subseteq B\subseteq C\,\text{ and }\,\nu(C\setminus A)<\varepsilon\right\}.$
In fact, it is clear that the family on the right hand side is contained in $\overline{\Sigma}$. To show the converse, pick $B\in\overline{\Sigma}$ and $\epsilon>0$. Then there is $D\in\Sigma$ such that $\nu^{\star}(B\triangle D)<\epsilon$. By definition of $\nu^{\star}$, there is $E\in\Sigma$ such that $B\triangle D\subseteq E$ and $\nu(E)<\epsilon$. Define $A:=D\setminus E\in\Sigma$ and $C:=D\cup E\in\Sigma$. Then $A\subseteq B\subseteq C$ and $\nu(C\setminus A)=\nu(E)<\epsilon$.
Lemma 3.9.
With the notations of Section 2.3, we have
[…]
Proof.
[…]
We remark that $\mathrm{Cl}_{\hat{\nu}}(\mathcal{C})$ is, in general, not a $\sigma$-field. More precisely, the following example shows that $\sigma(\mathcal{C})$ might not be contained in $\mathrm{Cl}_{\hat{\nu}}(\mathcal{C})$, even if the underlying pseudometric space is complete.
Example 3.10.
Set $X:=\omega$, $\Sigma:=\mathcal{P}(\omega)$, and define $\nu:\Sigma\to\mathbb{R}$ by
[…]
Note that $(\Sigma,d_{\nu})$ is complete. We use the same notations of Section 2.3, e.g., $S$ is the Stone space of $\Sigma/\nu$ and $\mathcal{C}$ is its field of clopen subsets.
If $B=\emptyset$ then $\hat{\nu}(B)=0$. If $B\neq\emptyset$ and $(A_{n})\in\Delta(B)$, then at least one $\phi(\pi(A_{n}))$ is nonempty, hence $\pi(A_{n})\neq\pi(\emptyset)$ and $\nu(A_{n})=1$. Thus $\sum_{n}\nu(A_{n})\geq 1$, and consequently $\hat{\nu}(B)\geq 1$. On the other hand, $B\subseteq S=\phi(\pi(X))$ and $\nu(X)=1$, so $\hat{\nu}(B)\leq 1$. It follows that
[…]
This implies that the topology induced by $d_{\hat{\nu}}$ on $\mathcal{P}(S)$ is the discrete one. In particular, $\mathrm{Cl}_{\hat{\nu}}(\mathcal{C})=\mathcal{C}$. Hence, it is enough to show that $\mathcal{C}$ is not a $\sigma$-field. To this end, fix a sequence $(I_{n}:n\in\omega)$ of $\omega$ of pairwise disjoint infinite sets and define $G_{n}:=\phi(\pi(I_{n}))$ for each $n\in\omega$. Since $(G_{n}:n\in\omega)$ is a sequence of pairwise disjoint nonempty clopen subsets of $S$ and $S$ is compact, it follows that $G:=\bigcup_{n}G_{n}\notin\mathcal{C}$. Therefore $\mathcal{C}$ is not a $\sigma$-field.
Taking into account that both $\sigma(\mathcal{C})$ and $\mathrm{Cl}_{\hat{\nu}}(\mathcal{C})$ are stable under complements and choosing $S\setminus G$, it follows that there exists a complete pseudometric space which admits a compact $G_{\delta}$ set in $\sigma(\mathcal{C})\setminus\mathrm{Cl}_{\hat{\nu}}(\mathcal{C})$.
Additional properties of $\mathrm{Cl}_{\hat{\nu}}(\mathcal{C})$ are given below, provided that $\nu$ is finite.
Lemma 3.11.
With the notations of Section 2.3, suppose that $\nu(X)<\infty$, and let $\mathscr{M}_{\hat{\nu}}$ be the family of Carathéodory $\hat{\nu}-$measurable sets, that is,
$\mathscr{M}_{\hat{\nu}}:=\left\{B\subseteq S:\forall C\subseteq S,\quad\hat{\nu}(C)=\hat{\nu}(C\cap B)+\hat{\nu}(C\setminus B)\right\}.$
Then $\mathscr{M}_{\hat{\nu}}$ is a $\sigma$-field, $\hat{\nu}$ is $\sigma$-additive on $\mathscr{M}_{\hat{\nu}}$, and $\mathcal{N}_{\hat{\nu}}\subseteq\mathscr{M}_{\hat{\nu}}\subseteq\mathrm{Cl}_{\hat{\nu}}(\mathcal{C})$.
Proof.
Recall that $\hat{\nu}$ is an outer measure on $\mathcal{P}(S)$ (that is, it is monotone, $\sigma$-additive, and satisfies $\hat{\nu}(\emptyset)=0$). Hence it follows by Carathéodory’s Theorem that $\mathscr{M}_{\hat{\nu}}$ is a $\sigma$-field which contains $\mathcal{N}_{\hat{\nu}}$, and the restriction of $\hat{\nu}$ on $\mathscr{M}_{\hat{\nu}}$ is $\sigma$-additive, cf. e.g. [26, Theorem 1.11].
Claim 1.
$\nu=\hat{\nu}\circ\phi\circ\pi$.
Proof.
Pick $A\in\Sigma$. Since $(A,\emptyset,\emptyset,\ldots)\in\Delta(\phi(\pi(A)))$, we have $\hat{\nu}(\phi(\pi(A)))\leq\nu(A)$. Conversely, suppose $(A_{n}:n\in\omega)\in\Delta(\phi(\pi(A)))$. Since $\{\phi(\pi(A_{n})):n\in\omega\}$ is an open cover of the compact set $\phi(\pi(A))$, there exists $F\in\mathrm{Fin}$ such that $\phi(\pi(A))\subseteq\bigcup_{n\in F}\phi(\pi(A_{n}))$. Since $\phi$ is a Boolean isomorphism and $\nu$ is a submeasure, we get $\nu(A)\leq\sum_{n\in F}\nu(A_{n})\leq\sum_{n\in\omega}\nu(A_{n}).$ Therefore $\nu(A)\leq\hat{\nu}(\phi(\pi(A)))$. ∎
To show the last inclusion $\mathscr{M}_{\hat{\nu}}\subseteq\mathrm{Cl}_{\hat{\nu}}(\mathcal{C})$, fix a set $B\in\mathscr{M}_{\hat{\nu}}$ and a real $\varepsilon>0$. Observe by Claim 1 that $\hat{\nu}(B)\leq\hat{\nu}(S)=\nu(X)<\infty$. By the definition of $\hat{\nu}$, there exists a sequence $(A_{n}:n\in\omega)\in\Delta(B)$ such that $\sum_{n}\nu(A_{n})<\hat{\nu}(B)+\nicefrac{{\varepsilon}}{{2}}$. In particular, the latter series is convergent and we can pick $n_{0}\in\omega$ such that $\sum_{n\geq n_{0}}\nu(A_{n})<\nicefrac{{\varepsilon}}{{2}}$. By $\sigma$-subadditivity of $\hat{\nu}$ and Claim 1 we get
$\hat{\nu}\left(B_{\infty}\right)\leq\sum_{n\in\omega}\hat{\nu}(\phi(\pi(A_{n})))=\sum_{n\in\omega}\nu(A_{n})<\hat{\nu}(B)+\frac{\varepsilon}{2},$
where $B_{\infty}:=\bigcup_{n}\phi(\pi(A_{n}))$. In addition, since $B\in\mathscr{M}_{\hat{\nu}}$ and $B\subseteq B_{\infty}$, we get $\hat{\nu}(B_{\infty})=\hat{\nu}(B)+\hat{\nu}(B_{\infty}\setminus B)$, hence $\hat{\nu}(B_{\infty}\setminus B)<\nicefrac{{\varepsilon}}{{2}}$. At this point, define the clopen set $C:=\bigcup_{n\in n_{0}}\phi(\pi(A_{n}))$. It follows by the above premises that
$\begin{split}\hat{\nu}(B\bigtriangleup C)&\leq\hat{\nu}(B\setminus C)+\hat{\nu}(C\setminus B)\\ &\leq\hat{\nu}\left(B_{\infty}\setminus C\right)+\hat{\nu}\left(B_{\infty}\setminus B\right)<\sum_{n\geq n_{0}}\nu(A_{n})+\frac{\varepsilon}{2}<\varepsilon.\end{split}$
Since $\varepsilon$ is arbitrary, it follows that $B\in\mathrm{Cl}_{\hat{\nu}}(\mathcal{C})$. Therefore $\mathscr{M}_{\hat{\nu}}\subseteq\mathrm{Cl}_{\hat{\nu}}(\mathcal{C})$. ∎
As an application, we show that $\mathrm{Cl}_{\hat{\nu}}(\mathcal{C})$ coincides with $\mathscr{M}_{\hat{\nu}}$ if and only if $\nu$ is additive.
Proposition 3.12.
Let $\nu:\Sigma\to\overline{\mathbb{R}}$ be a submeasure such that $\nu(X)<\infty$. Then the following are equivalent:
  1. (i)
    $\nu$ is finitely additive;
  2. (ii)
    $\mathcal{C}\subseteq\mathscr{M}_{\hat{\nu}}$;
  3. (iii)
    $\mathrm{Cl}_{\hat{\nu}}(\mathcal{C})=\mathscr{M}_{\hat{\nu}}$.
Proof.
(i) $\implies$ (ii). Pick $C\in\mathcal{C}$ and $A\in\Sigma$ such that $C=\phi(\pi(A))$. We need to show that $C\in\mathscr{M}_{\hat{\nu}}$. To this aim, fix a subset $B\subseteq S$. We proceed as in [26, Proposition 1.13(b)]: for each $k\in\omega$, there exists $(A_{k,n}:n\in\omega)\in\Delta(B)$ such that $\sum_{n}\nu(A_{k,n})\leq\hat{\nu}(B)+2^{-k}$. By the additivity of $\nu$, we obtain
$\hat{\nu}(B)+2^{-k}\geq\sum_{n\in\omega}\nu(A_{k,n}\cap A)+\sum_{n\in\omega}\nu(A_{k,n}\setminus A)\geq\hat{\nu}(B\cap C)+\hat{\nu}(B\setminus C),$
where the last inequality follows by the definition of $\hat{\nu}$. Since $\hat{\nu}$ is subadditive and $k$ is arbitrary, we conclude that $C\in\mathscr{M}_{\hat{\nu}}$.
(ii) $\implies$ (iii). Thanks to Lemma 3.11 and the standing hypothesis, $\mathscr{M}_{\hat{\nu}}$ is a $\sigma$-field containing both $\mathcal{N}_{\hat{\nu}}$ and $\mathcal{C}$. Hence $\mathrm{Cl}_{\hat{\nu}}(\mathcal{C})\subseteq\sigma(\mathcal{C}\cup\mathcal{N}_{\hat{\nu}})\subseteq\mathscr{M}_{\hat{\nu}}$ by Lemma 3.9. The opposite inclusion is contained also in Lemma 3.11.
(iii) $\implies$ (i). Pick disjoint sets $A,B\in\Sigma$. Then $\phi(\pi(A))$ and $\phi(\pi(B))$ are disjoint clopen subsets of $S$. In particular, by item (iii), the latter sets belong to $\mathscr{M}_{\hat{\nu}}$. At this point, recall that $\hat{\nu}$ is, in particular, additive on $\mathscr{M}_{\hat{\nu}}$, thanks to Lemma 3.11. Therefore
$\nu(A\cup B)=\hat{\nu}(\phi(\pi(A\cup B)))=\hat{\nu}(\phi(\pi(A)))+\hat{\nu}(\phi(\pi(B)))=\nu(A)+\nu(B),$
where we used Claim 1. ∎
As it follows by Example 3.10, it is possible that $\mathrm{Cl}_{\hat{\nu}}(\mathcal{C})$ is a proper subset of $\mathcal{P}(S)$. Below, we show that same conclusion holds also in the additive case.
Corollary 3.13.
(AC) Let $\mu:\mathcal{P}(\omega)\to\mathbb{R}$ be a normalized shift-invariant finitely additive map. Then $\mathrm{Cl}_{\hat{\mu}}(\mathcal{C})\neq\mathcal{P}(S)$.
Proof.
Suppose for the sake of contradiction that $\mathrm{Cl}_{\hat{\mu}}(\mathcal{C})=\mathcal{P}(S)$. We will use a Vitaly–type construction. Let $\lambda:\omega\to\omega$ be the shift map $n\mapsto n+1$. This induces a homeomorphism $\tau:S\to S$. Observe that, for each clopen $B=\phi(\pi(A))$ with $A\in\Sigma$, $\tau[B]=\phi(\pi(\lambda[A])))$ and $\mu(A)=\mu(\lambda[A])$. Consequently $\hat{\mu}$ is $\tau$-invariant. Now, let $\sim$ be the equivalence relation on $S$ so that $x\sim y$ if and only if $x=\tau^{k}(y)$ for some $k\in\mathbb{Z}$. Using the Axiom of Choice, we can construct a set $V\subseteq S$ which selects exactly one element from each equivalence class. It follows by Lemma 3.11 that $\hat{\mu}$ is $\sigma$-additive on $\mathscr{M}_{\hat{\mu}}=\mathrm{Cl}_{\hat{\mu}}(\mathcal{C})=\mathcal{P}(S)$. Taking into account that $\{\tau^{k}[V]:k\in\mathbb{Z}\}$ is a partition of $S$ by construction, we obtain that
$1=\hat{\mu}(S)=\sum_{k\in\mathbb{Z}}\hat{\mu}(\tau^{k}[V])=\sum_{k\in\mathbb{Z}}\hat{\mu}(V),$
which is impossible for every value of $\hat{\mu}(V)$. ∎
Examples of finitely additive maps $\mu$ satisfying the hypothesis of Corollary 3.13 (which are upper densities as in Section 2.2) can be found in [38, Remark 3].
Several properties of the analogous families with respect to $\tilde{\nu}$ are listed below. To this aim, given a set $B\subseteq S$, we denote by $B^{\circ}$ its interior and $\overline{B}$ its closure.
Lemma 3.14.
The following properties hold:
  1. (i)
    $\tilde{\nu}(\partial B)=0$ for every $B\in\mathrm{Cl}_{\tilde{\nu}}(\mathcal{C})$;
  2. (ii)
    $\partial B$, $B^{\circ}$, and $\overline{B}$ are elements of $\mathrm{Cl}_{\tilde{\nu}}(\mathcal{C})$ for every $B\in\mathrm{Cl}_{\tilde{\nu}}(\mathcal{C})$;
  3. (iii)
    $\tilde{\nu}(\overline{B})=\tilde{\nu}(B)$ for every $B\in\mathrm{Cl}_{\tilde{\nu}}(\mathcal{C})$;
  4. (iv)
    $\mathcal{N}_{\tilde{\nu}}=\mathrm{nwd}(S)\cap\mathrm{Cl}_{\tilde{\nu}}(\mathcal{C})$;
  5. (v)
    A subset of $S$ is the boundary of some element of $\mathrm{Cl}_{\tilde{\nu}}(\mathcal{C})$ if and only if it is closed and $\tilde{\nu}-$null;
  6. (vi)
    If $B\subseteq S$ is a compact $G_{\delta}$ set, then $B\in\sigma(\mathcal{C})$;
  7. (vii)
    For every $B\in\mathrm{Cl}_{\hat{\nu}}(\mathcal{C})$, there are sets $A,C\in\sigma(\mathcal{C})$ such that $A\subseteq B\subseteq C$ and $\hat{\nu}(C\setminus A)=0$;
  8. (viii)
    For every $B\in\mathrm{Cl}_{\tilde{\nu}}(\mathcal{C})$, there are sets $A,C\subseteq S$ such that AA is a countable union of clopen sets, $C$ is a countable intersection of clopen sets, $A\subseteq B\subseteq C$, and $\tilde{\nu}(C\setminus A)=0$.
Proof.
(i) Fix $B\in\mathrm{Cl}_{\tilde{\nu}}(\mathcal{C})$ and $k\in\omega$. By Remark 3.8, there are $A_{k},C_{k}\in\mathcal{C}$ with $A_{k}\subseteq B\subseteq C_{k}$ and $\tilde{\nu}(C_{k}\setminus A_{k})\leq 2^{-k}$. But since $A_{k}$ and $C_{k}$ are clopen, we have $A_{k}\subseteq B^{\circ}$ and $\overline{B}\subseteq C_{k}$. Hence $\partial B\subseteq C_{k}\setminus A_{k}$ and $\tilde{\nu}(\partial B)\leq 2^{-k}$. The conclusion follows by the arbitrariness of $k$.
(ii) Since any $\tilde{\nu}-$null set is the $d_{\tilde{\nu}}-$limit of a sequence of empty sets, we have
(21) $\mathcal{N}_{\tilde{\nu}}\subseteq\mathrm{Cl}_{\tilde{\nu}}(\mathcal{C}).$
Now, fix $B\in\mathrm{Cl}_{\tilde{\nu}}(\mathcal{C})$. Since $\partial B\in\mathcal{N}_{\tilde{\nu}}$ by item (i), then $\partial B\in\mathrm{Cl}_{\tilde{\nu}}(\mathcal{C})$. Since $\mathrm{Cl}_{\tilde{\nu}}(\mathcal{C})$ is a field of sets, then also $B^{\circ}=B\setminus\partial B$ and $\overline{B}=B\cup\partial B$ belong to $\mathrm{Cl}_{\tilde{\nu}}(\mathcal{C})$.
(iii) Taking into account that $\tilde{\nu}$ is a submeasure, this is immediate by item (i).
(iv) Pick $B\in\mathrm{nwd}(S)\cap\mathrm{Cl}_{\tilde{\nu}}(\mathcal{C})$. Then $\overline{B}\in\mathrm{Cl}_{\tilde{\nu}}(\mathcal{C})$ by item (ii). In addition $\overline{B}=\partial B$ since $B$ is nowhere dense, hence $\tilde{\nu}(B)=0$ by item (i). This shows that $\mathrm{nwd}(S)\cap\mathrm{Cl}_{\tilde{\nu}}(\mathcal{C})\subseteq\mathcal{N}_{\tilde{\nu}}$.
To show the converse inclusion, pick $B\in\mathcal{N}_{\tilde{\nu}}$. Observe that $0=\tilde{\nu}(B)=\tilde{\nu}(\overline{B})$ by item (iii), hence $\overline{B}$ cannot contain a nonempty clopen set (since they have positive $\tilde{\nu}-$measure). Hence $B$ is nowhere dense. Taking into account also (21), we obtain that $\mathcal{N}_{\tilde{\nu}}\subseteq\mathrm{nwd}(S)\cap\mathrm{Cl}_{\tilde{\nu}}(\mathcal{C})$.
(v) Suppose that $B\subseteq S$ is closed and $\tilde{\nu}-$null. Thanks to item (iv), we have $B$ is a closed set in $\mathrm{nwd}(S)\cap\mathrm{Cl}_{\tilde{\nu}}(\mathcal{C})$. Then $B$ is the boundary of the dense open set $B^{c}\in\mathrm{Cl}_{\tilde{\nu}}(\mathcal{C})$. This proves the If part. The Only If part follows by item (i).
(vi) Pick a compact $G_{\delta}$ set $B\subseteq S$. Since $S$ is compact Hausdorff, it follows by [16, Corollary 1.5.2 and Theorem 3.1.9] that there exists a continuous function $f:S\to[0,1]$ such that $B=f^{-1}(0)$. Let $\mathscr{F}$ be the family of functions $g\in\mathbb{R}^{S}$ such that the image $g[S]$ is finite and $g^{-1}(x)\in\mathcal{C}$ for each $x\in\mathbb{R}$. Of course, such maps belong to the Banach lattice $C(S)$ of continuous functions $S\to\mathbb{R}$. Since $\mathscr{F}$ is dense in $C(S)$ (with respect to uniform convergence) by Stone–Weierstrass theorem, see e.g. [16, Theorem 3.2.21], it follows that $f\in C(S)$ is the limit of a sequence of elements in $\mathscr{F}$. Taking into account the maps in $\mathscr{F}$ are $\sigma(\mathcal{C})$-measurable, we conclude that $f$ is $\sigma(\mathcal{C})$-measurable, hence $B\in\sigma(\mathcal{C})$.
(vii) Fix $B\in\mathrm{Cl}_{\hat{\nu}}(\mathcal{C})$ and $\epsilon>0$. By definition of $\mathrm{Cl}_{\hat{\nu}}(\mathcal{C})$, there is $D\in\mathcal{C}$ and a sequence $(E_{n})_{n\in\omega}$ in $\mathcal{C}$ such that $B\triangle D\subseteq\bigcup_{n}E_{n}$ and $\sum_{n}\hat{\nu}(E_{n})<\epsilon$. Define $E:=\bigcup_{n}E_{n}\in\sigma(\mathcal{C})$, so that
$D\setminus E\subseteq B\subseteq D\cup E\quad\text{ and }\quad\hat{\nu}((D\cup E)\setminus(D\setminus E))=\hat{\nu}(E)\leq\sum_{n\in\omega}\hat{\nu}(E_{n})<\epsilon.$
By the arbitrariness of $\epsilon>0$, we conclude that there exist $A,C\in\sigma(\mathcal{C})$ such that $A\subseteq B\subseteq C$ and $\hat{\nu}(C\setminus A)=0$.
(viii) Fix $B\in\mathrm{Cl}_{\tilde{\nu}}(\mathcal{C})$ and, for each $k\in\omega$, choose $A_{k}$ and $C_{k}$ as in the proof of item (i). Then $A^{c}:=\bigcap_{k}A_{k}^{c}$ and $C:=\bigcap_{k}C_{k}$ are countable intersections of clopen sets since $A_{k},C_{k}\in\mathcal{C}$ for each $k\in\omega$. Therefore $A=\bigcup_{k}A_{k}\subseteq B\subseteq C$ and $\tilde{\nu}(C\setminus A)=0$. ∎
We conclude with the proofs of Theorem 2.9 and Corollary 2.10.
Proof of Theorem 2.9.
Let us start defining also the following condition:
  1. (a1)
    For every increasing sequence $(A_{n}:n\in\omega)$ in $\overline{\Sigma}$ such that $\sum_{n}\nu^{\star}(A_{n+1}\setminus A_{n})<\infty$, there exists $A\in\overline{\Sigma}$ such that $\nu^{\star}(A_{n}\setminus A)=0$ for all $n\in\omega$ and $\lim_{n}\nu^{\star}(A\setminus A_{n})=0$.
Thus, it will be enough to prove all the implications depicted in Figure 1 below (here, the implication (a3) $\implies$ (a1) may seem superfluous, but it is a step in the proof of (b4) $\implies$ (c1) ).
(a1)(a1)(a2)(a3)(b1)(b2)(b4)(b3)(c2)(c3)(c1)(d1)(d6)(d3)(d5)(d2)(e1)(d4)(e2)(d7)(e3)
Figure 1. Implications in the proof of Theorem 2.9.
(a1) $\Longleftrightarrow$ (a1). It follows by Theorem 2.1.
(a1) $\implies$ (a2). Pick an increasing sequence $(A_{n}:n\in\omega)$ in $\Sigma$ such that $\sum_{n}\nu(A_{n+1}\setminus A_{n})<\infty$, and fix $A\in\overline{\Sigma}$ as in item (a1). By the definition of $\nu^{\star}$, for each $m\in\omega$ there exists $E_{m}\in\Sigma$ such that $A\setminus A_{m}\subseteq E_{m}$ and $\nu^{\star}(E_{m})\leq\nu^{\star}(A\setminus A_{m})+2^{-m}$. Define the decreasing sequence $(D_{m}:m\in\omega)$ with values in $\Sigma$ by $D_{m}:=\bigcap_{k\leq m}(A_{k}\cup E_{k})$ for each $m\in\omega$. Since $A\subseteq A_{k}\cup E_{k}$ for each $k\in\omega$, we have that $A\subseteq D_{m}$ for all $m\in\omega$. This implies that $\nu(A_{n}\setminus D_{m})\leq\nu^{\star}(A_{n}\setminus A)=0$ for all $n,m\in\omega$. Lastly, since
$\nu(D_{n}\setminus A_{n})\leq\nu((A_{n}\cup E_{n})\setminus A_{n})\leq\nu(E_{n})\leq\nu^{\star}(A\setminus A_{n})+2^{-n}$
for all $n\in\omega$, it follows that $\lim_{n}\nu(D_{n}\setminus A_{n})=0$.
(a2) $\implies$ (a3). Pick an increasing sequence $(A_{n}:n\in\omega)$ in $\Sigma$ such that $\sum_{n}\nu(A_{n+1}\setminus A_{n})<\infty$. By item (a2), there exists $(D_{m}:m\in\omega)$ in $\Sigma$ such that $\nu(A_{n}\setminus D_{m})=0$ for all $n,m\in\omega$ and $\lim_{n}\nu(D_{n}\setminus A_{n})=0$. We claim that
(22) $C:=\bigcup_{k\in\omega}(A_{k}\cap D_{k})$
satisfies the required property in item (a3). Since $A_{n}\setminus C\subseteq A_{n}\setminus D_{n}$, it follows by monotonicity that $\nu^{\star}(A_{n}\setminus C)=0$ for each $n\in\omega$. Moreover, since $C\subseteq A_{m}\cup D_{m}$ for all $m\in\omega$ and $\nu^{\star}$ is a submeasure, we get that $\nu^{\star}(C\setminus A_{n})\leq\nu((A_{n}\cup D_{n})\setminus A_{n})=\nu(D_{n}\setminus A_{n})$, which converges to 0 as $n\to\infty$.
(a3) $\implies$ (a1). Pick an increasing $(A_{n}:n\in\omega)$ in $\overline{\Sigma}$ such that $\sum_{n}\nu^{\star}(A_{n+1}\setminus A_{n})<\infty$. For each $n\in\omega$, there exists $B_{n}\in\Sigma$ such that $\nu^{\star}(A_{n}\bigtriangleup B_{n})<2^{-n}$. Then there is also $C_{n}\in\Sigma$ such that $A_{n}\bigtriangleup B_{n}\subseteq C_{n}$ and $\nu(C_{n})<2^{-n}$. At this point, set $D_{n}:=A_{n}\setminus C_{n}$ for each $n\in\omega$ and note that $D_{n}\subseteq A_{n}$ and $D_{n}=B_{n}\setminus C_{n}\in\Sigma$. Hence the sequence $(E_{n}:n\in\omega)$ defined by $E_{n}:=\bigcup_{k\in n+1}D_{k}$ is increasing with values in $\Sigma$. Now, observing that $E_{n}\subseteq A_{n}$ and $A_{n}\setminus E_{n}\subseteq C_{n}$, we have
$\begin{split}\nu(E_{n+1}\setminus E_{n})&\leq\nu^{\star}(E_{n+1}\setminus A_{n+1})+\nu^{\star}(A_{n+1}\setminus A_{n})+\nu^{\star}(A_{n}\setminus E_{n})\\ &\leq\nu^{\star}(A_{n+1}\setminus A_{n})+\nu^{\star}(C_{n})<\nu^{\star}(A_{n+1}\setminus A_{n})+2^{-n}.\end{split}$
Hence $\sum_{n}\nu(E_{n+1}\setminus E_{n})<\infty$. It follows, thanks to item (a3), that there exists $A\subseteq X$ such that $\nu^{\star}(E_{n}\setminus A)=0$ for all $n\in\omega$ and $\lim_{n}\nu^{\star}(A\setminus E_{n})=0$.
To complete the implication, it is enough to show that the set AA satisfies the property given in item (a1). First, observe that
$\limsup_{n\to\infty}d_{\nu^{\star}}(E_{n},A)\leq\limsup_{n\to\infty}\,(\nu^{\star}(E_{n}\setminus A)+\nu^{\star}(A\setminus E_{n}))=0.$
Therefore $A\in\overline{\Sigma}$. Moreover, since $A_{n}\triangle E_{n}=A_{n}\setminus E_{n}\subseteq A_{n}\setminus D_{n}\subseteq C_{n}$ for all $n\in\omega$, we have
$\begin{split}\limsup_{n\to\infty}\nu^{\star}(A_{n}\triangle A)&\leq\limsup_{n\to\infty}\nu^{\star}(A_{n}\triangle E_{n})+\limsup_{n\to\infty}\nu^{\star}(E_{n}\triangle A)\\ &\leq\limsup_{n\to\infty}\nu(C_{n})+\limsup_{n\to\infty}\nu^{\star}(E_{n}\triangle A)=0.\end{split}$
This implies that $\lim_{n}\nu^{\star}(A\setminus A_{n})=0$ and, since $(A_{n}:n\in\omega)$ is increasing, also $\nu^{\star}(A_{n}\setminus A)\leq\limsup_{k}\nu^{\star}(A_{k}\setminus A)=0$ for each $n\in\omega$.
(a3) $\implies$ (b4). Pick $(A_{n}:n\in\omega)$ as in item (b4). Hence by (a3) there exists $C\subseteq X$ such that $\nu^{\star}(A_{n}\setminus C)=0$ for all $n\in\omega$ and $\lim_{n}\nu^{\star}(C\setminus A_{n})=0$. Now, observe that
(23) $\forall n\in\omega,\quad\nu(A_{n})\leq\nu^{\star}(A_{n}\cup C)\leq\nu^{\star}(C)+\nu^{\star}(A_{n}\setminus C)=\nu^{\star}(C).$
In addition, we have that $\nu^{\star}(C)\leq\nu^{\star}(C\setminus A_{n})+\nu^{\star}(A_{n}\cap C)$ for each $n\in\omega$. Together with $\nu^{\star}(A_{n}\setminus C)=0$ and Inequality (23), we get
$0\leq\nu^{\star}(C)-\nu^{\star}(A_{n})\leq\nu^{\star}(C\setminus A_{n})+\nu^{\star}(A_{n}\cap C)-\nu^{\star}(A_{n})=\nu^{\star}(C\setminus A_{n})$
for each $n\in\omega$. Considering that $\lim_{n}\nu^{\star}(C\setminus A_{n})=0$ and $(A_{n}:n\in\omega)$ is increasing, we conclude that $\nu^{\star}(C)=\sup_{n}\nu(A_{n})$.
(b1) $\implies$ (b2). This is clear.
(b2) $\implies$ (b3). Pick an increasing sequence $(A_{n}:n\in\omega)$ in $\Sigma$ such that $A_{0}=\emptyset$ and $\sum_{n}\nu(A_{n+1}\setminus A_{n})<\infty$, and define the countable union of clopen sets
$U:=\bigcup_{n\in\omega}\phi(\pi(A_{n})).$
Considering that $(A_{n+1}\setminus A_{n}:n\in\omega)\in\Delta(U)$, it follows that
(24) $\hat{\nu}(\overline{U})=\hat{\nu}(U)\leq\sum_{n\in\omega}\nu(A_{n+1}\setminus A_{n})<\infty.$
By the definition of $\hat{\nu}$, for each $k\in\omega$ there exists a sequence $(C_{k,n}:n\in\omega)\in\Delta(\overline{U})$ such that $\sum_{n}\nu(C_{k,n})\leq\hat{\nu}(\overline{U})+2^{-k}$. Taking into account that $(\phi(\pi(C_{k,n})):n\in\omega)$ is an open cover of the compact set $\overline{U}$, there exists $I_{k}\in\mathrm{Fin}$ such that $\overline{U}\subseteq\bigcup_{n\in I_{k}}\phi(\pi(C_{k,n}))$. Thus, it is sufficient to show that the decreasing sequence $(D_{m}:m\in\omega)$ given by
$\forall m\in\omega,\quad D_{m}:=\bigcap_{k\in m+1}\bigcup_{n\in I_{k}}C_{k,n}$
satisfies the required properties.
To this aim, note that $D_{m}\in\Sigma$ and $\overline{U}\subseteq\phi(\pi(D_{m}))$ for each $m\in\omega$. In addition, for each $n,m\in\omega$, we have $\phi(\pi(A_{n}))\setminus\phi(\pi(D_{m}))\subseteq U\setminus\overline{U}=\emptyset$, hence $\pi(A_{n})\leq\pi(D_{m})$, which implies $\nu(A_{n}\setminus D_{m})=0$. Lastly, for each $k\in\omega$, using also Claim 1 and the fact that $(A_{k},A_{k+1}\setminus A_{k},A_{k+2}\setminus A_{k+1},\ldots)\in\Delta(U)$, we obtain
$\begin{split}0\leq\nu(D_{k})-\nu(A_{k})&=\hat{\nu}(\phi(\pi(D_{k}))-\nu(A_{k})\\ &\leq\hat{\nu}(\overline{U})-\nu(A_{k})+2^{-k}\\ &=\hat{\nu}(U)-\nu(A_{k})+2^{-k}\leq\sum_{n\geq k}\nu(A_{n+1}\setminus A_{n})+2^{-k}.\end{split}$
The implication follows by the fact the right hand side converges to 0 as $k\to\infty$.
(b3) $\implies$ (b4). Pick $(A_{n}:n\in\omega)$ and $(D_{m}:m\in\omega)$ as in item (b3), and define $C:=\bigcup_{k}(A_{k}\cap D_{k}).$ Since $A_{n}\setminus C\subseteq A_{n}\setminus D_{n}$, it follows by monotonicity that $\nu^{\star}(A_{n}\setminus C)=0$ for each $n\in\omega$. Now, observe that by Inequality (23) we get $\sup_{n}\nu(A_{n})\leq\nu^{\star}(C)$. Conversely, since $C\subseteq A_{m}\cup D_{m}$ for all $m\in\omega$ and $\nu^{\star}$ is a submeasure, we get that $\nu^{\star}(C)\leq\nu(A_{m}\setminus D_{m})+\nu(D_{m})=\nu(D_{m})$. Hence by item (b3) we conclude that $\nu^{\star}(C)\leq\inf_{m}\nu(D_{m})=\sup_{n}\nu(A_{n})$.
(b4) $\implies$ (b1). Fix a set $U\subseteq S$. Since $\hat{\nu}$ is a capacity, it is obvious that $\hat{\nu}(\overline{U})=\infty$ if $\hat{\nu}(U)=\infty$. Hence, let us suppose hereafter that $\hat{\nu}(U)<\infty$.
Fix $k\in\omega$. By the definition of $\hat{\nu}$, there exists $(F_{k,n}:n\in\omega)\in\Delta(U)$ such that
(25) $\sum_{n\in\omega}\nu(F_{k,n})\leq\hat{\nu}(U)+2^{-k}.$
Assume without loss of generality that $F_{k,0}=\emptyset$. For each $n\in\omega$, define $A_{k,n}:=F_{k,0}\cup\cdots\cup F_{k,n}$, so that $(A_{k,n}:n\in\omega)$ is increasing in $\Sigma$ and $\sum_{n}\nu(A_{k,n+1}\setminus A_{k,n})\leq\sum_{n}\nu(F_{k,n})<\infty$, where the latter follows by (25). By item (b4), there exists $C_{k}\subseteq X$ such that $\nu^{\star}(A_{k,n}\setminus C_{k})=0$ for all $n\in\omega$ and $\sup_{n}\nu(A_{k,n})=\nu^{\star}(C_{k})$. In particular, there exists $n_{k}\in\omega$ such that $\nu^{\star}(C_{k})\leq\nu(A_{k,n_{k}})+2^{-k}$. In addition, by the definition of $\nu^{\star}$, there exists $D_{k}\in\Sigma$ such that $C_{k}\subseteq D_{k}$ and $\nu(D_{k})\leq\nu^{\star}(C_{k})+2^{-k}$.
Claim 2.
$\nu^{\star}(C_{k})\leq\hat{\nu}(U)+2^{1-k}$.
Proof.
Since $A_{k,n_{k}}=\bigcup_{n\in n_{k}}A_{k,n+1}\setminus A_{k,n}$ and $\nu$ is a submeasure, we obtain that
$\nu(A_{k,n_{k}})\leq\sum_{n\in\omega}\nu(A_{k,n+1}\setminus A_{k,n})\leq\sum_{n\in\omega}\nu(F_{k,n}).$
The claim follows taking into account also Inequality (25) and that $\nu^{\star}(C_{k})\leq\nu(A_{k,n_{k}})+2^{-k}$. ∎
Claim 3.
$\overline{U}\subseteq\phi(\pi(D_{k}))$.
Proof.
Observe, for each $n\in\omega$, that $\nu(A_{k,n}\setminus D_{k})\leq\nu^{\star}(A_{k,n}\setminus C_{k})=0$ which implies that $\pi(A_{k,n})\leq\pi(D_{k})$, hence $\phi(\pi(A_{k,n}))\subseteq\phi(\pi(D_{k}))$. Since $(F_{k,n}:n\in\omega)\in\Delta(U)$, we obtain
$U\subseteq\bigcup_{n\in\omega}\phi(\pi(F_{k,n}))\subseteq\bigcup_{n\in\omega}\phi(\pi(A_{k,n}))\subseteq\phi(\pi(D_{k})).$
The claim follows by the fact that the set on the right hand side is closed in $S$. ∎
It follows by Claim 1, Claim 2, Claim 3, and the above observations that
$\begin{split}\nu^{\star}(C_{k})&\leq\hat{\nu}(U)+2^{1-k}\leq\hat{\nu}(\overline{U})+2^{1-k}\\ &\leq\hat{\nu}(\phi(\pi(D_{k})))+2^{1-k}=\nu(D_{k})+2^{1-k}\leq\nu^{\star}(C_{k})+2^{2-k}.\end{split}$
By letting $k\to\infty$, it follows that $\hat{\nu}(U)=\hat{\nu}(\overline{U})$.
(b4) $\implies$ (c1). Pick an increasing sequence $(A_{n}:n\in\omega)$ of subsets of $X$ such that $\sum_{n}\nu^{\star}(A_{n+1}\setminus A_{n})<\infty$. Then $(A_{n}:n\in\omega)$ is a Cauchy sequence and in $(\mathcal{P}(X),d_{\nu^{\star}})$ and we can suppose without loss of generality that $\nu^{\star}(A_{n+1}\setminus A_{n})<2^{-n}$ for all $n\in\omega$. By the definition of $\nu^{\star}$, for each $n\in\omega$ there exists $V_{n}\in\Sigma$ such that $A_{n+1}\setminus A_{n}\subseteq V_{n}$ and $\nu(V_{n})<2^{-n}$. Fix $k\in\omega$ and define $(W_{k,n}:n\in\omega)$ by
$\forall n\in\omega,\quad W_{k,n}:=V_{k}\cup V_{k+1}\cup\cdots\cup V_{k+n}.$
Of course, $W_{k,n}\in\Sigma$ and $\sum_{n}\nu(W_{k,n+1}\setminus W_{k,n})\leq\sum_{n}\nu(V_{n})<\infty$. It follows by item (b4) that there exists $C_{k}\subseteq X$ such that $\nu^{\star}(W_{k,n}\setminus C_{k})=0$ for all $n\in\omega$ and $\nu^{\star}(C_{k})=\sup_{n}\nu(W_{k,n})\leq\sum_{n\geq k}\nu(V_{n})<2^{1-k}$.
Claim 4.
$\nu^{\star}(A_{i}\setminus(A_{j}\cup C_{j}))=0$ for all $i,j\in\omega$.
Proof.
Fix $i,j\in\omega$. If $i\leq j$ the claim is clear since $A_{i}\subseteq A_{j}$. Hence, suppose hereafter that $i>j$. Then
$A_{i}\setminus(A_{j}\cup C_{j})\subseteq\bigcup_{k=j}^{i-1}(A_{k+1}\setminus A_{k})\setminus C_{j}\subseteq\bigcup_{k=j}^{i-1}V_{k}\setminus C_{j}\subseteq\bigcup_{n\in i-j}W_{j,n}\setminus C_{j}.$
The claim follows since $\nu^{\star}$ is a submeasure and $\nu^{\star}(W_{j,n}\setminus C_{j})=0$ for all $n\in\omega$. ∎
At this point, consider the decreasing sequence $(D_{m}:m\in\omega)$ of subsets of $X$ defined by $D_{m}:=\bigcap_{i\in m+1}(A_{i}\cup C_{i})$. Since $D_{m}\subseteq A_{m}\cup C_{m}$, we get by Claim 4
$\nu^{\star}(A_{n}\setminus D_{m})\leq\sum_{i\in m+1}\nu^{\star}(A_{n}\setminus(A_{i}\cup C_{i}))=0$
for all $n,m\in\omega$. Lastly, we have also
$\nu^{\star}(D_{n}\setminus A_{n})\leq\nu^{\star}((A_{n}\cup C_{n})\setminus A_{n})\leq\nu^{\star}(C_{n})<2^{1-n}$
for all $n\in\omega$. Therefore $\lim_{n}\nu^{\star}(D_{n}\setminus A_{n})=0$, and the pseudometric space $(\mathcal{P}(X),d_{\nu^{\star}})$ satisfies its corresponding item (a2). Since we already proved that (a1) $\Longleftrightarrow$ (a2) then $(\mathcal{P}(X),d_{\nu^{\star}})$ is complete.
(c2) $\implies$ (c3). This is clear.
(c1) $\implies$ (c2) and (c3) $\implies$ (a1). They follow by the fact that a closed subset of a complete pseudometric space is a complete subspace.
(b4) $\implies$ (d1). Pick a set $U\subseteq S$. Since $\hat{\nu}\leq\tilde{\nu}$, we can suppose hereafter that $\hat{\nu}(U)<\infty$. Proceeding verbatim and with the same notations as in the proof of the implication (b4) $\implies$ (b1), for each $k\in\omega$ it is possible to pick a set $D_{k}\in\Sigma$ such that $U\subseteq\overline{U}\subseteq\phi(\pi(D_{k}))$ and $\nu(D_{k})\leq\hat{\nu}(U)+2^{2-k}$. It follows that $\tilde{\nu}(U)\leq\inf_{k}\nu(D_{k})\leq\hat{\nu}(U)$, which proves the converse inequality $\tilde{\nu}\leq\hat{\nu}$.
(d1) $\Longleftrightarrow$ (d2). The equivalence follows by the facts that $\hat{\nu}\leq\tilde{\nu}$, $\hat{\nu}$ is $\sigma$-subadditive, and, by definition, it is also the maximum $\sigma$-subadditive (and monotone) extension of the submeasure $\phi(\pi(A))\mapsto\nu(A)$ to $\mathcal{P}(S)$, cf. Claim 1, whereas $\tilde{\nu}$ is the maximum monotone (and subadditive) extension of the same submeasure.
(d1) $\implies$ (d5). This is obvious.
(d5) $\implies$ (d3). Since $\hat{\nu}\leq\tilde{\nu}$, we have $\mathrm{Cl}_{\tilde{\nu}}(\mathcal{C})\subseteq\mathrm{Cl}_{\hat{\nu}}(\mathcal{C})$. Conversely, it follows by Remark 3.8 that item (d3) holds if and only if $\mathrm{Cl}_{\hat{\nu}}(\mathcal{C})\subseteq\mathrm{Cl}_{\tilde{\nu}}(\mathcal{C})$.
(d3) $\implies$ (e1). Pick $B\in\mathrm{Cl}_{\hat{\nu}}(\mathcal{C})$. By hypothesis, for each $k\in\omega$, there exist $A_{k},C_{k}\in\mathcal{C}$ such that $A_{k}\subseteq B\subseteq C_{k}$ and $\tilde{\nu}(C_{k}\setminus A_{k})<2^{-k}$. Since $A_{k}$ and $C_{k}$ are clopen, the interior of $B$ contains $A_{k}$ and the closure of $B$ is contained in $C_{k}$. Hence $\partial B\subseteq C_{k}\setminus A_{k}$. It follows that $\tilde{\nu}(\partial B)\leq\inf_{k}\tilde{\nu}(C_{k}\setminus A_{k})=0$.
(e1) $\implies$ (e2). This is clear since $\hat{\nu}\leq\tilde{\nu}$.
(d1) $\implies$ (d4). This follows by Lemma 3.14(viii).
(d4) $\implies$ (e2). Fix $B\in\mathrm{Cl}_{\hat{\nu}}(\mathcal{C})$. By hypothesis, it is possible to pick sets $A,B\subseteq S$ such that AA is open, $C$ is closed, $A\subseteq B\subseteq C$ and $\hat{\nu}(C\setminus A)=0$. It follows that $A\subseteq B^{\circ}$ and $\overline{B}\subseteq C$. Since $\partial B\subseteq\overline{B}\setminus B^{\circ}\subseteq C\setminus A$, we conclude that $\hat{\nu}(\partial B)\leq\hat{\nu}(C\setminus A)=0$.
(d1) $\implies$ (d6). This follows by Lemma 3.14(v).
(d1) $\implies$ (d7). It follows by Lemma 3.14(ii) and Lemma 3.14(iv).
(d7) $\implies$ (e2). This follows by the fact that every $\partial B$ is nowhere dense.
(d6) $\implies$ (e2) $\implies$ (e3). They are obvious.
(e3) $\implies$ (a2). Pick an increasing sequence $(A_{n}:n\in\omega)$ with values in $\Sigma$ such that $A_{0}=\emptyset$ and $\sum_{n}\nu(A_{n+1}\setminus A_{n})<\infty$, and define
$U:=\bigcup_{n\in\omega}\phi(\pi(A_{n})).$
Then $U$ is a countable union of clopen sets. Hence by item (e3) we get $\hat{\nu}(\partial U)=0$. Moreover, $\hat{\nu}(U)<\infty$, as in (24). By the definition of $\hat{\nu}$, for each $k\in\omega$ there exists a sequence $(G_{k,n}:n\in\omega)\in\Delta(\overline{U}\setminus\phi(\pi(A_{k})))$ such that $\sum_{n}\nu(G_{k,n})\leq\hat{\nu}(\overline{U}\setminus\phi(\pi(A_{k})))+2^{-k}$. Taking into account that $(\phi(\pi(G_{k,n})):n\in\omega)$ is an open cover of the compact set $\overline{U}\setminus\phi(\pi(A_{k}))$, there exists $J_{k}\in\mathrm{Fin}$ such that $\overline{U}\setminus\phi(\pi(A_{k}))\subseteq\bigcup_{n\in J_{k}}\phi(\pi(G_{k,n}))$. Now, for each $k\in\omega$, define the set
$V_{k}:=U\cup\bigcup_{n\in J_{k}}\phi(\pi(G_{k,n}))$
Considering that $\{\phi(\pi(G_{k,n})):n\in J_{k}\}$ is also a cover of $\partial U$, it follows that $V_{k}=\overline{U}\cup\bigcup_{n\in J_{k}}\phi(\pi(G_{k,n}))$. Therefore each $V_{k}$ is a clopen subset of $S$ containing $\overline{U}$. Replacing, if necessary, $V_{k}$ with $\bigcap_{j\in k+1}V_{j}$, we can also assume without loss of generality that $(V_{k}:k\in\omega)$ is decreasing. Also, since $V_{k}\in\mathcal{C}$ and $\phi:\Sigma/\nu\to\mathcal{C}$ is a Boolean isomorphism, there exists a decreasing sequence $(D_{k}:k\in\omega)$ in $\Sigma$ such that
$\forall k\in\omega,\quad V_{k}=\phi(\pi(D_{k})).$
We claim that the above sequence satisfies the required properties. In fact, for each $n,m\in\omega$, we have $\phi(\pi(A_{n}))\setminus\phi(\pi(D_{m}))=\phi(\pi(A_{n}))\setminus V_{m}\subseteq U\setminus\overline{U}=\emptyset$, hence $\pi(A_{n})\leq\pi(D_{m})$, which implies $\nu(A_{n}\setminus D_{m})=0$. Lastly, using also Claim 1, for each $k\in\omega$ we get
$\begin{split}\nu(D_{k}\setminus A_{k})&=\hat{\nu}(\phi(\pi(D_{k}))\setminus\phi(\pi(A_{k})))\\ &=\hat{\nu}\left(U\cup\bigcup_{n\in J_{k}}\phi(\pi(G_{k,n}))\setminus\phi(\pi(A_{k}))\right)\leq\hat{\nu}(\phi(\pi(\bigcup_{n\in J_{k}}G_{k,n})))\\ &\leq\sum_{n\in J_{k}}\nu(G_{k,n})\leq\sum_{n\in\omega}\nu(G_{k,n})\leq\hat{\nu}(\overline{U}\setminus\phi(\pi(A_{k})))+2^{-k}\\ &\leq\hat{\nu}(U\setminus\phi(\pi(A_{k})))+\hat{\nu}(\partial U)+2^{-k}=\hat{\nu}\left(\bigcup_{n\geq k}\phi(\pi(A_{n+1}\setminus A_{n}))\right)+2^{-k}\\ &\leq\sum_{n\geq k}\nu(A_{n+1}\setminus A_{n})+2^{-k}.\end{split}$
Therefore $\lim_{k}\nu(D_{k}\setminus A_{k})=0$. ∎
Proof of Corollary 2.10.
(i) $\implies$ (ii). Suppose that $(\Sigma,d_{\nu})$ is complete, and fix $B\in\mathrm{Cl}_{\hat{\nu}}(\mathcal{C})$. By the equivalence (a1) $\Longleftrightarrow$ (d3) in Theorem 2.9, for each $n\in\omega$, there exist $A_{n},C_{n}\in\Sigma$ such that $\phi(\pi(A_{n}))\subseteq B\subseteq\phi(\pi(C_{n}))$ and $\tilde{\nu}(\phi(\pi(C_{n}\setminus A_{n})))=\nu(C_{n}\setminus A_{n})<2^{-n}$. Hence, $\lim_{n}\tilde{\nu}(\phi(\pi(A_{n}))\bigtriangleup B)=0$. Without loss of generality, it is possible to assume that $(A_{n}:n\in\omega)$ is increasing and $(C_{n}:n\in\omega)$ is decreasing. It follows that
$\nu(A_{m}\setminus A_{n})\leq\sum_{k\in\omega}\nu(A_{n+k+1}\setminus A_{n+k})\leq\sum_{k\in\omega}\nu(C_{n+k}\setminus A_{n+k})<2^{1-n}$
for all $n,m\in\omega$ with $m\geq n$, hence $(A_{n}:n\in\omega)$ is $d_{\nu}$-Cauchy. Thus it is convergent to some $A\in\Sigma$, which can be rewritten as $\lim_{n}\tilde{\nu}(\phi(\pi(A_{n}))\bigtriangleup\phi(\pi(A)))=0$. Therefore $\tilde{\nu}(\phi(\pi(A))\bigtriangleup B)=0$.
(ii) $\implies$ (iii). This is clear since $\hat{\nu}\leq\tilde{\nu}$.
(iii) $\implies$ (iv). This is immediate.
[…]
Finally, in Remark 3.15 and Remark 3.16 below we provide a necessary condition and a sufficient condition for the completeness of $(\overline{\Sigma},d_{\nu^{\star}})$, respectively. However, in both cases, the implications cannot be reversed.
Remark 3.15.
By the equivalence (a1) $\Longleftrightarrow$ (d7) of Theorem 2.9, the following implication holds:
[…]
However, the converse implication of (26) fails in general. To this aim, let $(a_{n}:n\in\omega)$ be a strictly positive sequence of reals such that $\sum_{n}a_{n}=1$, and let $\nu:\mathcal{P}(\omega)\to\overline{\mathbb{R}}$ be the submeasure defined by
[…]
In particular, $\Sigma=\overline{\Sigma}=\mathcal{P}(\omega)$ and $\nu=\nu^{\star}$. We get by Proposition 2.6 that $(\overline{\Sigma},d_{\nu^{\star}})$ is not complete. Now, observe that $\nu(A)=0$ if and only if $A=\emptyset$, hence the Stone space $S$ is isomorphic to the classical Stone–Čech compactification of $\omega$, that is, the set of principal ultrafilters $P$ and free ultrafilters $F$ on $\omega$, so that $S=P\cup F$ (hence, $P$ and $F$ are identified with their images under isomorphism). Let us fix $B\subseteq S$. First, suppose that $B\subseteq P=\bigcup_{n}P_{n}$ where $P_{n}:=\phi(\pi(\{n\}))$ is the singleton containing the unique principal ultrafilter at $\pi(\{n\})$. Define $J:=\{j\in\omega:P_{j}\subseteq B\}$, so that $B=\bigcup_{j\in J}P_{j}$ and $(\{j\}:j\in J)\in\Delta(B)$. It follows that $\hat{\nu}(B)=\sum_{j\in J}\nu(\{j\})=\sum_{j\in J}a_{j}$, hence
(27) $\forall B\subseteq S,\quad B\cap F=\emptyset\implies 0\leq\hat{\nu}(B)\leq 1.$
In the opposite case that $B\cap F\neq\emptyset$, i.e., $B$ contains a free ultrafilter $x$. Pick an arbitrary sequence $(A_{n}:n\in\omega)\in\Delta(B)$. Then $x\in B\subseteq\bigcup_{n}\phi(\pi(A_{n}))$, so we can pick $m\in\omega$ such that $x\in\phi(\pi(A_{m}))$. Equivalently, this means that $\pi(A_{m})\in x$, which is possible only if $A_{m}$ is infinite since $x\notin P$. Considering that $\nu(A_{m})>1$, it follows that $\hat{\nu}(B)\geq\hat{\nu}(\{x\})\geq 1$. Together with the monotonicity of $\hat{\nu}$ and the fact by Claim 1 that $\hat{\nu}(S)=\nu(\omega)=2$, we obtain
(28) $\forall B\subseteq S,\quad B\cap F\neq\emptyset\implies 1\leq\hat{\nu}(B)\leq 2.$
At this point, pick $B\subseteq S$ with $\hat{\nu}(B)=0$. It follows by (27) and (28) that $B\subseteq P$. By the minimality argument above, we get $B=\emptyset$, that is, $\mathcal{N}_{\hat{\nu}}$$\{\emptyset\}$. Therefore $\mathcal{N}_{\hat{\nu}}$ is a proper subset of $\mathrm{nwd}(S)$, while $(\overline{\Sigma},d_{\nu^{\star}})$ is not complete.
Remark 3.16.
Since countable unions of clopen sets are open, one might considering the following streghtening of item (e3):
  1. (e4)
    $\hat{\nu}(\partial U)=0$ for every open set $U\subseteq S$.
Thanks to Theorem 2.9, the following implication holds:
[…]
However, the converse implication of (29) fails in general. In fact, in the complete pseudometric space given in Example 3.10 we proved that there exists an open set $U\subseteq S$ which is a countable union of clopen sets and $U\notin\mathcal{C}$. This implies that $\partial U\neq\emptyset$, so that $\hat{\nu}(\partial U)=1$. Therefore item (e4) does not hold.
3.1. Acknowledgments
The authors are grateful to K. P. S. Bhaskara Rao (Indiana University Northwest, US) and Ilijas Farah (York University, CA) for several helpful communications during the preparation of this manuscript.